### v2

# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

Abstract
This paper presents a novel approach to four fundamental problems in the field of cosmological perturbation theory. Firstly, the issue of gauge dependence has been addressed by demonstrating the existence of unique and gauge-invariant quantities corresponding to the actual perturbations. Secondly, the formation of primordial structures after decoupling of matter and radiation is dependent on the existence of local fluid flows resulting from local pressure gradients. To take pressure gradients into account, it is necessary to consider both the energy density and the particle number density. Thirdly, the novel relativistic perturbation theory applies to an open, flat, and closed Friedmann-Lemaître-Robertson-Walker universe. The derivation of the novel perturbation theory definitively reveals the inherent limitations of Newtonian gravitation as a framework for investigating cosmological density perturbations. Finally, the application of the perturbation theory to a flat universe demonstrates that, prior to decoupling, perturbations in both Cold Dark Matter and ordinary matter are coupled to perturbations in radiation. Therefore, the universe’s earliest structures formed only after decoupling, at which point local nonadiabatic random pressure fluctuations became a significant factor. Negative nonadiabatic pressure fluctuations resulted in a brief, rapid growth of density fluctuations until the total pressure fluctuations became positive. In contrast, positive nonadiabatic pressure fluctuations led to the formation of voids. Perturbations with masses about $2.2\times 10^{4}\,\text{M}_{\odot}$ became nonlinear already $60$ million years after the Big Bang, and perturbations with masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$ became nonlinear within about $600$ million years.
Cosmology; Early Universe; Structure Formation; Perturbation Theory; Gauge Problem.
pacs:
98.80.-k; 98.80.Cq; 97.10.Bt; 04.25.Nx.
I Introduction
The linearized Einstein equations and conservation laws of Friedmann-Lemaître-Robertson-Walker (flrw) universes are crucial to cosmology. They can explain the formation of all kinds of structures in the expanding universe, including stars, galaxies, and microwave background fluctuations.
Lifshitz [1] and Lifshitz & Khalatnikov [2] were the first researchers to develop a cosmological perturbation theory to study the evolution of density perturbations in the universe. They encountered the problem that the solutions of the linearized Einstein equations and conservation laws can be modified by linear coordinate transformations, which are known as gauge transformations. Consequently, the solutions are contingent upon the selection of a coordinate system, rendering them gauge-dependent and thus devoid of physical significance. This is the notorious gauge problem of cosmology, which demands urgent resolution. A multitude of potential solutions to this problem have been proposed in the literature. Stewart [3] noted in his paper, […] perturbation theory within general relativity is generally regarded as a somewhat arcane subject […] . This observation is further substantiated by the lack of progress observed to date in this field. The purpose of this paper is to elucidate the hitherto puzzling aspects of cosmological perturbation theory.
Before delving into an examination of the approaches outlined in the extant literature, it is imperative to acknowledge the fundamental criteria that a comprehensive and precise cosmological perturbation theory must fulfill. Primarily, as Bardeen elucidated in his seminal paper [4], a quantity signifying the real physical energy density perturbation must be manifestly gauge-invariant, that is, it must be defined independently of any reference system. Secondly, in the nonrelativistic limit, the linearized energy density constraint equation, when considered together with the linearized momentum constraint equation, must yield the time-independent Poisson equation of Newtonian theory with the aforementioned gauge-invariant quantity as the source term. Thirdly, in the nonrelativistic limit, it is essential that the freedom to select a coordinate system is present. That is to say, relativistic gauge transformations must converge to Newtonian gauge transformations in the nonrelativistic limit. Finally, and equally important, the perturbation theory must encompass all previously established and well-known physical solutions for large-scale density perturbations.
The present paper proposes a novel cosmological perturbation theory that has been demonstrated to satisfy all four of the aforementioned criteria. A review of the existing literature reveals that there is no theory that meets all four of the above criteria.
Bardeen [4] was the first to demonstrate that the linearized Einstein equations and conservation laws can be recast into evolution equations that have solutions independent of the choice of a coordinate system by using gauge-invariant quantities. In his paper, Bardeen constructed two distinct gauge-invariant quantities for the perturbation to the energy density, designated as $\epsilon_{m}$ and $\epsilon_{g}$. In the limit of small scales, these quantities converge to the usual gauge-dependent quantity, denoted as $\varepsilon_{{\scriptscriptstyle(1)}}$. Bardeen presupposed that on small scales, Newtonian gravitation is a valid description of cosmological density perturbations, and therefore, the gauge-dependent quantity, $\varepsilon_{{\scriptscriptstyle(1)}}$, becomes equal to its Newtonian counterpart. It is, however, a well-known fact that in Newtonian gravitation the coordinates of time and space can be chosen at will. Correspondingly, this must also be true in the nonrelativistic limit of a cosmological perturbation theory. It can thus be concluded that $\varepsilon_{{\scriptscriptstyle(1)}}$, which depends on the choice of coordinates in general relativity, is also dependent on the choice of coordinates in Newtonian gravitation. Thus, quantities that lack physical significance in general relativity are similarly devoid of physical meaning in Newtonian gravitation. It follows that the small-scale limit is not equivalent to the nonrelativistic limit. In consequence, the quantities $\epsilon_{m}$ and $\epsilon_{g}$ do not correspond to the actual density perturbations.
Bardeen’s paper has inspired a number of works, including those in Refs. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These researchers proposed alternative perturbation theories using gauge-invariant quantities that differ from those used by Bardeen.
The authors of Refs. [5, 6, 7, 9, 11] define gauge-invariant density perturbations by gradients of the energy density. For a pressureless fluid, their equations are identical to the standard perturbation equation derived from Newton’s theory of gravitation. This equation yields the well-known solutions of the standard equation (150) with $w=0$ and $\beta=0$, including the gauge mode. In addition to the fact that the solution is gauge-dependent, even in the nonrelativistic limit, the evolution of density perturbations in a pressureless fluid is precluded by the absence of pressure gradients. This is a consequence of the linearized momentum conservation laws. It can thus be concluded that the standard equation (150) with $w=0$ and $\beta=0$ cannot be taken to represent the nonrelativistic limit of a cosmological perturbation theory. Moreover, as indicated in Ref. [6], the constraint equations have not been integrated into the analysis. It is of crucial importance to acknowledge that the constraint equations represent relations between initial values, thereby playing a pivotal role in the analysis. In the absence of the constraint equations, the solutions to the system of dynamical equations and conservation laws lack physical significance. Furthermore, it is not possible to achieve a correct Newtonian limit without the constraint equations (for further details, see Sec. IX). In this paper, the complete set of Einstein equations and conservation laws is employed to develop a novel perturbation theory.
In order to interpret the gauge-invariant definition of the density perturbation $\overline{\delta\varepsilon}$, as proposed by Mukhanov et al. [8, 10], it is necessary to set the Hubble function $\cal{H}$ to zero in Eq. (7.47) in Ref. [10]. Nevertheless, this yields a Poisson-like equation that differs from the Newtonian Poisson equation, as the potential is time-dependent in the Poisson-like equation, as evidenced by equations (7.48) and (7.49) in Ref. [10]. Thus, $\overline{\delta\varepsilon}$ does not represent the actual physical density perturbation. Furthermore, the Hubble function cannot be set to zero, as this would violate the Friedmann equation (10a), which is inconsistent with the assumption that the universe is not empty. The present paper introduces a novel perturbation theory that achieves the time-independent Poisson equation and thus the correct nonrelativistic limit by considering zero pressure alone. This obviates the need to set the Hubble function to zero.
As has been stated in Refs. [12, 13], the linearized Einstein equations and conservation laws for density perturbations yield the usual evolution equation (150) derived from Newtonian gravitation when the pressure is zero. The current study will demonstrate that in a pressureless fluid, the linearized Einstein equations and conservation laws are reduced to a single equation: the time-independent Poisson equation of Newtonian gravitation.
The authors of Refs. [14, 15] present a method for the construction of gauge-invariant quantities in the synchronous reference system. However, the definition of gauge invariance presented in these works does not align with the conventional understanding of what constitutes a gauge-invariant quantity. As defined in Refs. [4, 3], a gauge-invariant quantity is independent of the choice of a coordinate system. Consequently, the definition of gauge-invariant quantities is formulated in this paper without the utilization of a coordinate system.
In their study of the evolution of density perturbations in a flat flrw universe in its radiation-dominated phase, Ma & Bertschinger [16] employ two distinct coordinate systems: the synchronous ($g_{0i}=0$) gauge and the conformal Newtonian gauge. The latter gauge is, in fact, a restricted synchronous gauge. Given the existence of gauge transformations between different synchronous reference systems, the solutions are dependent on the selected synchronous coordinates. Consequently, gauge modes emerge in the solutions, rendering them devoid of physical significance even when physical initial values are imposed. Their findings pertaining to the density fluctuation, represented by $\delta$, deviate between the synchronous and conformal Newtonian coordinates. In addition, Eq. (29) in Ref. [16] is flawed in two respects. Firstly, the factor $(1-3w)$ in the expansion term should, in fact, be $(2-3w)$, which is critical in the context of a radiation-dominated universe, where $w=\tfrac{1}{3}$. Secondly, it is crucial to replace the expression $k^{2}$ with $k^{2}/a^{2}$. This substitution is necessary because the background three-space metric of a flat flrw universe is $a^{2}\delta_{ij}$ rather than $\delta_{ij}$. This can be demonstrated by considering the line element (1) in Ref. [16], which illustrates that three-space is expanding. Once these corrections are made, the resulting equation is equal to the covariant divergence of Eq. (43) in this paper. The latter equation is correct, as demonstrated in Appendices B and C.
The aforementioned treatments resulted in a plethora of additional contributions to this field. See, e.g., Refs. [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In a review, Ellis [28] discusses the work of Lifshitz and Khalatnikov and other research on the subject. A review of the literature reveals a lack of consensus concerning the accuracy of the quantities employed to represent physical density perturbations.
This paper addresses the issue of gauge dependence in the context of cosmology. It demonstrates the existence of unique gauge-invariant quantities that correspond to the true physical density perturbations. Subsequently, it presents equations for the study of the evolution of density perturbations in universes with an open, flat, or closed flrw geometry. In light of the general covariance of general relativity, the evolution equations for gauge-invariant quantities are mathematically identical in all coordinate systems. Consequently, they yield the same solutions in any given coordinate system.
The novel perturbation theory, when applied to a flat flrw universe, reveals that prior to the decoupling of radiation and matter, Cold Dark Matter (cdm) perturbations were coupled to the radiation energy density perturbations in such a way that they could not clump together. Consequently, cdm proved to be ineffective in facilitating the process of structure formation subsequent to the decoupling of radiation and matter. The formation of the first stars, designated as Population iii stars [29], can be explained by the rapid growth of small density perturbations over a brief period of time. This rapid growth arises from local nonadiabatic random pressure fluctuations in the early universe following the decoupling. The James Webb Space Telescope (jwst), which was launched on December 25, 2021, is designed primarily for near-infrared astronomy and may therefore provide insights regarding the existence and nature of the earliest stars. The data from the jwst [30] suggest that the galaxies observed in the early universe appear to have grown to a size greater than anticipated, forming massive and compact structures at a rate exceeding the expectations of the hierarchical $\Lambda\textsc{cdm}$ structure formation paradigm. This may indicate that the existence of cdm is not supported by the data.
II A Novel Cosmological Perturbation Theory
This paper presents a novel cosmological perturbation theory, one that differs substantially from the perturbation theories in the extant literature, including the treatment of the cosmic fluid. The aim of this section is to provide an overview of the novel approach and to elucidate the method by which the new theory is constructed.
II.1 Cosmic Fluid
It is well established that in the era following the decoupling of matter and radiation, the pressure in the cosmological fluid is negligible in comparison to the energy density. Therefore, it can be concluded that the pressure can be disregarded in the Einstein equations and conservation laws of the background, i.e., the unperturbed, flrw universe. In the context of cosmology, it is commonly assumed that in the perturbed universe, the pressure and, therefore, local pressure gradients can also be neglected. However, it is well known that in the absence of local pressure gradients, local fluid flows cannot occur. This is equally true in the case of the cosmic fluid. As will be demonstrated in Sec. IX, the linearized momentum conservation laws indicate that, in the absence of pressure, local fluid flows will not occur and density perturbations will be static. It is thus crucial to consider pressure gradients in order to examine the evolution of density perturbations. Given that, subsequent to decoupling, the cosmic fluid is to be considered a perfect gas where the pressure is dependent on the mean kinetic energy of its particles, it is essential that the kinetic energy density be incorporated into the analysis. Both the pressure and the kinetic energy density are dependent on the particle number density and temperature. Consequently, in addition to the energy density, it is necessary to include the particle number density and the temperature in the equation of state. See Ref. [31], Sec. 2.10 on relativistic hydrodynamics, and Ref. [32]. As indicated in Ref. [33], the formation of structures in the universe requires a focus on the clumping of a set of particles rather than on the evolution of energy density instabilities. The approach presented in this paper yields novel insights into the evolution of density perturbations in both the pre-decoupling and post-decoupling eras. This is achieved by examining the evolution of energy density and particle number density perturbations, as well as their interactions. A detailed explanation of this can be found in Sec. XII.
II.2 Perturbation Theory
The equations governing the evolution of the unperturbed universe include three scalars: the energy density, the particle number density, and the expansion. As will be demonstrated in Sec. III, the perturbations to these three scalars transform under general linear coordinate transformations in the same way. This ultimately provides a solution to the gauge problem. As will be demonstrated in Sec. IV, there are only three sets of three gauge-invariant quantities that can be constructed from these scalars and their gauge-dependent perturbations. Each of these sets contains precisely one gauge-invariant quantity that is equal to zero. It will be shown that if the gauge-invariant perturbation of the expansion is zero, then it follows from the nonrelativistic limit in Sec. IX that the other two gauge-invariant quantities, denoted by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, are the true physical perturbations of the energy density and the particle number density, respectively.
The subsequent step is to derive the evolution equations for $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$. In order to do this, it is necessary to employ a suitable coordinate system. In the context of general relativity, there is no a priori preferred coordinate system. Furthermore, the perturbations of the energy and particle number densities are both gauge-invariant, allowing for the freedom to choose any coordinate system. In order to facilitate the derivation of the aforementioned equations, it is typical to select a coordinate system that aligns with the specific characteristics of the problem in question. In Section Sec. V, a detailed justification is provided for the selection of synchronous coordinates. In Sec. VI, the Einstein equations, conservation laws, and their linearized counterparts with respect to this coordinate system are presented.
In Sec. VII, the synchronous reference system is employed to decompose the coupled system of linearized Einstein equations and conservation laws into three independent systems. These systems describe the evolution of tensor, vector, and scalar perturbations, respectively. In the extant literature, it is commonly assumed that the decomposition of the linearized equations can be achieved by decomposing only the perturbation of the metric tensor into its tensor, vector, and scalar parts. However, this is insufficient. Additionally, the perturbation of the spatial Ricci tensor, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}$, must similarly be decomposed into its tensor, vector, and scalar parts, as demonstrated in Ref. [34].
In the case of tensor and vector perturbations, the perturbation of the spatial Ricci scalar, or the spatial curvature perturbation, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, is identically zero. In contrast, for scalar perturbations, the value of ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ is not zero. In addition, it will be demonstrated that the Newtonian potential is encapsulated within the expression for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$. This implies that only scalar perturbations are associated with density perturbations, specifically the physical density perturbations, represented by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
In Sec. VIII, the system of equations (37) is derived which describes the evolution of scalar perturbations. This system incorporates not only the usual conservation laws for energy density and momentum, but also a conservation law for the particle number density. Furthermore, it is found that this system comprises an algebraic energy density constraint equation (37a) that incorporates the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, in addition to an evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$. The latter equation is the covariant divergence of the three linearized momentum constraint equations (12b). In the nonrelativistic limit (i.e., the limit in which the pressure vanishes), the algebraic energy density constraint equation (37a) and the evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, when combined, yield the time-independent Poisson equation of Newtonian gravitation with source term $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x})$, as detailed in Sec. IX. Consequently, the constraint equations (37a)–(37b) are of great importance in the evolution of density perturbations and in the derivation of the nonrelativistic limit. The constraint equations presented in this specific formulation are notably absent from the extant literature on cosmological perturbation theories. This is the precise reason why an exact and correct nonrelativistic limit for a cosmological perturbation theory has not yet been discovered. The absence of a clearly defined nonrelativistic limit makes it impossible to interpret gauge-invariant quantities in previous perturbation theories. The issue is further discussed in Sec. IX, which addresses the nonrelativistic limit.
In Sec. X, the primary result is presented, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/n_{{\scriptscriptstyle(0)}}$ in closed, flat, and open flrw universes. As the quantities $\delta_{\varepsilon}$ and $\delta_{n}$ are gauge-invariant, the evolution equations will have the same mathematical form in all coordinate systems and yield solutions that are independent of the choice of a coordinate system.
In Sec. XII, the recently developed perturbation theory is applied to a flat flrw universe in its phases preceding and following the decoupling of radiation and matter. In order to proceed with the analysis, it is necessary to derive gauge-invariant thermodynamic quantities. These quantities will be presented in Sec. XI. Finally, in Sec. XIII, a summary of the results is provided.
The paper is accompanied by three appendices. In Appendix A, the primary result, Eqs. (60), is derived using the background equations (10), and the equations for scalar perturbations (37). The latter system represents a novel result and is indispensable for the proposed methodology. A proof of its correctness is provided in Appendices B and C. In Appendix C, the systems of equations (10) and (37) are employed once more. In this instance, however, the evolution equations (154) and (156) are recovered for the gauge-dependent quantity defined by $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, in a flat flrw universe. From these equations, the well-known physical solutions for large-scale perturbations in a flat flrw universe are derived. These include the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe following the decoupling of matter and radiation. A comparison of Eqs. (85) and (155a) and Eqs. (109) and (157a) reveals that the novel perturbation theory encompasses all previously established and well-known physical solutions for large-scale density perturbations. This constitutes a significant result, as the aforementioned solutions are widely documented in the literature. It follows that a comprehensive and accurate perturbation theory must also yield these solutions.
The homogeneous part of Eqs. (154a) and (156a) constitutes the conventional relativistic equation (150). Evidently, the latter equation is incomplete and has the gauge mode $\propto t^{-1}$ as solution. As will be demonstrated in Sec. IX, in the nonrelativistic limit, relativistic gauge transformations converge to Newtonian gauge transformations. Consequently, the quantity $\delta$ is also dependent on the choice of coordinates in Newtonian gravitation. Accordingly, it possesses no physical meaning. It can be concluded that the conventional relativistic evolution equation (150), is inadequate for examining the evolution of density fluctuations in the universe.
In the extant literature, the relativistic equation (150) is derived from Newtonian gravitation under two conditions: first, the expansion of the universe must be taken into account, and second, the equation of state must be nonrelativistic. The presence of the gauge mode in the solution of the aforementioned equation, when derived from the Newtonian theory of gravity, is the reason why Newtonian gravitation is not applicable to the study of density fluctuations in an evolving universe.
III Gauge Problem of Cosmology
Einstein’s equations and conservation laws are invariant under general coordinate transformations $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. This signifies that there are no preferred coordinate systems. Consequently, the linearized Einstein equations and conservation laws are invariant under a general linear infinitesimal space-time transformation, which is usually referred to as a gauge transformation. This transformation is given by
$x^{\mu}\rightarrow x^{\prime\mu}=x^{\mu}-\xi^{\mu}(t,\bm{x}),$ (1)
where the gauge functions $\xi^{\mu}(t,\bm{x})$ are four arbitrary infinitesimal functions of time, $x^{0}=ct$ with $c$ the speed of light, and space, $\bm{x}=(x^{1},x^{2},x^{3})$, coordinates.
Consider a closed, flat, or open flrw universe. The evolution equations for these universes contain three scalars: the energy density $\varepsilon$, the particle number density $n$, and the expansion $\theta$. These scalars are defined as follows
$\varepsilon:=T^{\mu\nu}u_{\mu}u_{\nu},\quad n:=N^{\mu}u_{\mu},\quad\theta:=u^{\mu}{}_{;\mu},$ (2)
where $u^{\mu}$ is the fluid four-velocity normalized to unity, ${u^{\mu}u_{\mu}=1}$, and $N^{\mu}:=nu^{\mu}$ represents the cosmological particle current four-vector, which satisfies the particle number conservation law ${N^{\mu}{}_{;\mu}=0}$. A semicolon denotes covariant differentiation.
Let $S_{\scriptscriptstyle(0)}(t)$ denote the quantities $\varepsilon_{\scriptscriptstyle(0)}$, $n_{\scriptscriptstyle(0)}$, and $\theta_{\scriptscriptstyle(0)}$. These quantities satisfy the background, that is to say, the unperturbed Einstein equations and conservation laws. Let $S_{\scriptscriptstyle(1)}(t,\bm{x})$ be their perturbed counterparts $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, which satisfy the linearized equations. As a consequence of the linearity of these equations, new solutions that are physically equivalent can be generated. These solutions are given by (see Weinberg [31], Sec. 10.9 for a detailed explanation)
$S^{\prime}_{\scriptscriptstyle(1)}=S_{\scriptscriptstyle(1)}+\mathcal{L}_{\xi}S_{\scriptscriptstyle(0)}=S_{\scriptscriptstyle(1)}+\xi^{0}\dot{S}_{\scriptscriptstyle(0)}.$ (3)
In this expression, the operator $\mathcal{L}_{\xi}$ is the Lie derivative with respect to the infinitesimal four-vector $\xi^{\mu}$. An overdot denotes differentiation with respect to $x^{0}=ct$. The solutions $S_{\scriptscriptstyle(1)}$ and $S^{\prime}_{\scriptscriptstyle(1)}$ contain the so-called gauge modes which are given by $\hat{S}_{\scriptscriptstyle(1)}=\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This indicates that the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are dependent on the choice of a system of reference, rendering them gauge-dependent and, consequently, devoid of physical significance. This is the gauge problem of cosmology, whereby the coordinate artifacts and the underlying physics are inextricably linked in the solution of the linearized equations. Even when physical initial values are imposed in order to solve the linearized equations, the resulting solution lacks physical meaning. It is therefore imperative to resolve the gauge problem in order to facilitate an investigation into the evolution of density perturbations in the universe.
IV Solving the Gauge Problem of Cosmology
In contrast to the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, their physical counterparts are independent of the choice of a coordinate system. This is to say that they are gauge-invariant. As the physics of the density perturbations is concealed within the general solution of the linearized Einstein equations and conservation laws, it follows that the physical counterparts of $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, must be linear combinations of gauge-dependent solutions to these equations. In order to identify these quantities, it is necessary to consider the fact that a quantity is gauge-invariant if it is independent of the coordinate transformation parameter $\xi^{\mu}$, as outlined by Stewart [3]. This implies that the quantity must be invariant under a linear coordinate transformation (1). As is evident from (3), the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are independent of the spatial components, $\xi^{i}$, of the transformation, but not of the temporal component, $\xi^{0}$. Given that the transformation of these quantities under the general infinitesimal transformation (1) is identical to that described by (3), it is possible to construct gauge-invariant quantities by combining these gauge-dependent quantities in a way that eliminates the gauge modes $\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This results in three distinct sets of linear combinations. In each of these sets, precisely one gauge-invariant quantity is identically zero. Given that the focus is on energy density and particle number density perturbations, the only possible set is
$\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\varepsilon_{\scriptscriptstyle(1)}-\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad n_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=n_{\scriptscriptstyle(1)}-\dfrac{\dot{n}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},$ (4a)
$\theta_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\theta_{\scriptscriptstyle(1)}-\dfrac{\dot{\theta}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}=0.$ (4b)
Local density perturbations exert no influence on the expansion of the universe, as evidenced by (4b). The transformation rule (3), illustrates that the quantities (4) are indeed gauge-invariant.
In Sec. IX, it is demonstrated that in the nonrelativistic limit, where the pressure is considered to be zero, the linearized conservation laws lack physical significance and are decoupled from the linearized energy density constraint equation and momentum constraint equations. The combination of these constraint equations results in the time-independent Poisson equation of Newtonian gravitation with a source term $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})$. Furthermore, the special relativistic relation $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}mc^{2}$, where $m$ is the rest mass, is recovered. In light of the aforementioned considerations, it can be concluded that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the actual physical perturbations of the energy density and the particle number density, respectively, in the context of general relativity.
The sole method for constructing gauge-invariant quantities for the energy density and particle number density perturbations is provided by the expressions (4). As these quantities also possess the correct nonrelativistic limit, it can be concluded that the gauge problem of cosmology has been resolved. This outcome is an inevitable result of the mathematical consistency of general relativity. That is to say, the general theory of relativity allows for only one gauge-invariant quantity that accurately represents a physical density perturbation.
V Selection of a Reference System
The quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ are defined without the use of a coordinate system. However, to derive evolution equations for these quantities, it is first necessary to establish a system of reference. Since general relativity is covariant, and the quantities in question are gauge-invariant, it follows that the evolution equations for these quantities can be derived in any chosen coordinate system. The selection of an appropriate coordinate system for this problem is dependent on two criteria, which are outlined below.
Firstly, the interpretation of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ requires taking the nonrelativistic limit. In the context of general relativity, coordinate transformations are in general space-time transformations, specifically $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. In the linear case, space-time transformations are given by (1). In Newtonian gravitation, the space and time transformations are regarded as distinct and independent entities. Consequently, when taking the nonrelativistic limit, the relativistic space-time transformations must be automatically separated into independent space and time transformations. Secondly, it would be advantageous to have a coordinate system that would facilitate the derivation of the evolution equations. It will now be demonstrated that both of these requirements can be met by employing the same coordinate system.
In Newtonian gravitation, where space and time are treated as distinct and independent entities, all coordinate systems are inherently synchronous. In view of the nonrelativistic limit, it can be seen that a synchronous reference system represents the optimal choice. In these coordinates, the metric tensor $g_{\mu\nu}(t,\bm{x})$ of the flrw universes is given by
$g_{00}=1,\quad g_{0i}=0,\quad g_{ij}=-a^{2}(t)\tilde{g}_{ij}(\bm{x}).$ (5)
The scale factor of the universe is designated as $a(t)$. Given that $g_{00}=1$, coordinate time is equal to proper time. The synchronicity condition is $g_{0i}=0$, as explained in Ref. [31], Sec. 11.8 and Ref. [35], § 84. The metric tensor of the three-dimensional maximally symmetric subspaces of constant time is denoted by $\tilde{g}_{ij}$. The Killing equations, $\xi_{\mu;\nu}+\xi_{\nu;\mu}=0$, and (5) demonstrate that the gauge functions $\xi^{\mu}(t,\bm{x})$ in the transformation (1) become
$\xi^{0}=\psi(\bm{x}),\quad\xi^{i}=\tilde{g}^{ik}(\bm{x})\dfrac{\partial\psi(\bm{x})}{\partial x^{k}}\int\!\!\frac{{\mathrm{d}}t}{a^{2}(t)}+\chi^{i}(\bm{x}),$ (6)
if only transformations between synchronous coordinates are allowed. In expressions (6), $\psi(\bm{x})$ and $\chi^{i}(\bm{x})$ are four arbitrary infinitesimal functions of the spatial coordinates. In Sec. IX, it will be proven that in the nonrelativistic limit, the function $\psi(\bm{x})$ is an arbitrary infinitesimal constant. This means that the transformations of space and time are independent of each other in this limit.
The second requirement will now be considered. Synchronous coordinates possess the property that the space-space components, $R_{ij}$, of the four-dimensional Ricci curvature tensor, $R_{\mu\nu}$, are partitioned into two distinct parts. One part exclusively comprises the time derivatives of the spatial metric tensor, $g_{ij}$, while the second part is the Ricci curvature tensor of the three-dimensional subspaces, denoted by ${{}^{3}\!}R_{ij}$. This is demonstrated in Ref. [35], § 97. As the perturbations of the spatial metric tensor and the spatial Ricci tensor are also tensors (see Sec. VI.2), the covariant decomposition theorems in Refs. [34, 36, 3] may be employed to decompose these perturbed tensors into their scalar, vector, and tensor parts. This decomposition permits the system of linearized Einstein equations and conservation laws to be decomposed into three independent systems, each of which describes the evolution of scalar, vector, and tensor perturbations. As will be demonstrated in Sec. VII, only scalar perturbations are associated with density perturbations. Accordingly, it is sufficient to consider the system for scalar perturbations (37). The aforementioned system is more straightforward to utilize than the original set of equations (12), thus facilitating the derivation of the evolution equations for density fluctuations in synchronous coordinates in comparison to any other coordinate system. In view of the fact that synchronous coordinates are likewise compatible with Newton’s theory of gravitation, they will be employed in the derivation of the primary result (60).
VI Einstein Equations and Conservation Laws
In the preceding section, it was concluded that synchronous coordinates are the optimal choice for addressing the problem at hand. In this section, the background Einstein equations and conservation laws, along with their linearized counterparts, will be expressed in this particular coordinate system.
The background energy density constraint equation and its linearized counterpart are expressed in a manner consistent with the contracted Bianchi identities. These identities state that the Riemann curvature tensor satisfies the equation $R^{\mu}{}_{\nu;\mu}=\tfrac{1}{2}R_{,\nu}$. This can be expressed in an equivalent manner as the four-divergence of the Einstein tensor, i.e., ${G^{\mu\nu}{}_{;\nu}=0}$. This result demonstrates that the constraint equations contain at most first-order time derivatives of the metric, which is precisely what is required for the development of the novel perturbation theory. For further details see Ref. [31], Secs. 6.8 and 7.5 and Ref. [35], § 92 and § 95.
It is understood from thermodynamic principles that the energy density, $\varepsilon$, and the pressure, $p$, are functions of both the particle number density, $n$, and the temperature, $T$:
$\varepsilon=\varepsilon(n,T),\quad p=p(n,T).$ (7)
In order to facilitate the calculations, it is advantageous to eliminate the temperature from the given equations of state. This yields the equation of state for the pressure:
$p=p(\varepsilon,n),$ (8)
which will be utilized in the subsequent analysis.
VI.1 Background Equations
The system of background Einstein equations and conservation laws for closed, flat, and open flrw universes filled with a perfect fluid with an energy-momentum tensor as defined by
$T^{\mu\nu}=(\varepsilon+p)u^{\mu}u^{\nu}-pg^{\mu\nu},\quad p=p(n,\varepsilon),$ (9)
is represented by the following equations:
$3H^{2}$ $=\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}+\kappa\varepsilon_{\scriptscriptstyle(0)}+\Lambda,\quad$ $\kappa$ $=8\pi G_{\text{N}}/c^{4}$ (10a)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(0)}$ $=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(0)},$ (10b)
$\dot{\varepsilon}_{\scriptscriptstyle(0)}$ $=-3H\varepsilon_{\scriptscriptstyle(0)}(1+w),$ $w$ $:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)},$ (10c)
$\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (10d)
It can be demonstrated that the $G_{0i}$ constraint equations and the $G_{ij}$ dynamical equations with $i\neq j$ are satisfied identically. The $G_{ii}$ dynamical equations are equivalent to the time derivative of the $G_{00}$ constraint equation (Friedmann equation) given by equation (10a). Therefore, the $G_{ij}$ dynamical equations can be omitted. The cosmological constant, $\Lambda$, the gravitational constant, $G_{\text{N}}$, and the speed of light, $c$, are fundamental constants.
The overdot is used to denote differentiation with respect to $x^{0}=ct$. The Hubble function $H$ is defined by $H:=\dot{a}/a$. For flrw universes, the Hubble function is given by $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, where $\theta_{\scriptscriptstyle(0)}$ is the background value of the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$, where $u^{\mu}$ is the four-velocity $u^{\mu}:=c^{-1}U^{\mu}$, normalized to unity, $u^{\mu}u_{\mu}=1$. A semicolon is used to denote covariant differentiation with respect to the background metric tensor $g_{{\scriptscriptstyle(0)}\mu\nu}$ (5). The equations (10c) and (10d) represent the energy density conservation law and the particle number conservation law, respectively. It is crucial to highlight that the variable $w$ in Eq. (10c) is an abbreviation for $p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, and it does not represent the equation of state. The spatial parts of the background Riemann tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}$, the Ricci curvature tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}$, and its contraction, ${{}^{3}\!}R_{\scriptscriptstyle(0)}$, are given by the following expressions:
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}=\tilde{R}^{i}{}_{jkl}=K(\delta^{i}{}_{k}\tilde{g}_{jl}-\delta^{i}{}_{l}\tilde{g}_{jk}),$ (11a)
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}=-\dfrac{2K}{a^{2}}\delta^{i}{}_{j},\quad{{}^{3}\!}R_{\scriptscriptstyle(0)}=-\dfrac{6K}{a^{2}},$ (11b)
where ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ is the spatial curvature. The value of $K$ determines the nature of the flrw universe. The universe is open for $K=-1$, flat for $K=0$, and closed for $K=+1$.
VI.2 Linearized Equations
The system of linearized Einstein equations and conservation laws for closed, flat, and open flrw universes is succinctly expressed as follows:
$H\dot{h}^{k}{}_{k}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\kappa\varepsilon_{\scriptscriptstyle(1)},$ (12a)
$\dot{h}^{k}{}_{k\mid i}-\dot{h}^{k}{}_{i\mid k}=2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}i},$ (12b)
$\ddot{h}^{i}{}_{j}+3H\dot{h}^{i}{}_{j}\,+$
$\qquad\delta^{i}{}_{j}H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}=-\kappa\delta^{i}{}_{j}(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}),$ (12c)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})+\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)}=0,$ (12d)
$\frac{1}{c}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}\right)\,-$
$\qquad g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}=0,$ (12e)
$\dot{n}_{\scriptscriptstyle(1)}+3Hn_{\scriptscriptstyle(1)}+n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}=0.$ (12f)
In these equations, the following notation is employed for the perturbed metric tensor: $h_{\mu\nu}:=-g_{{\scriptscriptstyle(1)}\mu\nu}$ and $h^{\mu\nu}:=g^{\mu\nu}_{{\scriptscriptstyle(1)}}$. In light of the use of synchronous coordinates, it follows that $h_{00}=0$ and $h_{0i}=0$. In addition, the raising and lowering of indices is performed by the spatial background metric tensor as defined in expression (5), namely, $h^{i}{}_{j}=g^{ik}_{\scriptscriptstyle(0)}h_{kj}$ where $g^{ik}_{\scriptscriptstyle(0)}=-\tilde{g}^{ik}/a^{2}$. A vertical bar denotes covariant differentiation with respect to the spatial background metric tensor $g_{{\scriptscriptstyle(0)}ij}=-a^{2}\tilde{g}_{ij}$ or, equivalently, with respect to $\tilde{g}_{ij}$. In the context of flrw universes, where $\Gamma^{k}_{{\scriptscriptstyle(0)}ij}=\tilde{\Gamma}^{k}{}_{ij}$, the operations of taking the time derivative and the covariant derivative are shown to commute. Equation (12a) represents the linearized energy density constraint equation (linearized Friedmann equation), while Eqs. (12b) correspond to the linearized momentum constraint equations. The linearized dynamical equations are presented in Eqs. (12c). Equations (12d) and (12e) represent the linearized energy density conservation law and the linearized energy-momentum conservation laws, respectively. Finally, the linearized particle number density conservation law is given by Eq. (12f).
As previously stated in Sec. II, the pressure perturbation, the spatial curvature perturbation, and the covariant divergence of the spatial fluid velocity play a pivotal role in a cosmological perturbation theory. In consequence, the expressions for these quantities are now derived.
VI.2.1 Pressure Perturbation
The gauge-dependent perturbation to the pressure (8) is given by
$p_{{\scriptscriptstyle(1)}}=p_{n}n_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}},$ (13a)
$p_{n}:=\left(\dfrac{\partial p}{\partial n}\right)_{\varepsilon},\quad p_{\varepsilon}:=\left(\dfrac{\partial p}{\partial\varepsilon}\right)_{n}.$ (13b)
The quantity representing the true physical perturbation, designated as $p^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, will be derived in Sec. XI on thermodynamics.
VI.2.2 Spatial Curvature Perturbation
Lifshitz’ expression for the perturbed connection coefficients presented in Ref. [2], Eq. (I.3) and Ref. [31], Eq. (10.9.1) is a tensor, which is given by
$\Gamma^{k}_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j}+h_{lj\mid i}-h_{ij\mid l}).$ (14)
The contracted Palatini identity as defined in Ref. [2], Eq. (I.5) and Ref. [31], Eq. (10.9.2) is given by
${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=\Gamma^{k}_{{\scriptscriptstyle(1)}ij\mid k}-\Gamma^{k}_{{\scriptscriptstyle(1)}ik\mid j}.$ (15)
By combining (14) and (15), the following expression for the perturbed spatial Ricci tensor may be derived:
${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j\mid k}+h_{lj\mid i\mid k}-h_{ij\mid l\mid k}-h_{lk\mid i\mid j}).$ (16)
By raising the index $i$, one arrives at the following result, employing (11b):
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}:=$ $\,(g^{ik}\,{{}^{3}\!}R_{kj})_{\scriptscriptstyle(1)}=g^{ik}_{\scriptscriptstyle(0)}\,{{}^{3}\!}R_{{\scriptscriptstyle(1)}kj}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}$
$=$ $-\tfrac{1}{2}g^{il}_{\scriptscriptstyle(0)}(h^{k}{}_{l\mid j\mid k}+h^{k}{}_{j\mid l\mid k}-h^{k}{}_{k\mid l\mid j})$
$+\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}h^{i}{}_{j\mid k\mid l}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}.$ (17)
By employing $g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid j\mid k}=g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid k\mid j}$, one arrives at the contraction of the perturbed spatial Ricci tensor:
${{}^{3}\!}R_{\scriptscriptstyle(1)}:={{}^{3}\!}R^{k}_{{\scriptscriptstyle(1)}k}=g_{\scriptscriptstyle(0)}^{ij}(h^{k}{}_{k\mid i\mid j}-h^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{k}{}_{k}.$ (18)
In Sec. VII it will be demonstrated that the expression (18) represents the local perturbation to the spatial curvature ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}$, or the spatial curvature perturbation, induced by local density perturbations. The spatial curvature perturbation ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ is incorporated into the evolution equations (60), as shown in Eq. (128).
VI.2.3 Covariant Divergence of the Spatial Fluid Velocity
In the background flrw universe, the fluid four-velocity is given by $u^{\mu}_{\scriptscriptstyle(0)}=\delta^{\mu}{}_{0}$, so that the perturbation $\theta_{\scriptscriptstyle(1)}$ to the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$ is
$\theta_{\scriptscriptstyle(1)}=\vartheta_{\scriptscriptstyle(1)}-\tfrac{1}{2}\dot{h}^{k}{}_{k},\quad\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k},$ (19)
where $\vartheta_{\scriptscriptstyle(1)}$ is the covariant divergence of the spatial part $u^{i}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity $u^{\mu}_{\scriptscriptstyle(1)}$. In deriving (19), the following identity is employed: $(u^{k}{}_{|k})_{{\scriptscriptstyle(1)}}=u^{k}_{{\scriptscriptstyle(1)}|k}$, which follows from $u^{i}_{{\scriptscriptstyle(0)}}=0$. The covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the evolution equations (60), as shown in Eq. (133).
VII Decomposition of Spatial Tensors
The following important theorem is used to decompose the linearized Einstein equations and conservation laws into three independent systems. In Refs. [34, 36, 3], it is demonstrated that any symmetric spatial tensor ${\cal{T}}_{ij}$ of rank two can be unambiguously decomposed in a covariant manner into three irreducible components:
${\cal{T}}^{i}{}_{j}={\cal{T}}^{i}_{\parallel\;j}+{\cal{T}}^{i}_{\perp\;j}+{\cal{T}}^{i}_{\ast\;j}.$ (20)
The constituents possess the following properties:
${\cal{T}}^{k}_{\perp\;\,k}=0,\quad{\cal{T}}^{k}_{\ast\;\;k}=0,\quad{\cal{T}}^{k}_{\ast\;\;i\mid k}=0,\quad{\cal{T}}^{i}_{\parallel\;j}=\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j},$ (21)
where $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$ are two independent potentials. Consequently, the component $h^{i}_{\parallel\,j}$ can be expressed in terms of the two independent potentials as follows:
$h^{i}_{\parallel\,j}=\frac{2}{c^{2}}\bigl{(}\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j}\bigr{)}.$ (22)
The factor $2/c^{2}$ is included in view of the nonrelativistic limit. Evidently, from (21), it can be inferred that an expression analogous to (22) exists for the tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\parallel\,j}$. In the context of cosmological perturbation theory, only the contraction ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ of the latter tensor is required. Upon substituting (22) into (18), the following result is obtained:
${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}=\dfrac{2}{c^{2}}\Bigl{[}2\phi^{|i}{}_{|i}$ $+\zeta^{|k}{}_{|k}{}^{|i}{}_{|i}-\zeta^{|k|i}{}_{|k|i}$
$+\tfrac{1}{3}\,{{}^{3}\!}R_{{\scriptscriptstyle(0)}}(3\phi+\zeta^{|k}{}_{|k})\Bigr{]},$ (23)
and for the perturbed expansion (19) it is found
$\theta_{{\scriptscriptstyle(1)}}=\vartheta_{{\scriptscriptstyle(1)}}-\dfrac{1}{c^{2}}\bigl{(}3\dot{\phi}+\dot{\zeta}^{|k}{}_{|k}\bigr{)}.$ (24)
As evidenced by (VII), the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ encompasses two distinct potentials $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$. In the nonrelativistic limit, as will be discussed in detail in Sec. IX, the potential $\phi$ becomes independent of time and equal to the Newtonian potential, while the potential $\zeta$ becomes inconsequential.
The spatial component $\bm{u}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity can be decomposed as follows
$\bm{u}_{{\scriptscriptstyle(1)}}=\bm{u}_{{\scriptscriptstyle(1)}\parallel}+\bm{u}_{{\scriptscriptstyle(1)}\perp},$ (25)
where the components $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ and $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ have the properties
~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel},\quad\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{{\scriptscriptstyle(1)}\perp}.$ (26)
In these expressions, the generalized vector differential operator, denoted by $\bm{\tilde{\nabla}}$, is defined by the relation ~${\tilde{\nabla}_{i}v_{k}:=v_{k\mid i}}$. The expressions (25)–(26) are referred to as the Helmholtz decomposition. The following section will demonstrate that the decomposition of ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}$ and $h^{i}{}_{j}$ as given by (20) and the properties (21) are consistent with the the Riemann tensor (11a) and the momentum constraint equations (12b).
It is important to note that the velocity vector $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is not uniquely determined by the system of equations (12). Since $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is irrotational, it is possible to supplement it with the gradient of an arbitrary function. Given that the system of equations (12) is invariant under the gauge transformation (1), where the gauge function $\xi^{\mu}(t,\bm{x})$ is given by (6), it follows that
$u^{\prime}_{{\scriptscriptstyle(1)}\parallel i}=u_{{\scriptscriptstyle(1)}\parallel i}+\psi_{|i},$ (27)
is also a solution of the system (12). As a result, the gauge modes associated with $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ are given by
$\hat{u}_{{\scriptscriptstyle(1)}\parallel i}=\psi_{|i},\quad\hat{u}^{i}_{{\scriptscriptstyle(1)}\parallel}=g_{{\scriptscriptstyle(0)}}^{ik}\psi_{|k}=-\dfrac{1}{a^{2}}\tilde{g}^{ik}\psi_{|k}.$ (28)
This result is of particular significance in the context of the derivation of the nonrelativistic limit, as presented in Sec. IX.
VII.1 Decomposition of the Linearized Equations
The system of equations (12) is split into three independent systems by using the decomposition (20) with the properties (21) for $h_{ij}$ and ${{}^{3}\!}R_{ij}$, as well as (25)–(26). The solutions to these systems are conventionally designated as tensor perturbations $\ast$, vector perturbations $\perp$, and scalar perturbations $\parallel$.
In the following three subsections, it is demonstrated that scalar perturbations are the only ones associated with density perturbations $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
VII.1.1 Tensor Perturbations
From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\ast}=0$. This, in combination with Eq. (12a), yields ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$. Eqs. (12b) yield the result that ${\bm{u}_{\scriptscriptstyle(1)}=\bm{0}}$. This, in turn, implies, with (19) that $\theta_{{\scriptscriptstyle(1)}}=0$. From (4a), it follows that ${\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0}$. Given that $\theta_{{\scriptscriptstyle(1)}}=0$, Eqs. (10d) and (12f) are identical. This implies that $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$. In consequence, the evolution equations for tensor perturbations are as follows:
$\ddot{h}^{i}_{\ast\,j}+3H\dot{h}^{i}_{\ast\,j}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\ast j}=0.$ (29)
Given the form of these equations, tensor perturbations are typically referred to as gravitational waves.
VII.1.2 Vector Perturbations
From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=0$. This implies that
$h^{kl}_{\perp\mid k\mid l}=0.$ (30)
Moreover, it is found from Eq. (12a), that ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$. Raising the index $i$ of Eqs. (12b) with $g^{ij}_{\scriptscriptstyle(0)}$, and subsequently taking the covariant derivative with respect to the index $j$, one finds, using that $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$,
$\dot{h}^{kj}_{\perp{\mid k\mid j}}+2Hh^{kj}_{\perp{\mid k\mid j}}=-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{j}_{{\scriptscriptstyle(1)}\mid j}.$ (31)
The combination of (30) and (31) results in ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=0$. This indicates that only the component $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ remains. From (19) one finds that $\theta_{{\scriptscriptstyle(1)}}=0$. This, combined with $\varepsilon_{{\scriptscriptstyle(1)}}=0$, yields $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. From $\theta_{{\scriptscriptstyle(1)}}=0$ one finds that Eq. (12f) is identical to the background equation (10d), implying that $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies with (4a) that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$. Therefore, the system of equations for vector perturbations is given by
$\!\dot{h}^{k}_{\perp i\mid k}+2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}\perp i}=0,$ (32a)
$\!\ddot{h}^{i}_{\perp j}+3H\dot{h}^{i}_{\perp j}+2R^{i}_{{\scriptscriptstyle(1)}\perp j}=0,$ (32b)
$\!\frac{1}{c}\frac{{\text{d}}}{{\text{d}}t}\Bigl{(}\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}\Bigr{)}+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}=0.$ (32c)
According to (26), vector perturbations are also referred to as rotational perturbations.
VII.1.3 Scalar Perturbations
In this particular instance one has ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}\neq 0$. By taking the covariant derivative of (12b) with respect to the index $j$ and subsequently substituting the expression (22), the following result is obtained:
$2\dot{\phi}_{\mid i\mid j}+\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}=\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}i\mid j}.$ (33)
Upon interchanging the indices $i$ and $j$ and subtracting the result from Eqs. (33) one finds
$\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}=-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}),$ (34)
where it is used that $\dot{\phi}_{\mid i\mid j}=\dot{\phi}_{\mid j\mid i}$ and $\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}=\dot{\zeta}^{\mid k}{}_{\mid k\mid j\mid i}$. By rearranging the covariant derivatives, Eqs. (34) can be transformed in the following form:
$(\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid j\mid k})-(\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i\mid k})\,+$
$\qquad(\dot{\zeta}^{\mid k}{}_{\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i})_{\mid k}\,=$
$\qquad-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}).$ (35)
The expressions for the commutator of second-order covariant derivatives (see Ref. [31], Sec. 6.5) are given by
$A^{i}{}_{j\mid p\mid q}-A^{i}{}_{j\mid q\mid p}=A^{i}{}_{k}\,{{}^{3}\!}R^{k}_{{\scriptscriptstyle(0)}jpq}-A^{k}{}_{j}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq},$ (36a)
$B^{i}{}_{\mid p\mid q}-B^{i}{}_{\mid q\mid p}=B^{k}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq}.$ (36b)
Upon substituting the background Riemann tensor (11a), it is found that the left-hand sides of Eqs. (35) vanished identically. This implies that ~×$\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$, so that only the component $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ remains. Since ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}\neq 0$ implies that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$, it can thus be concluded that only scalar perturbations are coupled to density perturbations.
VIII Evolution Equations for Scalar Perturbations
As demonstrated in the preceding section, the evolution of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is determined by the background equations (10) and the equations that govern the evolution of scalar perturbations. In this section, the findings of the preceding section are utilized to reformulate the system (12) into a novel set of equations that exclusively describe the evolution of scalar perturbations. Given the subsequent focus on scalar perturbations, the subscript $\parallel$ is omitted. The evolution equations for scalar perturbations are as follows:
$2H(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})=\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+\kappa\varepsilon_{\scriptscriptstyle(1)},$ (37a)
$\!\!{{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$
$\qquad\;\,+2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}-\tfrac{2}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)}),$ (37b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}=-3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})-\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)},$ (37c)
$\dot{\vartheta}_{\scriptscriptstyle(1)}=-H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}-\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}p_{\scriptscriptstyle(1)}}{a^{2}},$ (37d)
$\dot{n}_{\scriptscriptstyle(1)}=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (37e)
The symbol ~$\tilde{\nabla}^{2}$ denotes the Laplace-Beltrami operator with respect to the metric $\tilde{g}_{ij}(\bm{x})$ of three-dimensional subspaces of constant time:
~$\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}:=\tilde{g}^{ij}p_{{\scriptscriptstyle(1)}\mid i\mid j},\quad g^{ij}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|i|j}=-\dfrac{\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}}{a^{2}}.$ (38)
The parameter $\beta$ is defined as follows:
$\beta^{2}:=\dfrac{\dot{p}_{{\scriptscriptstyle(0)}}}{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}.$ (39)
By employing the expression $\dot{p}_{\scriptscriptstyle(0)}=p_{n}\dot{n}_{\scriptscriptstyle(0)}+p_{\varepsilon}\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and eliminating the time derivatives of $\varepsilon_{\scriptscriptstyle(0)}$ and $n_{\scriptscriptstyle(0)}$ with the aid of the conservation laws (10c) and (10d), the following result is obtained:
$\beta^{2}=p_{\varepsilon}+\dfrac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (40)
where $p_{\varepsilon}$ and $p_{n}$ are given by  (13b).
The derivation of Eqs. (37) will now be conducted. To initiate this process, the term $\dot{h}^{k}{}_{k}$ is eliminated from Eq. (12a) using the expression (19), which results in the algebraic energy density constraint equation (37a).
Given that only $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is associated with scalar perturbations, it is possible to simplify the momentum constraint equations (12b) and the momentum conservation laws (12e) by employing the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}:=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}$, (19), as an alternative to $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$. By multiplying both sides of Eqs. (12b) by $g^{ij}_{\scriptscriptstyle(0)}$ and taking the covariant derivative with respect to the index $j$, one finds
$g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})=2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}.$ (41)
The left-hand side of Eq. (41) will appear as a component of the time derivative of the local spatial curvature perturbation, ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Indeed, differentiating (18) with respect to time, and using the relation $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$ and (10b), yields the following result:
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\dot{h}^{k}{}_{k}.$ (42)
By combining (41) and (42) and using (19) to eliminate $\dot{h}^{k}{}_{k}$, one obtains Eq. (37b). In consequence, the three $G^{0}_{{\scriptscriptstyle(1)}i}$ momentum constraint equations (12b) have been reformulated in the form of a first-order ordinary differential equation (37b) for the perturbed spatial Ricci scalar (18).
The momentum conservation laws (12e) are rewritten by performing the differentiation with respect to time and using (10c), then dividing by $\varepsilon_{{\scriptscriptstyle(0)}}(1+w)$. The result is as follows:
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}+H(2-3w)u^{i}_{{\scriptscriptstyle(1)}}+\dfrac{\dot{w}}{1+w}u^{i}_{{\scriptscriptstyle(1)}}-\dfrac{g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}}{\varepsilon_{{\scriptscriptstyle(0)}}(1+w)}=0.$ (43)
This equation can be simplified by eliminating the time derivative of $w$. From the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and the energy conservation law (10c) one obtains
$\dot{w}=3H(1+w)(w-\beta^{2}).$ (44)
Upon substituting this expression into Eq. (43), it can be seen that the momentum conservation laws can be reformulated as
$\dot{u}^{i}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})u^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}=0.$ (45)
By taking the covariant divergence of (45) with respect to the background metric tensor (5), and employing (19) and (38), one obtains Eq. (37d).
As demonstrated in Appendix B, the dynamical equations (12c) are superfluous since the system (37) comprises the constraint equations and conservation laws, and thus its solution automatically satisfies the dynamical equations. Consequently, the derivation of the equations (37) has been successfully completed.
In Appendix A, the primary result, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}/n_{{\scriptscriptstyle(0)}}$, will be derived using the systems (10) and (37). Prior to proceeding, it is crucial to demonstrate that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the true physical energy density and particle number density perturbations, respectively. This necessitates the nonrelativistic limit being employed for the interpretation of these quantities.
IX Nonrelativistic Limit
The standard nonrelativistic limit is defined by four fundamental criteria (see, e.g., Ref. [37], Sec. 4.1):
  1. 1.
    The subspaces of constant time are flat.
  2. 2.
    When the pressure is reduced to zero, there are no local fluid flows.
  3. 3.
    Relativistic gauge transformations are reduced to Newtonian gauge transformations.
  4. 4.
    The gravitational field of a density perturbation is required to be constant.
The third requirement is of considerable importance, as it permits the selection of a coordinate system in Newtonian gravitation. This fact is not referenced in the existing literature. The fourth requirement is a direct consequence of the first and second requirement.
In the case of flat three-spaces, it is evident that ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$. In a flat flrw universe, covariant derivatives are equivalent to ordinary derivatives, thereby simplifying the expression (VII) to the following form:
${{}^{3}\!}R_{{\scriptscriptstyle(1)}}=\dfrac{4}{c^{2}}\phi^{\mid k}{}_{\mid k}=-\dfrac{4}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}.$ (46)
The symbol $\nabla^{2}$ represents the conventional Laplace operator. Upon substituting the expression (46) into the perturbation equations (37), and employing the relation $H:=\dot{a}/a$, the following result is obtained:
$-H\vartheta_{\scriptscriptstyle(1)}$ $=-\dfrac{1}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)},\quad\kappa=8\pi G_{\text{N}}/c^{4},$ (47a)
$\dfrac{1}{c^{2}}\dfrac{\nabla^{2}\dot{\phi}}{a^{2}}$ $=-\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)},\quad w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}},$ (47b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $=-3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})-\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)},$ (47c)
$\dot{\vartheta}_{\scriptscriptstyle(1)}$ $=-H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}-\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\nabla^{2}p_{\scriptscriptstyle(1)}}{a^{2}},$ (47d)
$\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)},\quad\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (47e)
In order to derive the $G^{0}_{{\scriptscriptstyle(1)}0}$ constraint equation (47a) from Eq. (37a), it is first necessary to eliminate the function $\varepsilon_{{\scriptscriptstyle(1)}}$ using the expression (4a). The result is
$2H(\theta_{{\scriptscriptstyle(1)}}-\vartheta_{{\scriptscriptstyle(1)}})=-\dfrac{2}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\kappa\left(\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\dfrac{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}{\dot{\theta}_{{\scriptscriptstyle(0)}}}\theta_{{\scriptscriptstyle(1)}}\right).$ (48)
Subsequently, the Friedmann equation (10a) with the Ricci scalar ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ set to zero, the energy conservation law (10c), and the relation between the Hubble parameter and the expansion scalar, namely, $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, are employed to obtain the result $\dot{\varepsilon}_{\scriptscriptstyle(0)}/\dot{\theta}_{\scriptscriptstyle(0)}=2H/\kappa$. Upon substituting this expression into (48), the desired result, Eq. (47a), is obtained. With the condition of ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, the first requirement is satisfied.
Subsequently, the second requirement will be implemented. Upon substituting $p=0$, i.e., $w=0$ and $\beta=0$, into the momentum conservation laws (45), the result is
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}=-2Hu^{i}_{{\scriptscriptstyle(1)}}.$ (49)
By lowering the index $i$ with the background metric $g_{{\scriptscriptstyle(0)}ij}$ given by (5), and using the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$, one finds
$\dot{u}_{{\scriptscriptstyle(1)}i}=0.$ (50)
These equations admit only functions of the spatial coordinates as solutions. As these equations are part of the system (12), it can be concluded that the gauge modes, $\psi_{|i}$, given by (28), are the only solutions of equations (50). Consequently, it can be deduced that
$p\rightarrow 0\quad\Rightarrow\quad u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0.$ (51)
In the absence of pressure, local pressure gradients are nonexistent. Consequently, the emergence of local fluid flows is precluded. This implies that
$u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0\quad\Rightarrow\quad\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})\rightarrow\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x}).$ (52)
Therefore, in the event of a zero fluid flow, the density perturbations are observed to remain static. It can be concluded that the second requirement is fulfilled by the limits (51)–(52).
Given that the physical parts of the spatial components of the fluid four-velocity are identically zero, only the gauge modes $\psi_{|i}$ remain. It can thus be concluded with confidence that the gauge modes may be taken to be equal to zero without any loss of physical information. Taking $\hat{u}_{{\scriptscriptstyle(1)}i}=\psi_{|i}=0$ implies that $\psi$ is an arbitrary infinitesimal constant. Upon substituting $\psi=C$ into the expressions (6), the result is that the relativistic gauge transformation between synchronous coordinates reduces in the limits given in (51) to the gauge transformation of Newtonian gravitation:
$t\rightarrow t^{\prime}=t-C,\quad x^{i}\rightarrow x^{\prime i}=x^{i}-\chi^{i}(\bm{x}).$ (53)
This implies that within the Newtonian theory of gravitation, there is the possibility of shifting the time coordinate and selecting the spatial coordinates at will. In accordance with the expressions (53), the third requisite is thereby fulfilled.
The fourth and final requirement will now be discussed. Given that $u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}=0$ and $\hat{u}^{i}_{{\scriptscriptstyle(1)}}=0$, it follows that $\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k}=0$. By combining this result with (52), the linearized Einstein equations and conservation laws (47) are reduced to
$\nabla^{2}\phi(t,\bm{x})$ $=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}),$ (54a)
$\nabla^{2}\dot{\phi}(t,\bm{x})$ $=0,$ (54b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $=-3H\varepsilon_{\scriptscriptstyle(1)}-\varepsilon_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)},$ (54c)
$\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (54d)
The constraint equations (54a) and (54b) are combined to yield the following result:
$\nabla^{2}\phi(\bm{x})=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t_{0})\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}).$ (55)
The potential $\varphi(\bm{x})$ is defined as $\varphi(\bm{x}):=\phi(\bm{x})/a^{2}(t_{0})$, which results in the Poisson equation of Newtonian gravitation:
$\nabla^{2}\varphi(\bm{x})=4\pi G_{\text{N}}\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{c^{2}}.$ (56)
With (56), the fourth and final requirement has been met.
In the nonrelativistic limit, the background equations (10) are given by
$3H^{2}=\kappa\varepsilon_{\scriptscriptstyle(0)}+\Lambda,\quad\dot{\varepsilon}_{\scriptscriptstyle(0)}=-3H\varepsilon_{\scriptscriptstyle(0)},\quad\dot{n}_{\scriptscriptstyle(0)}=-3Hn_{\scriptscriptstyle(0)}.$ (57)
Consequently, the conservation laws imply that $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$. As may be observed, Eqs. (54c)–(54d) yield solely gauge modes as solutions: specifically, $\hat{\varepsilon}_{\scriptscriptstyle(1)}=C\dot{\varepsilon}_{\scriptscriptstyle(0)}$, $\hat{n}_{{\scriptscriptstyle(1)}}=C\dot{n}_{{\scriptscriptstyle(0)}}$, and $\hat{\theta}_{{\scriptscriptstyle(1)}}=C\dot{\theta}_{{\scriptscriptstyle(0)}}$, where $C$ is an infinitesimal constant of the gauge transformation (53) of Newtonian gravitation. This, combined with $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$ yields $\varepsilon_{{\scriptscriptstyle(1)}}=n_{{\scriptscriptstyle(1)}}mc^{2}$. Combining the latter two relations with the definitions (4a) one finds the well-known special relativistic relation between the rest mass density and energy density:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})mc^{2}.$ (58)
The equations (54c) and (54d) represent the residual form of the conservation laws (47c) and (47e) in the nonrelativistic limit. It is evident that equations Eqs. (54c) and (54d) have no physical significance. These equations are decoupled from the physical constraint equations (54a) and (54b) and thus do not form part of Newtonian gravitation. Therefore, the potential $\zeta$ which appears in $\theta_{{\scriptscriptstyle(1)}}$, (24), is no longer relevant. In view of the aforementioned considerations, it becomes evident that solely the Poisson equation (56) and the special relativistic relation (58) have survived.
Following an examination of the four points regarding the nonrelativistic limit, the following conclusion is reached. It has been demonstrated that the system of linearized Einstein equations and conservation laws for scalar perturbations (37), when combined with the gauge-invariant quantities (4a), reduces in the nonrelativistic limit to the Poisson equation (56) of Newtonian gravitation and the special relativistic relation (58). This indicates that the gauge-invariant quantities $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, defined by expressions (4a), represent the local perturbations of the energy density and the particle number density, respectively, in Newtonian gravitation. Therefore, it can be stated that $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ represent the actual local density perturbations of the energy density and the particle number density in general relativity. In conclusion, since the gauge-invariant perturbation of the expansion, $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ (4b), is identically zero, it can be stated that density perturbations have no influence on the expansion of the universe.
The subsequent section of this study will proceed to derive the evolutionary equations for $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$. These equations permit the investigation of the evolution of density fluctuations in closed, flat, and open flrw universes independently of the choice of coordinate system.
X Evolution Equations for Density Fluctuations
In light of the preceding analysis, it can be concluded that the quantities represented by $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, are the real physical perturbations to the energy density and particle number density, respectively. Furthermore, it has been demonstrated that the evolution equations for scalar perturbations (37), when combined with the background equations (10), govern the evolution of the energy density perturbation and the particle number density perturbation. In Appendix A, the derivation of evolution equations for their corresponding density fluctuations will be presented. These density fluctuations are defined as follows:
$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{\scriptscriptstyle(0)}(t)},\quad\delta_{n}(t,\bm{x}):=\dfrac{n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{n_{\scriptscriptstyle(0)}(t)}.$ (59)
The final result is a system of evolution equations for the density fluctuations $\delta_{\varepsilon}$ and $\delta_{n}$, in closed, flat, and open flrw universes. These equations are given by
$\ddot{\delta}_{\varepsilon}+b_{1}\dot{\delta}_{\varepsilon}+b_{2}\delta_{\varepsilon}$ $=b_{3}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right),$ (60a)
$\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right)$ $=\frac{3Hn_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right).$ (60b)
The coefficients $b_{1}$, $b_{2}$ and $b_{3}$ are as follows:
$b_{1}=$ $\,\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{H}-2\dfrac{\dot{\beta}}{\beta}-H(2+6w+3\beta^{2})$
$+{{}^{3}\!}R_{\scriptscriptstyle(0)}\left(\dfrac{1}{3H}+\dfrac{2H(1+3\beta^{2})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right),$ (61a)
$b_{2}=$ $-\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)(1+3w)$
$+H^{2}\left(1-3w+6\beta^{2}(2+3w)\right)$
$+6H\dfrac{\dot{\beta}}{\beta}\left(w+\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)$
$-{{}^{3}\!}R_{\scriptscriptstyle(0)}\left(\tfrac{1}{2}w+\dfrac{H^{2}(1+6w)(1+3\beta^{2})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)$
$-\beta^{2}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right),$ (61b)
$b_{3}=$ $\,\Biggl{[}\dfrac{-18H^{2}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\Biggl{(}\varepsilon_{\scriptscriptstyle(0)}p_{\varepsilon n}(1+w)+\dfrac{2p_{n}}{3H}\dfrac{\dot{\beta}}{\beta}$
$+p_{n}(p_{\varepsilon}-\beta^{2})+n_{\scriptscriptstyle(0)}p_{nn}\Biggr{)}+p_{n}\Biggr{]}$
×$\times\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right).$ (61c)
In these expressions the partial derivatives of the pressure, $p_{\varepsilon}$ and $p_{n}$, are defined by (13b). The second-order partial derivatives are defined by $p_{nn}:={\partial^{2}p}/{\partial n^{2}}$ and $p_{\varepsilon n}:={\partial^{2}p}/{\partial\varepsilon\,\partial n}$.
It is essential to recognize that, with regard to the nonrelativistic limit, the aforementioned equations apply solely in the case of $p\not\equiv 0$.
XI Thermodynamic Quantities
The combined first and second laws of thermodynamics are expressed in terms of density fluctuations, denoted by $\delta_{\varepsilon}$ and $\delta_{n}$ and gauge-invariant expressions are provided for the local pressure and temperature perturbations.
XI.1 Thermodynamic Laws
The combined first and second laws of thermodynamics for a simple, single-species system is given by (see, for example, Ref. [38], Sec. 2.1)
${\mathrm{d}}E=T{\mathrm{d}}S-p{\mathrm{d}}V+\mu{\mathrm{d}}N,$ (62)
where $E$, $S$, and $N$ are the energy, the entropy and the number of particles of a system with volume $V$, temperature $T$, and pressure $p$. The thermal—or chemical—potential $\mu$, is the energy required to add one particle to the system. In terms of the particle number density $n=N/V$, the energy per particle $E/N=\varepsilon/n$ and the entropy per particle $s=S/N$ the law (62) can be rewritten in the form
${\mathrm{d}}\left(\dfrac{\varepsilon}{n}N\right)=T{\mathrm{d}}(sN)-p{\mathrm{d}}\left(\dfrac{N}{n}\right)+\mu{\mathrm{d}}N,$ (63)
where $\varepsilon$ is the energy density. The system is extensive, i.e., $S(\lambda E,\lambda V,\lambda N)=\lambda S(E,V,N)$, which implies that the entropy of the gas is given by $S=(E+pV-\mu N)/T$. Upon dividing this relation by $N$ the Euler relation is obtained:
$\mu=\dfrac{\varepsilon+p}{n}-Ts.$ (64)
Eliminating $\mu$ in (63) with the aid of (64) reveals that the combined first and second laws of thermodynamics  (62) can be expressed in a form that does not include $\mu$ or $N$ (see Ref. [33]):
$T{\mathrm{d}}s={\mathrm{d}}\Bigl{(}\dfrac{\varepsilon}{n}\Bigr{)}+p{\mathrm{d}}\Bigl{(}\dfrac{1}{n}\Bigr{)}.$ (65)
From the background equations (10) and the thermodynamic law (65) it can be demonstrated that $\dot{s}_{\scriptscriptstyle(0)}=0$, indicating that the expansion of the universe takes place without generating entropy. From the transformation (3) it can be concluded that $s_{\scriptscriptstyle(1)}=s_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ is automatically gauge-invariant. By employing (138) and $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$, the thermodynamic relation (65) can be rewritten as
$T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n^{2}_{\scriptscriptstyle(0)}}\left(n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\right).$ (66)
By employing the definitions (59), the following is obtained:
$T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n_{\scriptscriptstyle(0)}}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (67)
Therefore, $s_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}$ represents the local perturbation to the entropy per particle.
XI.2 Temperature and Pressure Perturbations
Given that both the background temperature, denoted by $T_{{\scriptscriptstyle(0)}}$, and the pressure, represented by $p_{{\scriptscriptstyle(0)}}$, are scalars, their gauge-invariant temperature and pressure perturbations can be defined in a manner analogous to that of the expressions (4):
$T^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=T_{\scriptscriptstyle(1)}-\dfrac{\dot{T}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad p^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=p_{\scriptscriptstyle(1)}-\dfrac{\dot{p}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}.$ (68)
In order to arrive at the gauge-invariant counterpart of (13a), it is first necessary to eliminate $\varepsilon_{{\scriptscriptstyle(1)}}$ and $n_{{\scriptscriptstyle(1)}}$ from (13a) using (4a). The utilization of $\dot{p}_{{\scriptscriptstyle(0)}}=p_{n}\dot{n}_{{\scriptscriptstyle(0)}}+p_{\varepsilon}\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and (68), results in
$p^{\textnormal{phys}}_{\scriptscriptstyle(1)}=p_{n}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (69)
Eliminating $p_{\varepsilon}$ with the aid of (40) and utilizing the definitions provided in (59) for density fluctuations results in the following expression:
$p_{\scriptscriptstyle(1)}^{\textnormal{phys}}=\beta^{2}\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}+n_{\scriptscriptstyle(0)}p_{n}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (70)
The first term represents the adiabatic component of the pressure perturbation, while the second term denotes the nonadiabatic component.
XII Application: Structure Formation in a Flat FLRW Universe
The evolution equations presented in (60), are applicable to open, flat, and closed flrw universes. Henceforth, the discussion will be focused on the flat, ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, universe. Given that the initial density fluctuations occurred in the early universe when $\Lambda\ll\kappa\varepsilon_{\scriptscriptstyle(0)}$, the cosmological constant, $\Lambda$, will be neglected. Consequently, the background equations, as given in Eq. (10), are reduced to
$3H^{2}$ $=\kappa\varepsilon_{\scriptscriptstyle(0)},\quad$ $\kappa$ $=8\pi G_{\text{N}}/c^{4},$ (71a)
$\dot{\varepsilon}_{\scriptscriptstyle(0)}$ $=-3H\varepsilon_{\scriptscriptstyle(0)}(1+w),$ $\quad w$ $:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)},$ (71b)
$\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (71c)
By employing the Friedmann equation (71a), the coefficients (61) of Eq. (60a) result in
$b_{1}$ $=H(1-3w-3\beta^{2})-2\dfrac{\dot{\beta}}{\beta},$ (72a)
$b_{2}$ $=\kappa\varepsilon_{\scriptscriptstyle(0)}\Bigl{(}2\beta^{2}(2+3w)-\tfrac{1}{6}(1+18w+9w^{2})\Bigr{)}$
$+2H\dfrac{\dot{\beta}}{\beta}(1+3w)-\beta^{2}\dfrac{\nabla^{2}}{a^{2}},$ (72b)
$b_{3}$ $=\Biggl{[}\dfrac{-2}{1+w}\Biggl{(}\varepsilon_{\scriptscriptstyle(0)}p_{\varepsilon n}(1+w)+\dfrac{2p_{n}}{3H}\dfrac{\dot{\beta}}{\beta}$
$+p_{n}(p_{\varepsilon}-\beta^{2})+n_{\scriptscriptstyle(0)}p_{nn}\Biggr{)}+p_{n}\Biggr{]}\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}\dfrac{\nabla^{2}}{a^{2}},$ (72c)
where $\nabla^{2}$ represents the conventional Laplace operator.
XII.1 Era before Decoupling of Matter and Radiation
In this era the primordial fluid is a mixture of radiation and matter, wherein the contribution of matter to the pressure is negligible. Consequently, the equations of state (see, for example, Ref. [8], Eq. (5.49), Ref. [10], Eq. (1.78), and Ref. [17], § V-1) are given by:
$\varepsilon=a_{\text{B}}T_{\gamma}^{4}+nmc^{2},\quad p=\tfrac{1}{3}a_{\text{B}}T_{\gamma}^{4}.$ (73)
The black body constant is represented by $a_{\text{B}}$, the radiation temperature is denoted by $T_{\gamma}$ and the particle number density of ordinary matter or cdm is given by $n$. Upon eliminating $T_{\gamma}$, the following result is obtained [see Ref. [31], Eq. (2.10.27)]:
$p=\tfrac{1}{3}(\varepsilon-nmc^{2}).$ (74)
By making use of (13b), one has
$p_{n}=-\tfrac{1}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{1}{3}.$ (75)
For the parameter $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and the nonadiabatic speed of sound $\beta$, (40), the following expressions are derived:
$w=\frac{\tfrac{1}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}},\quad\beta^{2}=\frac{1}{3}-\frac{\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{\tfrac{4}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}}.$ (76)
The ensuing subsections provide a concise overview of the radiation-dominated and matter-dominated eras preceding decoupling.
XII.1.1 Radiation-dominated Era
The universe was radiation-dominated when
$a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}\gg n_{{\scriptscriptstyle(0)}}mc^{2}.$ (77)
In this case one has $\beta^{2}\approx\tfrac{1}{3}$, and $w\approx\tfrac{1}{3}$, implying that $\dot{\beta}\approx 0$ and $p_{\varepsilon}\approx w$. Upon substituting the aforementioned values and expressions (75) into the coefficients (72), the following result is obtained:
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}$
$\qquad\qquad\qquad\;\,\approx\dfrac{-\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\dfrac{\nabla^{2}}{a^{2}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)},$ (78a)
$\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}\approx-\tfrac{3}{4}H\dfrac{n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}.$ (78b)
From (77), it can be deduced that the pressure term on the left-hand side of equation (78a) is, in absolute value, significantly larger than the source term. Consequently, the source term may be neglected. In the radiation-dominated phase, the system of equations (60) can be expressed as follows:
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}\approx 0,$ (79a)
$\left|\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\right|\rightarrow 0.$ (79b)
The limit in Eq. (79b) is a consequence of the fact that the coefficient in Eq. (78b) is negative, since $p_{n}<0$.
In order to solve Eq. (79a), it is necessary to employ the solutions of the background equations (71). These solutions are provided by:
$H=\tfrac{1}{2}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{3}{4}(ct)^{-2},\quad a\propto t^{1/2}.$ (80)
The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{0}\geq 1$. This definition implies that
$\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{0}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl{(}2H(t_{0})\bigr{)}^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (81)
By employing the Helmholtz equation $\nabla^{2}\delta_{\varepsilon}=-|\bm{q}|^{2}\delta_{\varepsilon}$ and the result derived in (81), it can be shown that the equation (79a) can be rewritten as
$\delta_{\varepsilon}^{\prime\prime}-\dfrac{1}{2\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{\mu_{\text{r}}^{2}}{4\tau}+\dfrac{1}{2\tau^{2}}\right)\delta_{\varepsilon}=0.$ (82)
In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{r}}$ is defined as follows:
$\mu_{\text{r}}:=\dfrac{2\pi}{\lambda_{0}}\dfrac{1}{H(t_{0})}\dfrac{1}{\sqrt{3}}.$ (83)
In this expression, $\lambda_{0}:=\lambda a(t_{0})$ represents the physical scale of a fluctuation at $t=t_{0}$, with the magnitude of the fluctuation given by $\lvert\bm{q}\rvert=2\pi/\lambda_{0}$.
In order to solve Eq. (82), it is first necessary to replace the variable $\tau$ with the variable defined by $x:=\mu_{\text{r}}\sqrt{\tau}$. Upon transforming back to $\tau$, the general solution of Eq. (82) with constants of integration $A_{1}(\bm{q})$ and $A_{2}(\bm{q})$ is obtained, namely
$\delta_{\varepsilon}(\tau,\bm{q})=\bigl{[}A_{1}(\bm{q})\sin\left(\mu_{\text{r}}\sqrt{\tau}\right)+A_{2}(\bm{q})\cos\left(\mu_{\text{r}}\sqrt{\tau}\right)\bigr{]}\sqrt{\tau},$ (84)
It can be inferred from this solution that, during the radiation-dominated era of the universe, density fluctuations oscillated with an amplitude that grew in a manner proportional to the square root of time.
In the case of large-scale fluctuations, that is, ${\lambda_{0}\rightarrow\infty}$, which leads to $\mu_{\text{r}}\rightarrow 0$, (83), the general solution of Eq. (82) is given by
$\delta_{\varepsilon}(\tau)=-\bigl{[}\delta_{\varepsilon}(1)-2\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau+2\bigl{[}\delta_{\varepsilon}(1)-\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau^{1/2}.$ (85)
This solution, with the exception of the precise factors of proportionality, has been derived by a significant number of authors, as evidenced in Refs. [2, 39, 40, 41, 42].
A comparison of the solutions (85) and (155a) reveals that the solution ${\propto\tau}$ is a consequence of the homogeneous part of equation (154a). The solution $\propto\tau^{1/2}$ is the particular solution of the inhomogeneous equation. This solution can only be obtained if the covariant divergence, $\vartheta_{{\scriptscriptstyle(1)}}$, is taken into account, as demonstrated in Appendix C.1. The minus sign preceding the solution $\propto\tau$ is a direct consequence of the fact that ${\vartheta_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}>0}$, as follows from (155b). The presence of the gauge function $\psi$ precludes the derivation of the solution (85) from (154). As demonstrated in Appendix A, $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the left-hand side of (78a), as evidenced by the derivation of (60a).
XII.1.2 Matter-dominated Era
The radiation-dominated phase came to an end and the matter-dominated phase commenced at $z(t_{\text{eq}})=3387$, as indicated in Table 1, well before decoupling [43], when
$a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}=n_{{\scriptscriptstyle(0)}}mc^{2}.$ (86)
Given that $p_{n}<0$, it can be shown from Eq. (60b) that
$\left|\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right|\rightarrow 0.$ (87)
It can thus be seen that fluctuations in the energy density are coupled to those in the particle number density. In light of the fact that the evolution equations (60) are general relativistic, it follows that the weak equivalence principle is valid. This implies that the evolution of fluctuations in the particle number density is independent of the composition of matter. Therefore, if cdm consists of particles that interact via gravitation, the fluctuations in the cdm are also linked to the fluctuations in the energy density. This implies that it is impossible for cdm to have contracted before decoupling. Consequently, cdm cannot have initiated the formation of structure after decoupling.
XII.2 Era after Decoupling of Matter and Radiation
Once protons and electrons have combined to form hydrogen, the radiation pressure will be negligible, and the equations of state will be those of a nonrelativistic monatomic perfect gas with three degrees of freedom. These are given by the equations of state [see Ref. [31], Eqs. (15.8.20)–(15.8.21), and Ref. [33], Eq. (13)]:
$\varepsilon(n,T)=nmc^{2}+\tfrac{3}{2}nk_{\text{B}}T,\quad p(n,T)=nk_{\text{B}}T.$ (88)
In these expressions, the symbol $k_{\text{B}}$ represents Boltzmann’s constant, the quantity $m$ denotes the proton rest mass, and the temperature of the matter is indicated by the symbol $T$. The rest mass energy density is $nmc^{2}$ and the kinetic energy density is $\tfrac{3}{2}nk_{\text{B}}T$. The elimination of the temperature, $T$, from the equations of state (88), yields the equation of state for the pressure [see Ref. [31], Eq. (2.10.27)]:
$p(n,\varepsilon)=\tfrac{2}{3}(\varepsilon-nmc^{2}).$ (89)
The partial derivatives are determined by the utilization of the expressions (13b) and are given by
$p_{n}=-\tfrac{2}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{2}{3}.$ (90)
The parameter $w$ is defined by
$w:=\dfrac{p_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}=\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}+\tfrac{3}{2}k_{\text{B}}T_{\scriptscriptstyle(0)}}\approx\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}\ll 1.$ (91)
Upon substituting (88), (90) and (91) into (40), one arrives at the well-known result [see Ref. [31], Eq. (15.8.22)]:
$\quad\beta\approx\dfrac{v_{\text{s}}}{c}=\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}}.$ (92)
In this expression, $v_{\text{s}}$ represents the adiabatic speed of sound. From $\beta^{2}\approx\tfrac{5}{3}w$, it is found that Eq. (44) results in $\dot{w}\approx-2Hw$. Therefore, with $H:=\dot{a}/a$, it can be concluded that $w\propto a^{-2}$. Given that $w$ is proportional to $T_{{\scriptscriptstyle(0)}}$, the well-known result (see Ref. [31], Eq. (15.5.16) with $\gamma=\tfrac{5}{3}$, where ${\gamma:=c_{\text{p}}/c_{\text{v}}}$ is the adiabatic index) is obtained:
$T_{\scriptscriptstyle(0)}\propto a^{-2}.$ (93)
In light of the relatively minor values of both $w$ and $\beta^{2}$, it is now possible to justify the equations of state (88). From the proportionality (93) and the fact that $n_{{\scriptscriptstyle(0)}}\propto a^{-3}$, as can be seen from Eq. (71c), it follows from (88) that ${p_{{\scriptscriptstyle(0)}}\propto a^{-5}}$. Given that pressure decayed rapidly and was already negligible in comparison to the energy density at the time of decoupling, it can be concluded that the unperturbed pressure had a minimal impact on structure formation. It is therefore reasonable to neglect the quantities $w$ and $\beta^{2}$ with respect to the constants of order one that appear in the momentum conservation laws (45). In doing so, one finds
$\dot{u}^{i}_{\scriptscriptstyle(1)}+2Hu^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}-\dfrac{p_{{\scriptscriptstyle(1)}|j}}{\varepsilon_{\scriptscriptstyle(0)}}=0.$ (94)
The second equation is derived by lowering the index $i$ in the first equation using the background metric $g_{{\scriptscriptstyle(0)}ij}$ and the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$. Accordingly, the evolution of density fluctuations is highly dependent upon the presence of pressure gradients, represented by $p_{{\scriptscriptstyle(1)}|j}$. In the event that the pressure gradients are identically zero, Eqs. (94) reduce to the form of Eqs. (49) and (50) in the nonrelativistic limit. This implies that the physical component of the fluid velocity is zero, and consequently, density perturbations cease to evolve. Given that structure formation in the universe was predominantly determined by pressure gradients, it is necessary to consider the kinetic energy density and pressure in addition to the energy density. This is the reason why the particle number density must be taken into account in the equations of state. In conclusion, it can be inferred from the equations (94) that the gauge modes (28) are zero. This result demonstrates that the relativistic gauge transformations (1), where $\xi^{\mu}$ is given by (6), are equivalent to the Newtonian gauge transformations (53). This conclusion is consistent with expectations, given that the equations of state (88) are nonrelativistic.
The subsequent step is to derive the evolution equations for density perturbations. The proportionality (93) implies with (92) that $\dot{\beta}/\beta=-H$. The system of equations (60) with coefficients (72) can now be rewritten in the following form:
$\ddot{\delta}_{\varepsilon}+3H\dot{\delta}_{\varepsilon}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{5}{6}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}=-\dfrac{2}{3}\dfrac{\nabla^{2}}{a^{2}}\left(\delta_{n}-\delta_{\varepsilon}\right),$ (95a)
$\dfrac{1}{c}\dfrac{\mathrm{d}}{{\mathrm{d}}t}\left(\delta_{n}-\delta_{\varepsilon}\right)=-2H\left(\delta_{n}-\delta_{\varepsilon}\right),$ (95b)
In this calculation, the approximation is made that $w$ and $\beta^{2}$ are negligible with respect to the constants of order one.
Given that the term $(\delta_{n}-\delta_{\varepsilon})$ occurs within the source term of Eq. (95a), it is necessary to first solve for the solution to Eq. (95b). By making the substitution $H:=\dot{a}/a$, the following result is obtained:
$\delta_{n}-\delta_{\varepsilon}\propto a^{-2}.$ (96)
In order to relate this proportionality to thermodynamic quantities, it is necessary to express the perturbed equation of state for the energy density in terms of density fluctuations. It can be deduced from $\varepsilon=\varepsilon(n,T)$ that
$\dot{\varepsilon}_{\scriptscriptstyle(0)}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}\dot{n}_{\scriptscriptstyle(0)}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}\dot{T}_{\scriptscriptstyle(0)},$ (97a)
$\varepsilon_{\scriptscriptstyle(1)}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}.$ (97b)
By multiplying (97a) by $\theta_{\scriptscriptstyle(1)}/\dot{\theta}_{\scriptscriptstyle(0)}$ and subtracting the result from (97b), one obtains the following result:
$\varepsilon_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}.$ (98)
This result is based on the use of expressions (4a) and (68). By employing the expression for $\varepsilon$ in (88) to eliminate the partial derivatives, one obtains the perturbed equation of state:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}mc^{2}+\tfrac{3}{2}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}k_{\text{B}}T_{\scriptscriptstyle(0)}+\tfrac{3}{2}n_{\scriptscriptstyle(0)}k_{\text{B}}T^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (99)
By dividing expression (99) by $\varepsilon_{\scriptscriptstyle(0)}$, and using the exact value of the ratio $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, the perturbed equation of state for the energy density expressed in fluctuations (59) is obtained:
$\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})=-\tfrac{3}{2}w(t)\delta_{T}(t,\bm{x}).$ (100)
In this context, the quantity $\delta_{T}$ is defined by $\delta_{T}:=T^{\textnormal{phys}}_{\scriptscriptstyle(1)}/T_{\scriptscriptstyle(0)}$. The solution to Eq. (95b) is obtained by combining (96) and (100) and using the approximate value of $w$, (91), which implies that $w\propto a^{-2}$. The result is as follows:
$\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})\approx-\tfrac{3}{2}w(t)\delta_{T}(t_{\text{dec}},\bm{x}).$ (101)
The quantity $\delta_{T}(t_{\text{dec}},\bm{x})$ represents the temperature fluctuation of matter at the time $t_{\text{dec}}$, which marks the decoupling of matter from radiation. Since $w\ll 1$, (101) is in alignment with (87).
According to (67), (91) and (101) the entropy per particle is as follows:
$s^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})\approx\tfrac{3}{2}k_{\text{B}}\delta_{T}(t_{\text{dec}},\bm{x}).$ (102)
Therefore, it can be concluded that the quantity $\delta_{T}$ is random and that density fluctuations are nonadiabatic if $\delta_{T}$ is nonzero.
Upon substituting (90), (92), (101), and the approximation ${w\ll 1}$, the local pressure fluctuation, defined by $\delta_{p}:=p^{\textnormal{phys}}_{\scriptscriptstyle(1)}/p_{\scriptscriptstyle(0)}$, can be calculated from expression (70). One obtains
$\delta_{p}(t,\bm{x})\approx\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})+\delta_{T}(t_{\text{dec}},\bm{x}).$ (103)
In this context, the terms $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ and $\delta_{T}(t_{\text{dec}},\bm{x})$ represent the adiabatic and nonadiabatic pressure fluctuations, respectively.
The solution to Eq. (95b) has been determined. This allows for the reformulation of the second-order equation (95a) to facilitate analysis of the evolution of density fluctuations. To that end, the solutions to the background equations (71) are required. Given that the pressure with respect to the rest mass energy density can be disregarded ($w\ll 1$), the following solutions are obtained:
$H=\tfrac{2}{3}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{4}{3}(ct)^{-2},\quad a\propto t^{2/3}.$ (104)
The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{\text{dec}}\geq 1$. This definition implies that
$\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{\text{dec}}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl{(}\tfrac{3}{2}H(t_{\text{dec}})\bigr{)}^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (105)
By employing Eqs. (71a) and (101), as well as the expressions (92), (104) (105), the Helmholtz equation $\nabla^{2}\delta=-|\bm{q}|^{2}\delta$, and $w\propto a^{-2}$, it can be demonstrated that Eq. (95a) can be expressed in the following manner:
$\delta_{\varepsilon}^{\prime\prime}+\dfrac{2}{\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{4}{9}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}-\dfrac{10}{9\tau^{2}}\right)\delta_{\varepsilon}=-\dfrac{4}{15}\dfrac{\mu^{2}_{\text{m}}}{\tau^{8/3}}\delta_{T}(t_{\text{dec}},\bm{q}).$ (106)
In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{m}}$ is defined as follows:
$\mu_{\text{m}}:=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{H(t_{\text{dec}})}\dfrac{v_{\text{s}}(t_{\text{dec}})}{c}.$ (107)
In this expression, $\lambda_{\text{dec}}:=\lambda a(t_{\text{dec}})$ represents the physical scale of a fluctuation at $t=t_{\text{dec}}$. The magnitude of the fluctuation is given by $\lvert\bm{q}\rvert=2\pi/\lambda_{\text{dec}}$. Furthermore, one has $v_{\text{s}}(t_{\text{dec}})/c\approx\beta(t_{\text{dec}})$, as stated in expression (92).
Finally, the expression (103) enables the reformulation of Eq. (106) in a form suitable for the study of the evolution of density fluctuations:
$\delta^{\prime\prime}_{\varepsilon}+\frac{2}{\tau}\delta^{\prime}_{\varepsilon}+\dfrac{4}{15}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}\delta_{p}-\dfrac{10}{9\tau^{2}}\delta_{\varepsilon}=0,\quad\tau:=\dfrac{t}{t_{\text{dec}}}\geq 1.$ (108)
The second term in this equation represents the expansion, the third term is the pressure term, where $\delta_{p}$ is given by (103), and the fourth term represents gravitation.
In the limit of large scales, that is, $\lambda_{\text{dec}}\rightarrow\infty$, which leads to $\mu_{\text{m}}\rightarrow 0$, the solution of Eq. (108) is
$\delta_{\varepsilon}(\tau)=\bigl{[}\tfrac{5}{7}\delta_{\varepsilon}(1)+\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau^{2/3}+\bigl{[}\tfrac{2}{7}\delta_{\varepsilon}(1)-\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau^{-5/3}.$ (109)
Therefore, pressure fluctuations $\delta_{p}$ played a negligible role in the evolution of the large-scale fluctuations. In other words, these fluctuations were almost adiabatic and evolved solely under the influence of gravitation and expansion.
A comparison of the solutions (109) and (157a) reveals that the solution ${\propto\tau^{2/3}}$ is a consequence of the homogeneous part of Eq. (156a). The solution $\propto\tau^{-5/3}$ is the particular solution to the inhomogeneous equation. This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.2. In contrast to the standard equation (150), the quantity $\vartheta_{{\scriptscriptstyle(1)}}$ is included in Eq. (95a). The presence of the gauge constant $C$, precludes the derivation of solution (109) from the system (156).
XII.2.1 Cosmological Quantities
To investigate the growth of density fluctuations, it is first necessary to express the parameter $\mu_{\text{m}}$ (107) in observable quantities.
The redshift $z(t)$ is defined by the expression
$z(t):=\dfrac{a(t_{\text{p}})}{a(t)}-1,$ (110)
where $a(t_{\text{p}})$ is the present value of the scale factor. In the case of a flat flrw universe, it is permissible to take $a(t_{\text{p}})=1$. Using the background solutions (104), it is possible to express the relevant parameters of the universe in the redshift. The result is
$H(t)$ $=H(t_{\text{p}})\bigl{(}z(t)+1\bigr{)}^{3/2},$ (111a)
$t$ $=t_{\text{p}}\bigl{(}z(t)+1\bigr{)}^{-3/2},$ (111b)
$T_{{\scriptscriptstyle(0)}\gamma}(t)$ $=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})\bigl{(}z(t)+1\bigr{)},$ (111c)
$n_{{\scriptscriptstyle(0)}}(t)$ $=n_{{\scriptscriptstyle(0)}}(t_{\text{p}})\bigl{(}z(t)+1\bigr{)}^{3}.$ (111d)
In deriving relation (111c), it is utilised that, subsequent to decoupling, $T_{{\scriptscriptstyle(0)}\gamma}\propto a^{-1}$. The dimensionless time, defined as $\tau:=t/t_{\text{dec}}$, is given by
$\tau=\left(\dfrac{z(t_{\text{dec}})+1}{z(t)+1}\right)^{3/2},$ (112)
where relation (111b) with $z(t_{\text{p}})=0$ is used. Upon substituting the result (92) into expression (107) and utilizing the relations (111), the following is obtained:
$\mu_{\text{m}}=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{cH(t_{\text{p}})}\dfrac{1}{z(t_{\text{dec}})+1}\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})}{m}},$ (113)
where it is assumed that at decoupling, the matter and radiation temperatures were equal, that is, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. In this manner, the parameter $\mu_{\text{m}}$ has been expressed in observable quantities. Upon substituting the numerical values from Tables 1 and 2, one obtains the result:
$\mu_{\text{m}}=\dfrac{16.48}{\lambda_{\text{dec}}},\quad\lambda_{\text{dec}}\text{ in pc}.$ (114)
It can thus be concluded that the parameter $\mu_{\text{m}}$ is dependent solely on the initial scale $\lambda_{\text{dec}}$ of a density fluctuation.
Table 1: Planck satellite results
$z(t_{\text{dec}})$ = 1090
$z(t_{\text{eq}})$ = 3387
$cH(t_{\text{p}})$ = $67.66\,\text{km}\,\text{s}^{-1}\text{Mpc}^{-1}$
[…]
$\lvert\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})\rvert$ $\lesssim$ $10^{-5}$
XII.2.2 Initial Values
In order to solve Eq. (108), it is necessary to determine the initial values of the quantities $\delta_{\varepsilon}$ and $\delta^{\prime}_{\varepsilon}$. Furthermore, it is imperative to ascertain the values of the local nonadiabatic random pressure fluctuation $\delta_{T}$ in expression (103).
Planck Satellite
The Planck observations [43] of the fluctuations $\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})$ in the background radiation temperature imply that $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$. In the absence of knowledge regarding the initial growth rate, it is assumed that $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$. Therefore, the initial values for Eq. (108) are as follows:
$\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\lesssim 10^{-5},\quad\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0.$ (115)
As demonstrated in the calculations presented in Sec. XII.2.3, the outcome of Eq. (108) is largely independent of the initial value for $\delta_{\varepsilon}$, provided that it satisfies the condition of having a maximum value of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\leq 10^{-4}$.
Nonadiabatic Pressure Fluctuations
The values of the quantities $\delta_{\varepsilon}(t,\bm{q})$ and $\delta_{n}(t,\bm{q})$ are related to the local nonadiabatic random pressure fluctuation $\delta_{T}(t_{\text{dec}},\bm{q})$ by Eq. (101). Immediately following the decoupling, this equation reads:
$\delta_{n}(t_{\text{dec}},\bm{q})-\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx-4.1\times 10^{-10}\delta_{T}(t_{\text{dec}},\bm{q}),$ (116)
where it has been assumed that at decoupling, the matter temperature was equal to the radiation temperature, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. Expressions (91) and (111c) were employed in conjunction with the values presented in Tables 1 and 2.
The transition of the universe from the era preceding decoupling to the era following decoupling was rapid and chaotic. A notable decline was observed in the mean particle velocity, and the pressure. Furthermore, the values of both $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ exhibited minor irregularities on the surface of last scattering. Therefore, it is probable that after the transition, the initial values of the difference between fluctuations, represented by ${(\delta_{n}-\delta_{\varepsilon})(t_{\text{dec}},\bm{q})}$, were randomly distributed among all density fluctuations. From expression (116), it can be seen that very small differences between the initial energy density and particle number density fluctuations result in relatively large positive or negative fluctuations in the nonadiabatic component of the pressure fluctuations in expression (103). Therefore, the nonadiabatic pressure fluctuations, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, are randomly (102) distributed among the multitude of density fluctuations.
Table 2: Physical constants
$m$ = $1.6726\times 10^{-27}\,\text{kg}$
pc = $3.0857\times 10^{16}\,\text{m}=3.2616\,\text{ly}$
$c$ = $2.9979\times 10^{8}\,\text{m}\,\text{s}^{-1}$
$k_{\text{B}}$ = $1.3806\times 10^{-23}\,\text{J}\,\text{K}^{-1}$
$G_{\text{N}}$ = $6.6743\times 10^{-11}\,\text{m}^{3}\,\text{kg}^{-1}\,\text{s}^{-2}$
$\text{M}_{\odot}$ = $1.9889\times 10^{30}\,\text{kg}$
$a_{\text{B}}$ = $7.5657\times 10^{-16}\,\text{J}\,\text{m}^{-3}\,\text{K}^{-4}$
XII.2.3 Structure Formation
The subsequent analysis will employ Eq. (108) to examine the evolution of density fluctuations. This equation demonstrates that the evolution of a fluctuation $\delta_{\varepsilon}(t,\bm{q})$ with an initial scale $\lambda_{\text{dec}}$, was influenced by three key factors: the expansion of the universe, fluctuations in pressure, and gravitation. Fluctuations in pressure, as indicated by (103), were found to consist of two discrete components: an adiabatic contribution, represented by $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{q})$, and a nonadiabatic contribution, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, where the adiabatic component was initially negligible (115) and the nonadiabatic component was observed to be random.
[…]
Fig. 1 provides a clear summary of the evolution of density fluctuations. It was created using the following methodology. For each value of ${\delta_{T}=-0.01}$, $-0.05$, $-0.10$, $-0.15$, and $-0.20$, Eq. (108) is numerically solved for a large number of values for the initial fluctuation scale $\lambda_{\text{dec}}$ using the initial values (115).
[…]
$d_{\text{H}}(t_{\text{dec}})=a(t_{\text{dec}})\int_{0}^{ct_{\text{dec}}}\!\!\dfrac{\mathrm{d}\tau}{a(\tau)}=3ct_{\text{dec}}\approx 3.5\times 10^{5}\,\text{pc}.$ (117)
This value has been calculated using Eq. (111b) and Tables 1 and 2.
Figure 2: The graphs show the growth rates $\delta^{\prime}_{\varepsilon}$, with initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$, as function of the redshift $z$, or time in million of years. The initial scales $\lambda_{\text{dec}}$ of the fluctuations are measured in parsec. The evolution of density fluctuations started at $z=1090$.
[…]
$M(t_{\text{dec}})=\dfrac{4\pi}{3}\bigl{(}\tfrac{1}{2}\lambda_{\text{dec}}\bigr{)}^{3}n_{\scriptscriptstyle(0)}(t_{\text{dec}})m.$ (118)
It is first necessary to determine the particle number density at the time of decoupling denoted by $n_{\scriptscriptstyle(0)}(t_{\text{dec}})$, in order to proceed. This quantity is derived from the Friedmann equation (71a). For $w\ll 1$ and at time $t_{\text{p}}$ this equation is given by:
$3H^{2}(t_{\text{p}})=\kappa n_{{\scriptscriptstyle(0)}}(t_{\text{p}})mc^{2},\quad\kappa=8\pi G_{\text{N}}/c^{4}.$ (119)
By employing the relation (111d), one obtains
$n_{\scriptscriptstyle(0)}(t_{\text{dec}})=\frac{3\bigl{[}cH(t_{\text{p}})\bigr{]}^{2}\bigl{[}z(t_{\text{dec}})+1\bigr{]}^{3}}{8\pi mG_{\text{N}}}.$ (120)
[…]
$\lambda_{\text{Jeans}}(t_{\text{dec}}):=\lambda_{\text{Jeans}}a(t_{\text{dec}})\approx 6.4\,\text{pc}.$ (121)
Accordingly, the Jeans mass at decoupling is thus given by
$M_{\text{Jeans}}(t_{\text{dec}})\approx 2.2\times 10^{4}\,\text{M}_{\odot}.$ (122)
As evidenced by the preceding analysis, fluctuations in density, with initial values of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$ and $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$, result in the formation of structures in a flat flrw universe, provided that pressure fluctuations are taken into account. This marks the conclusion of the analysis concerning the evolution of density fluctuations in a flat flrw universe.
XIII Summary and Conclusion
This paper addresses four issues in the field of cosmological perturbation theory for Friedmann-Lemaître-Robertson-Walker (flrw) universes. The first and most significant issue concerns the gauge problem. It is well-established that the linearized Einstein equations and conservation laws yield solutions that are both physical and nonphysical. Consequently, the imposition of physical initial conditions does not result in solutions with physical significance. To resolve this issue, it is crucial to consider gauge-invariant quantities, that is, quantities that are independent of the choice of a coordinate system. It has been demonstrated that perturbations in the energy density and the particle number density can be constructed in a unique manner. In the nonrelativistic limit, it is demonstrated that these quantities represent the actual physical perturbations of the energy density and particle number density in the general theory of relativity. Similarly, as has been done for density perturbations, gauge-invariant perturbations in temperature, pressure, and entropy can be constructed. This addresses the gauge issue in the field of cosmology.
The second issue is that of the growth of minor density fluctuations subsequent to the decoupling of matter and radiation. In a fluid with zero pressure, density fluctuations do not evolve. To address this issue, it is necessary to incorporate the particle number density in addition to the energy density in the linearized Einstein equations and conservation laws. This allows for the kinetic energy density and pressure to be accounted for, resulting in the emergence of local pressure gradients. These gradients give rise to local fluid flows that are instrumental in the formation of structures after decoupling.
A third result is the novel perturbation theory (60), which provides a clear and unambiguous description of the evolution of fluctuations in the particle number density the energy density, as well as their interaction in a perfect fluid where the pressure depends on both the particle number density and the energy density. The aforementioned equations are not only applicable to a flat flrw universe, but also to open and closed universes. Furthermore, the cosmological constant need not be zero. It has been demonstrated that Eqs. (60) yield all previously established physical solutions for large-scale density perturbations in a flat flrw universe in its radiation-dominated phase and in the era following the decoupling of matter and radiation. In the development of the new cosmological perturbation theory, it has been demonstrated that in the nonrelativistic limit, relativistic gauge transformations reduce to Newtonian gauge transformations. Consequently, the conventional relativistic evolution equation—whether derived from general relativity or Newtonian gravitation modified to account for the expansion of the universe—also exhibits gauge-dependent solutions. This finding indicates that Newtonian gravitation is inadequate in accurately describing the evolution of density perturbations.
[…]
The investigation into cosmological density perturbations in a flat flrw universe reveals that if a nonrelativistic monatomic perfect gas with three degrees of freedom accurately represents the cosmic fluid in our universe following the decoupling of matter and radiation, then a perturbation theory based on general relativity provides an explanation for the origin of the first structures in the universe that does not rely on cdm. In contrast, the theory considers the local nonadiabatic random pressure fluctuations that emerged during the turbulent transition to the era that followed the decoupling of matter and radiation. Subsequent observations with the James Webb Space Telescope are expected to yield novel insights into the conditions that ensued immediately after the decoupling of radiation and matter, as well as the formation of the earliest structures in the universe.
Appendix A Derivation of the Evolution Equations
Table 3: The coefficients $\alpha_{ij}$ figuring in the equations (125)
$3H(1+p_{\varepsilon})+\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{2H}$ $3Hp_{n}$ $\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{4H}$
$\dfrac{\kappa n_{\scriptscriptstyle(0)}}{2H}$ $3H$ $n_{\scriptscriptstyle(0)}$ $\dfrac{n_{\scriptscriptstyle(0)}}{4H}$
$\dfrac{p_{\varepsilon}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $\dfrac{p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $H(2-3\beta^{2})$ 00
$\dfrac{\kappa\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}$ 00 $-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $2H+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{6H}$
$\dfrac{-\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ 00 $\dfrac{6\varepsilon_{\scriptscriptstyle(0)}H(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ $\dfrac{\tfrac{3}{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$
In this appendix, the perturbation equations (60) are derived. Prior to undertaking the subsequent analysis, it is necessary to carry out two preliminary steps.
Firstly, as $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$, (4b), it is unnecessary to consider the gauge-dependent quantity $\theta_{{\scriptscriptstyle(1)}}$. As a result, the system (37) can be rewritten in a more suitable form. By eliminating the variable $\theta_{\scriptscriptstyle(1)}$ from the differential equations  (37b)–(37e) using the algebraic constraint equation (37a), a system of four first-order ordinary differential equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})$
$+\varepsilon_{\scriptscriptstyle(0)}(1+w)\Bigl{[}\vartheta_{\scriptscriptstyle(1)}+\frac{1}{2H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)\Bigr{]}=0,$ (123a)
$\dot{n}_{\scriptscriptstyle(1)}$ $+3Hn_{\scriptscriptstyle(1)}+n_{\scriptscriptstyle(0)}\Bigl{[}\vartheta_{\scriptscriptstyle(1)}+\frac{1}{2H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)\Bigr{]}=0,$ (123b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}$ $+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}p_{\scriptscriptstyle(1)}}{a^{2}}=0,$ (123c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$ $+2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$
$-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}+\frac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)=0.$ (123d)
Secondly, the energy density and particle number density perturbations (4a) are expressed in the four quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Employing the background equations (10) to eliminate all time derivatives and the linearized constraint equation (37a) to eliminate $\theta_{\scriptscriptstyle(1)}$, the following results are obtained:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\dfrac{\varepsilon_{\scriptscriptstyle(1)}{{}^{3}\!}R_{\scriptscriptstyle(0)}-3\varepsilon_{\scriptscriptstyle(0)}(1+w)(2H\vartheta_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (124a)
$n^{\textnormal{phys}}_{\scriptscriptstyle(1)}=n_{\scriptscriptstyle(1)}-\dfrac{3n_{\scriptscriptstyle(0)}(\kappa\varepsilon_{\scriptscriptstyle(1)}+2H\vartheta_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}.$ (124b)
From an algebraic standpoint, it is more straightforward to begin by deriving equations for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. This will be accomplished in Sec. A.1. Subsequently, Eqs. (60) for the fluctuations (59) will be derived in Sec. A.2. The expression (44) will be utilized to eliminate the time derivative of $w$.
A.1 Evolution Equation for Energy Density Perturbations
The system (123) and the expression (124a) are reformulated using (13a) in a form suitable for implementation in a computer algebra program:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+\alpha_{11}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{12}n_{\scriptscriptstyle(1)}+\alpha_{13}\vartheta_{\scriptscriptstyle(1)}+\alpha_{14}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125a)
$\dot{n}_{\scriptscriptstyle(1)}+\alpha_{21}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{22}n_{\scriptscriptstyle(1)}+\alpha_{23}\vartheta_{\scriptscriptstyle(1)}+\alpha_{24}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+\alpha_{31}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{32}n_{\scriptscriptstyle(1)}+\alpha_{33}\vartheta_{\scriptscriptstyle(1)}+\alpha_{34}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+\alpha_{41}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{42}n_{\scriptscriptstyle(1)}+\alpha_{43}\vartheta_{\scriptscriptstyle(1)}+\alpha_{44}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\alpha_{51}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{52}n_{\scriptscriptstyle(1)}+\alpha_{53}\vartheta_{\scriptscriptstyle(1)}+\alpha_{54}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0.\!$ (125e)
The coefficients $\alpha_{ij}$ are given in Table 3.
The quantities in question (124) do not include the gauge function $\psi(\bm{x})$. Consequently, the gauge modes (147) will naturally disappear as a result of deriving the evolution equations for $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$. This process will be carried out in three steps.
Step 1.
The first step is to remove the explicit occurrence of ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ from Eqs. (125). By differentiating Eq. (125e) with respect to time and eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{{\scriptscriptstyle(1)}}$, $\vartheta_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ with the help of Eqs. (125a)–(125d), one obtains the following equation:
$\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+p_{1}\varepsilon_{\scriptscriptstyle(1)}+p_{2}n_{\scriptscriptstyle(1)}+p_{3}\vartheta_{\scriptscriptstyle(1)}+p_{4}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=0,$ (126)
where the coefficients $p_{1},\ldots,p_{4}$ are given by
$p_{i}=\dot{\alpha}_{5i}-\alpha_{51}\alpha_{1i}-\alpha_{52}\alpha_{2i}-\alpha_{53}\alpha_{3i}-\alpha_{54}\alpha_{4i}.$ (127)
As can be seen from Eq. (126), it follows that
${{}^{3}\!}R_{\scriptscriptstyle(1)}=-\dfrac{1}{p_{4}}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{p_{1}}{p_{4}}\varepsilon_{\scriptscriptstyle(1)}-\dfrac{p_{2}}{p_{4}}n_{\scriptscriptstyle(1)}-\dfrac{p_{3}}{p_{4}}\vartheta_{\scriptscriptstyle(1)}.$ (128)
In this manner, ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ is expressed as a linear combination of the quantities $\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\vartheta_{\scriptscriptstyle(1)}$. Upon replacing ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ in Eqs. (125) by the right-hand side of (128), the following system of equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+q_{1}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{11}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{12}n_{\scriptscriptstyle(1)}+\gamma_{13}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129a)
$\dot{n}_{\scriptscriptstyle(1)}+q_{2}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{21}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{22}n_{\scriptscriptstyle(1)}+\gamma_{23}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+q_{3}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{31}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{32}n_{\scriptscriptstyle(1)}+\gamma_{33}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+q_{4}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\gamma_{41}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{42}n_{\scriptscriptstyle(1)}+\gamma_{43}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+q_{5}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\gamma_{51}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{52}n_{\scriptscriptstyle(1)}+\gamma_{53}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129e)
where the coefficients $q_{i}$ and $\gamma_{ij}$ are given by
$q_{i}=-\dfrac{\alpha_{i4}}{p_{4}},\quad\gamma_{ij}=\alpha_{ij}+q_{i}p_{j}.$ (130)
It has been achieved that ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occurs explicitly only in Eq. (129d), whereas ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occurs implicitly in the remaining equations. Therefore, Eq. (129d) is no longer required. Accordingly, the remaining four ordinary differential equations are as follows: (129a)–(129c) and (129e) for the four unknown quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
Step 2.
In a manner analogous to the approach undertaken in Step 1, the explicit appearance of $\vartheta_{{\scriptscriptstyle(1)}}$ is removed from the system of equations (129). Differentiating Eq. (129e) with respect to time and subsequently eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{\scriptscriptstyle(1)}$ and $\vartheta_{\scriptscriptstyle(1)}$ with the help of Eqs. (129a)–(129c), results in the following equation:
$q_{5}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+r\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+s_{1}\varepsilon_{\scriptscriptstyle(1)}+s_{2}n_{\scriptscriptstyle(1)}+s_{3}\vartheta_{\scriptscriptstyle(1)}=0,$ (131)
where the coefficients $r$ and $s_{i}$ are given by
$s_{i}$ $=\dot{\gamma}_{5i}-\gamma_{51}\gamma_{1i}-\gamma_{52}\gamma_{2i}-\gamma_{53}\gamma_{3i},$ (132a)
$r$ $=1+\dot{q}_{5}-\gamma_{51}q_{1}-\gamma_{52}q_{2}-\gamma_{53}q_{3}.$ (132b)
As can be seen from Eq. (131), it follows that
$\vartheta{\scriptscriptstyle(1)}=-\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{r}{s_{3}}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{s_{1}}{s_{3}}\varepsilon_{\scriptscriptstyle(1)}-\dfrac{s_{2}}{s_{3}}n_{\scriptscriptstyle(1)}.$ (133)
In this manner, $\vartheta_{\scriptscriptstyle(1)}$ is expressed as a linear combination of $\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$. Upon replacing $\vartheta_{\scriptscriptstyle(1)}$ in Eqs. (129) by the right-hand side of (133), the follwing system of equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{13}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{1}-\gamma_{13}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{11}-\gamma_{13}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{12}-\gamma_{13}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134a)
$\dot{n}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{23}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{2}-\gamma_{23}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{21}-\gamma_{23}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{22}-\gamma_{23}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{33}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{3}-\gamma_{33}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{31}-\gamma_{33}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{32}-\gamma_{33}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{43}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{4}-\gamma_{43}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{41}-\gamma_{43}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{42}-\gamma_{43}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{53}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{5}-\gamma_{53}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{51}-\gamma_{53}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{52}-\gamma_{53}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0.$ (134e)
It has been achieved that the quantities $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occur explicitly only in Eqs. (134c) and (134d), whereas they occur implicitly in the remaining equations. Consequently, Eqs. (134c) and (134d) are no longer required. The remaining equations (134a), (134b) and (134e) are three ordinary differential equations for the three unknown quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
Step 3.
The subsequent steps would be to eliminate, in sequence, $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$ from Eq. (134e) with the aid of Eqs. (134a) and (134b). This would result in a fourth-order differential equation for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. It must be noted, however, that this is not a possibility, given that the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$ do not occur explicitly in Eq. (134e). To demonstrate this, Eq. (134e) will be rewritten in the subsequent form:
$\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{1}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{2}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=a_{3}\left(n_{\scriptscriptstyle(1)}+\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}\varepsilon_{\scriptscriptstyle(1)}\right),$ (135)
where the coefficients $a_{1}$, $a_{2}$ and $a_{3}$ are given by
$a_{1}=-\dfrac{s_{3}}{\gamma_{53}}+\dfrac{r}{q_{5}},\quad a_{2}=-\dfrac{s_{3}}{\gamma_{53}q_{5}},\quad a_{3}=\dfrac{\gamma_{52}s_{3}}{\gamma_{53}q_{5}}-\dfrac{s_{2}}{q_{5}}.$ (136)
As a consequence of the gauge-invariance of the quantity $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, the left-hand side of Eq. (135) is also gauge-invariant. It thus follows that the right-hand side of this equation is also gauge-invariant, as will be demonstrated subsequently. Indeed, the result is as follows:
$\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}=-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}.$ (137)
It can be seen that the conservation laws (10c) and (10d) and the expressions (4a) lead to the following conclusion:
$n_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (138)
It follows from expressions (137) and (138) that the right-hand side of Eq. (135) does not explicitly contain the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$. Consequently, it is gauge-invariant. Therefore, equation (135) governs the evolution of $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
The proof of the equality (137) is straightforward but requires significant computational effort. The algebraic task was performed using the computer algebra system Maxima [44].
The final step in this process will be deriving an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. The linearized equations (123), from which the evolution equations are derived, are of fourth order. From this system, a second-order equation (135) for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ has been derived. It thus follows that the remaining system from which an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived is at most of second order. As the gauge-invariant quantities, namely, $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, have been employed, one degree of freedom, specifically the gauge function $\psi(\bm{x})$ in (147), has been eliminated. Consequently, only a first-order evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived. In place of deriving an equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, an evolution equation will be derived for the expression (138), which contains $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. By differentiating the left-hand side of equation (138) with respect to time and employing the background equations (10c) and (10d), the first-order equations (123a) and (123b) and the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$, the following result is obtained:
$\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left(n_{\scriptscriptstyle(1)}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}\right)$
$=-3H\left(1-\frac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)\left(n_{\scriptscriptstyle(1)}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}\right).$ (139)
The algebraic task was performed using the Maxima [44] computer algebra software. Due to the equality in Eq. (138), it is possible to replace $n_{\scriptscriptstyle(1)}$ and $\varepsilon_{\scriptscriptstyle(1)}$ in Eq. (A.1) by $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
A.2 Evolution Equations for Density Fluctuations
First Eq. (60a) will be derived. Upon substituting the expression $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}$ in Eq. (135) and dividing by $\varepsilon_{\scriptscriptstyle(0)}$, the result is
$b_{1}=2\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1},\quad b_{2}=\dfrac{\ddot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1}\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{2},\quad b_{3}=a_{3}\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}.$ (140)
The coefficients $a_{1}$, $a_{2}$ and $a_{3}$ are given by (136). With the aid of the computational software Maxima [44], the coefficients (61) of the main text have been calculated.
Finally, Eq. (60b) will be derived. From the definitions (59), it can be deduced that
$n_{\scriptscriptstyle(1)}^{\textnormal{phys}}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}=n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (141)
By employing the equality (138) and substituting the expression (141) into Eq. (A.1), the following is obtained:
$\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right]$
$=-3H\left(1-\frac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right].$ (142)
Derivation of Eq. (60b) of the main text was accomplished through the utilization of Eq. (10d).
Appendix B Verifying the Evolution Equations for Scalar Perturbations
The present paper does not provide a derivation of the linearized equations (12). A demonstration of the validity of the system of equations (37), which is derived from the aforementioned linearized equations and is fundamental to the new perturbation theory, will therefore be given via three different methods.
The initial step in this process is to utilize the well-established property of the Einstein equations and conservation laws. This implies that if the constraint equations are satisfied at an initial time, then the solutions of the dynamical equations and conservation laws will satisfy the constraint equations for all subsequent times. A more detailed discussion can be found in Weinberg’s book [31], in Sec. 7.5 on the Cauchy problem. This is also true in reverse: if solutions of the constraint equations and conservation laws can be found for all times, they must necessarily satisfy the dynamical equations. The system (37) comprises the conservation laws and constraint equations, and consists of four first-order ordinary differential equations and one algebraic equation. It thus follows that the system (37), can be solved without the use of the dynamical equations, since the latter are inherently satisfied by the solution of the aforementioned system. Accordingly, the latter of the two observations will be employed to substantiate the validity of equations (37).
The off-diagonal ($i\neq j$) equations of (12c) are not associated with scalar perturbations. For the diagonal ($i=j$) equations, it is sufficient to consider the contraction of the dynamical equations (12c):
$\ddot{h}^{k}{}_{k}+6H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-3\kappa(\varepsilon_{{\scriptscriptstyle(1)}}-p_{{\scriptscriptstyle(1)}}).$ (143)
Eliminating the quantity $\dot{h}^{k}{}_{k}$ with the help of (19), yields the following result:
$(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})+6H(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})-{{}^{3}\!}R_{\scriptscriptstyle(1)}=\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}).$ (144)
Using the energy density constraint equation (37a) to eliminate the second term of (144), results in the following equation:
$\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)}).$ (145)
This equation is identical to the time derivative of the energy density constraint equation (37a), as will now be demonstrated. Differentiation of (37a) with respect to time yields
$2\dot{H}(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})+2H(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})=\tfrac{1}{2}\,{{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+\kappa\dot{\varepsilon}_{\scriptscriptstyle(1)}.$ (146)
Eliminating the time derivatives $\dot{H}$, ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$, and $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ with the help of (10a)–(10c), the momentum constraint equation (37b) and the energy density conservation law (37c), respectively, yields the dynamical equation (145). Consequently, Eqs. (37a)–(37c) are correct.
As indicated in the aforementioned verification, the correctness of the equations (37d) and (37e) has not yet been verified. This will now be addressed by employing the gauge modes. It is observed that Eqs. (37) remain invariant under the gauge transformations (1) with the gauge functions $\xi^{\mu}(t,\bm{x})$ given by (6). Furthermore, the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\theta_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ are known to be gauge-dependent. It is therefore evident that the gauge modes
$\hat{S}_{\scriptscriptstyle(1)}$ $=\psi\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,\theta,p,$ (147a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}$ =~$=-\frac{\tilde{\nabla}^{2}\psi}{a^{2}},$ (147b)
${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H\left(\frac{\tilde{\nabla}^{2}\psi}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\psi\right),$ (147c)
must be solutions to Eqs. (37). The gauge modes (147a) are derived directly from (3), with the understanding that $\xi^{0}=\psi$. The gauge mode (147b) is the covariant divergence of (28). The gauge mode (147c) is derived by substituting the gauge modes (147a)–(147b) into the perturbed Friedmann equation (37a):
${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H(\hat{\theta}_{{\scriptscriptstyle(1)}}-\hat{\vartheta}_{{\scriptscriptstyle(1)}})-2\kappa\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$
$=4H\left(\psi\dot{\theta}_{{\scriptscriptstyle(0)}}+\frac{\tilde{\nabla}^{2}\psi}{a^{2}}\right)-2\kappa\psi\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (148)
Using that $\theta_{{\scriptscriptstyle(0)}}=3H$, the Friedmann equation (10a), and Eq. (10b), one arrives at the gauge mode (147c). It will now be demonstrated that the gauge modes given by (147) are indeed solutions of Eqs. (37b)–(37e). In order to execute this task, the first step entails deriving the time derivatives of the gauge modes:
$\dot{\hat{\vartheta}}_{{\scriptscriptstyle(1)}}$ $=-2H\hat{\vartheta}_{{\scriptscriptstyle(1)}},$ (149a)
${{}^{3}\!}\dot{\hat{R}}_{{\scriptscriptstyle(1)}}$ $=-4(\dot{H}-2H^{2})\bigl{[}\hat{\vartheta}_{{\scriptscriptstyle(1)}}+\tfrac{1}{2}\,{{}^{3}\!}R_{{\scriptscriptstyle(0)}}\psi\bigr{]},$ (149b)
$\dot{\hat{\varepsilon}}_{{\scriptscriptstyle(1)}}$ $=\psi\ddot{\varepsilon}_{{\scriptscriptstyle(0)}}=-3\varepsilon_{{\scriptscriptstyle(0)}}(1+w)\bigl{[}\dot{H}-3H^{2}(1+\beta^{2})\bigr{]}\psi,$ (149c)
$\dot{\hat{n}}_{{\scriptscriptstyle(1)}}$ $=\psi\ddot{n}_{{\scriptscriptstyle(0)}}=-3n_{{\scriptscriptstyle(0)}}\bigl{[}\dot{H}-3H^{2}\bigr{]}\psi,$ (149d)
where Eqs. (10b)–(10d) and Eq. (44) have been used. Subsequent to the aforementioned preparation, the verification of the accuracy of the equations (37b)–(37e) can be undertaken. First, by substituting the gauge modes, namely $\hat{\theta}_{{\scriptscriptstyle(1)}}$, $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}\hat{R}_{{\scriptscriptstyle(1)}}$ into Eq. (37b), the result is an equation from which the variable $\psi$ disappears. Next, dividing the resulting equation by $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$ and multiplying by $\tfrac{3}{2}H$, one arrives at the time derivative of Eq. (10a). Second, the substitution of the gauge modes $\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$, $\hat{\theta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37c) results in the elimination of the term with $\dot{H}$. Next, dividing, the resulting equation by $\psi$ and employing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ followed by utilizing Eq. (10c) yields an identity. Third, by substituting the gauge modes $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37d), and utilizing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and subsequently using Eq. (10c) one arrives at an identity. Finally, upon substituting the gauge modes $\hat{n}_{{\scriptscriptstyle(1)}}$ and $\hat{\theta}_{{\scriptscriptstyle(1)}}$ into Eq. (37e) the term with $\dot{H}$ vanishes. Using Eq. (10d), one arrives at an identity. It has been demonstrated that the gauge modes (147) satisfy Eqs. (37b)–(37e). Consequently, it can be concluded that these equations are accurate. Given the established validity of Eq. (37a), it can be deduced that the system (37) is also accurate.
As a final check, it is observed in Appendix C that for a flat flrw universe, the evolution equations (152) for the gauge-dependent density fluctuation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ are obtained from Eqs. (37). Of particular note is the fact that the well-known evolution equations for $\delta$ are recovered in the cases of a radiation-dominated universe (154) and a universe after decoupling of matter and radiation (156).
Consequently, it can be concluded that the system of equations (37) has been verified in three distinct ways.
Appendix C Conventional Evolution Equation for Cosmological Density Fluctuations
The usual evolution equation for density perturbations in a flat, ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}=0$, flrw universe is given by
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\frac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)(1+3w)\right)\delta=0,$ (150)
where $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$. The derivation of this equation has four primary objectives. Firstly, it demonstrates the correctness of the system (37), at least in the context of a flat flrw universe. Secondly, it is demonstrated that the standard equation (150) is incomplete. That is, its solution is a linear combination of only one physical mode and one gauge mode, independent of the scale of a fluctuation. Consequently, Eq. (150) does not give an adequate description of cosmological density perturbations. Thirdly, it is shown that the novel perturbation theory (60) provides all well-known physical solutions for large-scale perturbations in a flat flrw universe. Specifically, it produces the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe after decoupling of matter and radiation. Finally, it will be demonstrated that Eq. (150), when derived from the Newtonian theory of gravitation, yields solutions that remain dependent on the selection of the coordinates.
In order to derive equation (150), it is necessary to utilize the background equations (71) and the equations (123) for scalar perturbations. These two systems of equations are complete, since they contain all conservation laws and constraint equations. In this appendix, it is assumed that $c=1$.
Equation (150) has been derived from an equation of state $p=p(\varepsilon)$ in the existing literature. Therefore, the particle number conservation law (71c) and its linearized counterpart (123b) are not required. Substituting $p_{n}=0$ into (13) and (40) yields $p_{{\scriptscriptstyle(1)}}=p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}}=\beta^{2}\varepsilon_{{\scriptscriptstyle(1)}}$. According to the definition $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$, Eqs. (123) can be rewritten as follows:
$\dot{\delta}+3H\delta\bigl{(}\beta^{2}+\tfrac{1}{2}(1-w)\bigr{)}+(1+w)\left(\vartheta_{\scriptscriptstyle(1)}+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(1)}}{4H}\right)=0,$ (151a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\frac{\beta^{2}}{1+w}\dfrac{\nabla^{2}\delta}{a^{2}}=0,$ (151b)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}=0,$ (151c)
where Eq. (71b) has been used.
In order to arrive at a generalization of (150), the following steps must be taken. First, it is necessary to differentiate equation (151a) with respect to time. Subsequently, the time derivatives of $w$, $H$, $\varepsilon_{\scriptscriptstyle(0)}$, $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ must be eliminated. This can be achieved by utilizing the equations (44), (71a) and (71b) and the perturbation equations (151b) and (151c). The final step in the process is the elimination of ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ through the application of (151a). This results in the following system of equations:
$\ddot{\delta}+2H\dot{\delta}\bigl{(}1+3\beta^{2}-3w\bigr{)}$
$\qquad-\biggl{[}\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\Bigl{(}(1+w)(1+3w)\Bigr{.}\Bigr{.}$
+ $\qquad+\,4w-6w^{2}+12\beta^{2}w-4\beta^{2}-6\beta^{4}\Bigr{)}-6\beta\dot{\beta}H\biggr{]}\delta$
$\quad=-3H\beta^{2}(1+w)\vartheta_{\scriptscriptstyle(1)},$ (152a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\dfrac{\beta^{2}}{1+w}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (152b)
In this procedure the computer algebra program Maxima [44] has been used.
The gauge modes (147) are for the system (152) given by
$\hat{\delta}(t,\bm{x}):=\dfrac{\psi(\bm{x})\dot{\varepsilon}_{\scriptscriptstyle(0)}(t)}{\varepsilon_{\scriptscriptstyle(0)}(t)}=-3H(t)\psi(\bm{x})\bigl{(}1+w(t)\bigr{)},$ (153a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}(t)}.$ (153b)
The fact that the gauge modes (153) satisfy the system (152) has been shown in Appendix B.
It can be inferred from the set of equations (152) and the gauge modes (153) that:
  • The solution of the homogeneous part of Eq. (152a) is a linear combination of a physical mode and the gauge mode (153a). The particular solution is obtained by solving Eq. (152b), and contains the gauge mode (153b).
  • It is impossible to determine the arbitrary infinitesimal gauge function $\psi(\bm{x})$ through physical means. Therefore, the imposition of physical initial values on the system (152) does not yield a physical solution.
  • In the small-scale limit, the gauge modes (153) do not vanish. This means that in the small-scale limit gauge-dependent quantities do not become gauge-invariant. Consequently, $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ does not become equal to its Newtonian counterpart $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, given by (56) and (58). This is consistent with the gauge transformation (53) of the Newtonian theory of gravitation.
  • A notable distinction between the conventional equation (150) and Eq. (152a) is that the former is homogeneous, whereas Eq. (152a) is inhomogeneous. This is because the system (152) is derived from the complete set of constraint equations and conservation laws. In the case where the right-hand side of Eq. (152a) is absent, as is the case in Eq. (150), then (153a) is a gauge mode if $\nabla^{2}\psi=0$. Consequently, $\psi=C$ is an arbitrary infinitesimal constant, so that the gauge mode (153a) does not vanish.
  • In the universe following decoupling, the values of $w\ll 1$ and $\beta^{2}\ll 1$ are negligible, allowing for the subsequent neglect of $w$ in relation to the constants of order one in Eq. (150). This is equivalent to deriving the usual equation (150) from the Newtonian theory of gravitation adapted to the expansion of the universe. See, for example, Ref. [10], Sec. 6.2, Ref. [31], Sec. 15.9, and Ref. [45], Eq. (3.17). However, the gauge transformation (53) in the nonrelativistic limit ensures that the gauge mode (153a) does not vanish.
From these points it is clear that the conventional equation (150) is unsuitable to study the evolution of density perturbations in the universe. Given that Eq. (150) can be derived from Newtonian gravitation for $w\ll 1$, it follows that the Newtonian theory is unable to provide an adequate description of density perturbations in the universe.
In the ensuing two sections, a comparison is presented between Eqs. (79) and (95), with the system of equations (152). This comparison will be conducted in the case of large-scale fluctuations. It will become evident that the gauge modes preclude the attainment of physical solutions.
C.1 Radiation-Dominated Era
In this era, the pressure is given by a linear barotropic equation of state $p=\tfrac{1}{3}\varepsilon$, so that $\beta^{2}=\tfrac{1}{3}$ and ${w=\tfrac{1}{3}}$. In this case Eqs. (152) result in the system
$\ddot{\delta}+2H\dot{\delta}-\left(\dfrac{1}{3}\frac{\nabla^{2}}{a^{2}}+\tfrac{4}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=-\tfrac{4}{3}H\vartheta_{\scriptscriptstyle(1)},$ (154a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H\vartheta_{\scriptscriptstyle(1)}+\dfrac{1}{4}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (154b)
The gauge modes (153) are solutions of the system (154) for $w=\tfrac{1}{3}$, as can be verified by substitution.
For large-scale fluctuations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution of the system (154) is
$\delta(t,\bm{x})=(c_{1}t-2\psi(\bm{x})t^{-1})+\tfrac{9}{8}t^{1/2},$ (155a)
$\vartheta_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}}+\tfrac{9}{8}t^{-1/2},$ (155b)
where it is used that $H=1/2t$ and $\kappa\varepsilon_{\scriptscriptstyle(0)}=3/4t^{2}$. The expression between parentheses in (155a) is the solution of the homogeneous part of Eq. (154a). The particular solution $\tfrac{9}{8}t^{1/2}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\tfrac{9}{8}t^{-1/2}$. The solution (85) derived from the novel perturbation equation (82) is not derivable from the standard equations (154) due to the presence of the gauge function $\psi$ in the solution (155a).
Since $\nabla^{2}\delta^{\textnormal{phys}}$ could have been large for small-scale fluctuations, it may have had a large influence on $\vartheta^{{\textnormal{phys}}}_{\scriptscriptstyle(1)}$ and this could have, in turn, a major impact on the evolution of $\delta^{\textnormal{phys}}$. That is why (79a) yields oscillating density fluctuations (84) with an increasing amplitude, instead of a constant amplitude which follows from (150).
C.2 Era after Decoupling of Matter and Radiation
In this era, the equation of state for the pressure is according to thermodynamics given by (88). Therefore, (91) and (92) imply that $w\approx\tfrac{3}{5}\beta^{2}\ll 1$, so that with $T_{\scriptscriptstyle(0)}\propto a^{-2}$ one obtains $\dot{\beta}/\beta=-H$. Using Eq. (71a) one can derive the following result: $6\beta\dot{\beta}H=-2\kappa\varepsilon_{\scriptscriptstyle(0)}\beta^{2}$. Upon substituting the latter expression into (152a) and neglecting $w$ and $\beta^{2}$ with respect to the constants of order one, the system (152) results in
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=-3H\beta^{2}\vartheta_{\scriptscriptstyle(1)},$ (156a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+2H\vartheta_{\scriptscriptstyle(1)}+\beta^{2}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (156b)
The gauge modes (153) are solutions of the system (156) for $w\ll 1$ and $\nabla^{2}\psi=0$, as can be verified by substitution. Consequently, for the system (156) $\psi$ is an arbitrary infinitesimal constant $C$. This implies that $\vartheta_{\scriptscriptstyle(1)}=\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is a physical quantity, since its gauge mode $\hat{\vartheta}_{\scriptscriptstyle(1)}$, (153), vanishes identically. Consequently, for large-scale fluctuations the solution of the system (156) is
$\delta(t)=(c_{1}t^{2/3}-2Ct^{-1})+\tfrac{7}{9}t^{-5/3},$ (157a)
$\vartheta_{\scriptscriptstyle(1)}^{\textnormal{phys}}(t)=-\tfrac{7}{9}t^{-4/3},$ (157b)
where $H=2/3t$, $\kappa\varepsilon_{\scriptscriptstyle(0)}=4/3t^{2}$, and $\beta^{2}\propto t^{-4/3}$ has been used. The latter proportionality follows from (91) and (92), $T_{\scriptscriptstyle(0)}\propto a^{-2}$ and (104). The expression between parentheses in (157a) is the solution of the homogeneous part of Eq. (156a). The particular solution $\tfrac{7}{9}t^{-5/3}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\tfrac{7}{9}t^{-4/3}$. The solution (109) derived from the novel perturbation equation (108) is not derivable from the standard equations (156) due to the presence of the gauge constant $C$ in the solution (157a).
Since $\psi=C$, the gauge transformation (1) with $\xi^{\mu}(t,\bm{x})$ given by (6) reduces to the Newtonian gauge transformation (53). This is to be expected, since a cosmological fluid for which ${w\ll 1}$ and $\beta^{2}\ll 1$ can be described by nonrelativistic equations of state (88).

### v3

@@
-The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe
+Earliest Structures in the Universe can be explained by a Relativistic Cosmological Perturbation Theory
 Abstract
-This paper presents a novel approach to four fundamental problems in the field of cosmological perturbation theory.
-Firstly, the issue of gauge dependence has been addressed by demonstrating the existence of unique and gauge-invariant quantities corresponding to the actual perturbations.
-Secondly, the formation of primordial structures after decoupling of matter and radiation is dependent on the existence of local fluid flows resulting from local pressure gradients.
-To take pressure gradients into account, it is necessary to consider both the energy density and the particle number density.
-Thirdly, the novel relativistic perturbation theory applies to an open, flat, and closed Friedmann-Lemaître-Robertson-Walker universe.
-The derivation of the novel perturbation theory definitively reveals the inherent limitations of Newtonian gravitation as a framework for investigating cosmological density perturbations.
-Finally, the application of the perturbation theory to a flat universe demonstrates that, prior to decoupling, perturbations in both Cold Dark Matter and ordinary matter are coupled to perturbations in radiation.
-Therefore, the universe’s earliest structures formed only after decoupling, at which point local nonadiabatic random pressure fluctuations became a significant factor.
-Negative nonadiabatic pressure fluctuations resulted in a brief, rapid growth of density fluctuations until the total pressure fluctuations became positive.
-In contrast, positive nonadiabatic pressure fluctuations led to the formation of voids.
-Perturbations with masses about $2.2\times 10^{4}\,\text{M}_{\odot}$ became nonlinear already $60$ million years after the Big Bang, and perturbations with masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$ became nonlinear within about $600$ million years.
+A relativistic cosmological perturbation theory for the Friedmann-Lemaître-Robertson-Walker universe is presented that explains the masses and formation times of the first structures in our universe.
+First, it is shown that, without a coordinate system being used, quantities intended to represent energy density and particle number density perturbations can be defined in only one way.
+The Newtonian limit, where the pressure becomes zero, proves that these quantities are indeed the perturbations of the energy density and the particle number density.
+Then, after selecting a reference frame, a perturbation theory will be formulated based on these quantities.
+This formulation considers the local perturbation to the spatial curvature resulting from a density perturbation, the local fluid velocity due to pressure gradients caused by the self-gravity of the density perturbation, and entropy perturbations, all of which are necessary for structure formation.
+Pressure perturbations consist of two components: an adiabatic component caused by the density perturbation itself, and a random, nonadiabatic component resulting from the rapid, chaotic transition to the era when matter and radiation were decoupled.
+Immediately after decoupling, negative nonadiabatic pressure perturbations in various density perturbations enabled their rapid growth over a short period of time.
+This brief period ended when the total pressure perturbation became positive.
+Subsequently, the density perturbations gradually grew toward their nonlinear phase, which was reached early on.
 Cosmology; Early Universe; Structure Formation; Perturbation Theory; Gauge Problem.
@@
 I Introduction
-The linearized Einstein equations and conservation laws of Friedmann-Lemaître-Robertson-Walker (flrw) universes are crucial to cosmology.
-They can explain the formation of all kinds of structures in the expanding universe, including stars, galaxies, and microwave background fluctuations.
-Lifshitz [1] and Lifshitz & Khalatnikov [2] were the first researchers to develop a cosmological perturbation theory to study the evolution of density perturbations in the universe.
-They encountered the problem that the solutions of the linearized Einstein equations and conservation laws can be modified by linear coordinate transformations, which are known as gauge transformations.
-Consequently, the solutions are contingent upon the selection of a coordinate system, rendering them gauge-dependent and thus devoid of physical significance.
-This is the notorious gauge problem of cosmology, which demands urgent resolution.
-A multitude of potential solutions to this problem have been proposed in the literature.
-Stewart [3] noted in his paper, […] perturbation theory within general relativity is generally regarded as a somewhat arcane subject […] .
-This observation is further substantiated by the lack of progress observed to date in this field.
-The purpose of this paper is to elucidate the hitherto puzzling aspects of cosmological perturbation theory.
-Before delving into an examination of the approaches outlined in the extant literature, it is imperative to acknowledge the fundamental criteria that a comprehensive and precise cosmological perturbation theory must fulfill.
-Primarily, as Bardeen elucidated in his seminal paper [4], a quantity signifying the real physical energy density perturbation must be manifestly gauge-invariant, that is, it must be defined independently of any reference system.
-Secondly, in the nonrelativistic limit, the linearized energy density constraint equation, when considered together with the linearized momentum constraint equation, must yield the time-independent Poisson equation of Newtonian theory with the aforementioned gauge-invariant quantity as the source term.
-Thirdly, in the nonrelativistic limit, it is essential that the freedom to select a coordinate system is present.
-That is to say, relativistic gauge transformations must converge to Newtonian gauge transformations in the nonrelativistic limit.
-Finally, and equally important, the perturbation theory must encompass all previously established and well-known physical solutions for large-scale density perturbations.
-The present paper proposes a novel cosmological perturbation theory that has been demonstrated to satisfy all four of the aforementioned criteria.
-A review of the existing literature reveals that there is no theory that meets all four of the above criteria.
-Bardeen [4] was the first to demonstrate that the linearized Einstein equations and conservation laws can be recast into evolution equations that have solutions independent of the choice of a coordinate system by using gauge-invariant quantities.
-In his paper, Bardeen constructed two distinct gauge-invariant quantities for the perturbation to the energy density, designated as $\epsilon_{m}$ and $\epsilon_{g}$.
-In the limit of small scales, these quantities converge to the usual gauge-dependent quantity, denoted as $\varepsilon_{{\scriptscriptstyle(1)}}$.
-Bardeen presupposed that on small scales, Newtonian gravitation is a valid description of cosmological density perturbations, and therefore, the gauge-dependent quantity, $\varepsilon_{{\scriptscriptstyle(1)}}$, becomes equal to its Newtonian counterpart.
-It is, however, a well-known fact that in Newtonian gravitation the coordinates of time and space can be chosen at will.
-Correspondingly, this must also be true in the nonrelativistic limit of a cosmological perturbation theory.
-It can thus be concluded that $\varepsilon_{{\scriptscriptstyle(1)}}$, which depends on the choice of coordinates in general relativity, is also dependent on the choice of coordinates in Newtonian gravitation.
-Thus, quantities that lack physical significance in general relativity are similarly devoid of physical meaning in Newtonian gravitation.
-It follows that the small-scale limit is not equivalent to the nonrelativistic limit.
-In consequence, the quantities $\epsilon_{m}$ and $\epsilon_{g}$ do not correspond to the actual density perturbations.
-Bardeen’s paper has inspired a number of works, including those in Refs.
-[5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
-These researchers proposed alternative perturbation theories using gauge-invariant quantities that differ from those used by Bardeen.
-The authors of Refs.
-[5, 6, 7, 9, 11] define gauge-invariant density perturbations by gradients of the energy density.
-For a pressureless fluid, their equations are identical to the standard perturbation equation derived from Newton’s theory of gravitation.
-This equation yields the well-known solutions of the standard equation (150) with $w=0$ and $\beta=0$, including the gauge mode.
-In addition to the fact that the solution is gauge-dependent, even in the nonrelativistic limit, the evolution of density perturbations in a pressureless fluid is precluded by the absence of pressure gradients.
-This is a consequence of the linearized momentum conservation laws.
-It can thus be concluded that the standard equation (150) with $w=0$ and $\beta=0$ cannot be taken to represent the nonrelativistic limit of a cosmological perturbation theory.
-Moreover, as indicated in Ref.
-[6], the constraint equations have not been integrated into the analysis.
-It is of crucial importance to acknowledge that the constraint equations represent relations between initial values, thereby playing a pivotal role in the analysis.
-In the absence of the constraint equations, the solutions to the system of dynamical equations and conservation laws lack physical significance.
-Furthermore, it is not possible to achieve a correct Newtonian limit without the constraint equations (for further details, see Sec.
-IX).
-In this paper, the complete set of Einstein equations and conservation laws is employed to develop a novel perturbation theory.
-In order to interpret the gauge-invariant definition of the density perturbation $\overline{\delta\varepsilon}$, as proposed by Mukhanov et al.
-[8, 10], it is necessary to set the Hubble function $\cal{H}$ to zero in Eq.
-(7.47) in Ref.
-[10].
-Nevertheless, this yields a Poisson-like equation that differs from the Newtonian Poisson equation, as the potential is time-dependent in the Poisson-like equation, as evidenced by equations (7.48) and (7.49) in Ref.
-[10].
-Thus, $\overline{\delta\varepsilon}$ does not represent the actual physical density perturbation.
-Furthermore, the Hubble function cannot be set to zero, as this would violate the Friedmann equation (10a), which is inconsistent with the assumption that the universe is not empty.
-The present paper introduces a novel perturbation theory that achieves the time-independent Poisson equation and thus the correct nonrelativistic limit by considering zero pressure alone.
-This obviates the need to set the Hubble function to zero.
-As has been stated in Refs.
-[12, 13], the linearized Einstein equations and conservation laws for density perturbations yield the usual evolution equation (150) derived from Newtonian gravitation when the pressure is zero.
-The current study will demonstrate that in a pressureless fluid, the linearized Einstein equations and conservation laws are reduced to a single equation: the time-independent Poisson equation of Newtonian gravitation.
-The authors of Refs.
-[14, 15] present a method for the construction of gauge-invariant quantities in the synchronous reference system.
-However, the definition of gauge invariance presented in these works does not align with the conventional understanding of what constitutes a gauge-invariant quantity.
-As defined in Refs.
-[4, 3], a gauge-invariant quantity is independent of the choice of a coordinate system.
-Consequently, the definition of gauge-invariant quantities is formulated in this paper without the utilization of a coordinate system.
-In their study of the evolution of density perturbations in a flat flrw universe in its radiation-dominated phase, Ma & Bertschinger [16] employ two distinct coordinate systems: the synchronous ($g_{0i}=0$) gauge and the conformal Newtonian gauge.
-The latter gauge is, in fact, a restricted synchronous gauge.
-Given the existence of gauge transformations between different synchronous reference systems, the solutions are dependent on the selected synchronous coordinates.
-Consequently, gauge modes emerge in the solutions, rendering them devoid of physical significance even when physical initial values are imposed.
-Their findings pertaining to the density fluctuation, represented by $\delta$, deviate between the synchronous and conformal Newtonian coordinates.
-In addition, Eq.
-(29) in Ref.
-[16] is flawed in two respects.
-Firstly, the factor $(1-3w)$ in the expansion term should, in fact, be $(2-3w)$, which is critical in the context of a radiation-dominated universe, where $w=\tfrac{1}{3}$.
-Secondly, it is crucial to replace the expression $k^{2}$ with $k^{2}/a^{2}$.
-This substitution is necessary because the background three-space metric of a flat flrw universe is $a^{2}\delta_{ij}$ rather than $\delta_{ij}$.
-This can be demonstrated by considering the line element (1) in Ref.
-[16], which illustrates that three-space is expanding.
-Once these corrections are made, the resulting equation is equal to the covariant divergence of Eq.
-(43) in this paper.
-The latter equation is correct, as demonstrated in Appendices B and C.
-The aforementioned treatments resulted in a plethora of additional contributions to this field.
-See, e.g., Refs.
-[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
-In a review, Ellis [28] discusses the work of Lifshitz and Khalatnikov and other research on the subject.
-A review of the literature reveals a lack of consensus concerning the accuracy of the quantities employed to represent physical density perturbations.
-This paper addresses the issue of gauge dependence in the context of cosmology.
-It demonstrates the existence of unique gauge-invariant quantities that correspond to the true physical density perturbations.
-Subsequently, it presents equations for the study of the evolution of density perturbations in universes with an open, flat, or closed flrw geometry.
-In light of the general covariance of general relativity, the evolution equations for gauge-invariant quantities are mathematically identical in all coordinate systems.
-Consequently, they yield the same solutions in any given coordinate system.
-The novel perturbation theory, when applied to a flat flrw universe, reveals that prior to the decoupling of radiation and matter, Cold Dark Matter (cdm) perturbations were coupled to the radiation energy density perturbations in such a way that they could not clump together.
-Consequently, cdm proved to be ineffective in facilitating the process of structure formation subsequent to the decoupling of radiation and matter.
-The formation of the first stars, designated as Population iii stars [29], can be explained by the rapid growth of small density perturbations over a brief period of time.
-This rapid growth arises from local nonadiabatic random pressure fluctuations in the early universe following the decoupling.
-The James Webb Space Telescope (jwst), which was launched on December 25, 2021, is designed primarily for near-infrared astronomy and may therefore provide insights regarding the existence and nature of the earliest stars.
-The data from the jwst [30] suggest that the galaxies observed in the early universe appear to have grown to a size greater than anticipated, forming massive and compact structures at a rate exceeding the expectations of the hierarchical $\Lambda\textsc{cdm}$ structure formation paradigm.
-This may indicate that the existence of cdm is not supported by the data.
-II A Novel Cosmological Perturbation Theory
-This paper presents a novel cosmological perturbation theory, one that differs substantially from the perturbation theories in the extant literature, including the treatment of the cosmic fluid.
-The aim of this section is to provide an overview of the novel approach and to elucidate the method by which the new theory is constructed.
-II.1 Cosmic Fluid
-It is well established that in the era following the decoupling of matter and radiation, the pressure in the cosmological fluid is negligible in comparison to the energy density.
-Therefore, it can be concluded that the pressure can be disregarded in the Einstein equations and conservation laws of the background, i.e., the unperturbed, flrw universe.
-In the context of cosmology, it is commonly assumed that in the perturbed universe, the pressure and, therefore, local pressure gradients can also be neglected.
-However, it is well known that in the absence of local pressure gradients, local fluid flows cannot occur.
-This is equally true in the case of the cosmic fluid.
-As will be demonstrated in Sec.
-IX, the linearized momentum conservation laws indicate that, in the absence of pressure, local fluid flows will not occur and density perturbations will be static.
-It is thus crucial to consider pressure gradients in order to examine the evolution of density perturbations.
-Given that, subsequent to decoupling, the cosmic fluid is to be considered a perfect gas where the pressure is dependent on the mean kinetic energy of its particles, it is essential that the kinetic energy density be incorporated into the analysis.
-Both the pressure and the kinetic energy density are dependent on the particle number density and temperature.
-Consequently, in addition to the energy density, it is necessary to include the particle number density and the temperature in the equation of state.
+Structures in our universe, such as stars, galaxies, and black holes, form through the gravitational growth of small density perturbations.
+Therefore, it is reasonable to expect that a cosmological perturbation theory based on general relativity — the most accurate theory of gravity to date — would explain the masses and formation times of the first structures.
+There are two main reasons why such a theory does not yet exist.
+First, the gauge problem of cosmology has not yet been solved.
+Second, particle number density has not been considered, so the pressure gradients necessary for fluid flow — which is required for structure formation after matter and radiation decoupled — cannot be taken into account.
+This article resolves both issues.
+The study of the evolution of density perturbations by way of linearization of the Einstein equations and conservation laws with respect to a flat Friedmann-Lemaître-Robertson-Walker (flrw) universe was pioneered by Lifshitz [1] and Lifshitz & Khalatnikov [2].
+They observed that solutions to the linearized Einstein equations and conservation laws are dependent on the choice of a coordinate system, i.e., the solutions are gauge-dependent.
+Consequently, these solutions can be changed by linear coordinate transformations, also known as gauge transformations.
+This means that these solutions have no physical meaning, which is a problem known as the gauge problem of cosmology.
+Solving this long-standing problem is the only way to obtain physically significant solutions from the linearized Einstein equations and conservation laws.
+Doing so is paramount to explaining how the earliest structures [3, 4] in the universe formed from small density perturbations.
+Physical phenomena such as density, temperature, pressure, and entropy perturbations do not depend on the choice of a coordinate system.
+Accordingly, the corresponding mathematical quantities describing these phenomena must be defined without using a reference frame.
+In other words, these quantities must be gauge-invariant.
+Two different pioneering methods to address the gauge problem of cosmology have been proposed in the literature.
+In his seminal article, Bardeen [5] introduced gauge-invariant quantities intended to represent energy density perturbations as linear combinations of gauge-dependent quantities.
+Building on Bardeen’s work, Mukhanov et al.[6, 7] made significant contributions to the subject.
+Ellis et al.[8, 9, 10] present a second method that takes advantage of the fact that the gradient of a gauge-dependent quantity is gauge-invariant.
+For a comprehensive overview of this approach and its further extension, see Tsagas et al.[11].
+Unfortunately, these approaches have some disadvantages.
+Firstly, they allow for different definitions of gauge-invariant quantities.
+Since these approaches lack a Newtonian limit, it is impossible to determine which gauge-invariant quantity corresponds to an energy density perturbation.
+Secondly, these approaches do not account for the particle number density.
+Thirdly, metric perturbations are explicitly present in the evolution equations for density perturbations.
+Finally, these approaches have not been extended to open and closed flrw universes.
+Many new ideas have been added to this field.
+See Ellis’s $2017$ review article [12] and other publications, e.g., Refs.
+[13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
+These researchers proposed alternative perturbation theories using different gauge-invariant quantities that differ from those used by Bardeen and Mukhanov et al.
+However, the most fundamental question regarding which gauge-invariant quantity accurately and unambiguously represents a density perturbation remains unanswered.
+This article presents a relativistic cosmological perturbation theory that does not have the aforementioned disadvantages.
+Firstly, it demonstrates that a gauge-invariant quantity describing a density perturbation can be defined in precisely one way.
+Secondly, in addition to energy density, particle number density [32] is considered.
+Both are crucial for structure formation.
+Finally, it is shown that perturbations in the spatial metric and its derivatives are included in the perturbations of the spatial Ricci scalar and expansion scalar.
+The use of these quantities significantly simplifies the system of linearized Einstein equations and conservation laws, leading not only to a perturbation theory that encompasses closed, flat, and open flrw universes but also allowing the Newtonian limit to be derived.
+Unlike the approaches in Refs.
+[33, 34, 35, 36, 37, 38], this article does not introduce any ad hoc modifications to the theory of relativity, free parameters, Cold Dark Matter (cdm), or the peculiar velocities [39, 40, 41] of matter.
+However, the linearized Einstein equations and conservation laws naturally encompass the velocity of matter caused by pressure gradients, as well as the perturbed Ricci scalar caused by a density perturbation.
+These must be taken into account.
+II Development and Overview of the Perturbation Theory
+Sec. III explains why, in addition to energy density, particle number density must also be considered.
+Sec.
+IV explains the gauge problem in cosmology and provides a solution by demonstrating that a gauge-invariant quantity that describes energy density perturbations can be defined in only one way.
+The same applies to the gauge-invariant perturbation of the particle number density.
+These definitions imply that the gauge-invariant perturbation to the expansion of the universe is equal to zero.
+In Sec.
+X, the Newtonian limit will be used to demonstrate that these quantities represent density perturbations.
+To formulate a perturbation theory based on these quantities, a coordinate system is necessary. The reasons why a synchronous reference frame is the best choice for this task are explained in Sec. V.
+In Sec.
+VI, the Einstein equations and conservation laws for flrw universes, as well as their linearized counterparts, are presented with respect to synchronous coordinates.
+Given that the evolution of density perturbations is considerably influenced by pressure perturbations, the covariant divergence of the fluid velocity, and the perturbation of the Ricci scalar, expressions are derived for these quantities.
+Since we are working in synchronous coordinates, we can apply the decomposition theorem for symmetric, second-rank spatial tensors, such as the perturbed metric tensor and the perturbed spatial Ricci tensor, as well as the Helmholtz decomposition for the fluid velocity vector.
+These theorems are discussed in Sec.
+VII.
+According to these theorems, the linearized Einstein equations and conservation laws decompose into three independent systems of equations: one for tensor perturbations, one for vector perturbations, and one for scalar perturbations.
+In the literature, the decomposition theorem is only applied to the spatial metric tensor perturbation.
+However, this is insufficient.
+Since the perturbed Ricci scalar plays an important role in the evolution of density perturbations and in the Newtonian limit, the decomposition theorem must also be applied to the perturbed Ricci tensor.
+Sec.
+VIII uses these theorems to show that only scalar perturbations are associated with density perturbations.
+In this case, the perturbed Ricci scalar contains the spatial derivatives of two independent potentials, while the perturbed expansion scalar contains their time derivatives.
+In the Newtonian limit, one of the potentials becomes the Newtonian potential while the other becomes irrelevant.
+Since we are interested in density perturbations, we derive in Sec.
+IX a system of equations describing the evolution of scalar perturbations.
+This system contains three conservation laws: one for the energy density perturbation, one for the particle number density perturbation, and one for the momentum.
+It also contains two constraint equations.
+The algebraic energy density constraint equation contains the perturbed Ricci scalar, and the momentum constraint equation describes how the perturbed Ricci scalar evolves.
+The Newtonian limit is reached when the velocities of the particles become negligible compared to the speed of light — that is, when the pressure approaches zero.
+Using this limit it is shown in Sec.
+X that the newly defined gauge-invariant quantities in Sec.
+IV represent perturbations of the energy and particle number densities.
+In other words, in the Newtonian limit, the momentum constraint equation ensures that the Newtonian potential and the energy density perturbation are independent of time.
+It also makes sure that the energy density constraint equation is identical to the Poisson equation of Newtonian gravity.
+Additionally, the well-known special relativistic equation relating rest mass density to energy density is obtained.
+The Newtonian limit can only be obtained using the reformulated evolution equations for scalar perturbations given in Sec.
+IX.
+Since these equations are not presented in this form in the literature, previous cosmological perturbation theories have no Newtonian limit.
+Without a clearly defined Newtonian limit, it is impossible to demonstrate which gauge-invariant quantity represents a density perturbation.
+This article resolves this issue.
+The Newtonian limit demonstrates furthermore that relativistic gauge transformations become Newtonian gauge transformations.
+Consequently, the conventional evolution equation for relative perturbations in a flat flrw universe filled with a nonrelativistic fluid — whether derived from general relativity or Newtonian gravity adapted to the expansion of the universe — also has gauge-dependent solutions devoid of physical meaning.
+Consequently, Newtonian gravity is not suitable for studying density perturbations in the universe, regardless of their scale.
+This means that no distinction can be made between density perturbations inside and outside the horizon.
+The perturbation theory is presented in Sec.
+XI.
+It describes the evolution of density perturbations in closed, flat, and open flrw universes.
+The theory incorporates the perturbed Ricci scalar, the pressure, the entropy, and the covariant divergence of the fluid velocity.
+This is demonstrated by the algorithm in Appendix A.
+In Sec.
+XII, expressions are derived for the perturbations of entropy, temperature, and pressure to connect the perturbation theory to an equation of state for a cosmological fluid.
+It will be shown that pressure perturbations have two components: an adiabatic component caused by the density perturbation itself, and a nonadiabatic component.
+In Sec. XIII, we apply the perturbation theory to a flat flrw universe in its radiation-dominated phase and after the decoupling of matter and radiation.
+Finally, we summarize the results and draw conclusions in Sec. XIV.
+The article is accompanied by three appendices.
+Appendix A presents an algorithm that, when executed with the assistance of a computer algebra program, derives the perturbation theory from the background Einstein equations and the system of equations for scalar perturbations.
+Since the latter system is a new result, its correctness is demonstrated in Appendices B and C.
+Finally, Appendix C compares the conventional perturbation equations with the perturbation theory developed in this article.
+III Equations of State
+We distinguish between two main eras of the universe: the period before the decoupling of matter and radiation and the post-decoupling era.
+Before decoupling, the cosmological fluid can be treated as a relativistic fluid consisting of radiation, ordinary matter, and cdm, where the contributions of matter and cdm to the pressure are negligible and the radiation pressure is of the same order of magnitude as the energy density.
+See, e.g., Ref.
+[6], Eq.
+(5.49), Ref.
+[7], Eq.
+(1.78), and Ref.
+[13], § V-1.
+The conditions in the universe became significantly different after decoupling of matter and radiation.
+This is because the pressure in the cosmological fluid has become negligible compared to its energy density.
+Therefore, pressure can be neglected in the Einstein equations and conservation laws describing the evolution of a flrw universe.
+However, pressure perturbations cannot be ignored when studying structure formation.
+The self-gravity of a local density perturbation creates pressure gradients.
+These gradients then cause particle flows, thereby increasing a local density perturbation.
+The linearized momentum conservation law shows that the absence of pressure perturbations implies an absence of fluid flows.
+This results in static density perturbations.
+Therefore, pressure perturbations are crucial to understanding how structures in the universe formed.
+After decoupling, the cosmological fluid can be treated as a perfect nonrelativistic gas.
+In this case, pressure depends on the average kinetic energy density of the particles.
+Therefore, it is essential to include, next to the rest energy density, also the kinetic energy density in the analysis.
+Both pressure and kinetic energy density depend on the particle number density and temperature.
+Because of this, particle number density and temperature must be included in the equations of state.
 See Ref.
 [42], Sec.
-2.10 on relativistic hydrodynamics, and Ref.
-[32].
-As indicated in Ref.
-[33], the formation of structures in the universe requires a focus on the clumping of a set of particles rather than on the evolution of energy density instabilities.
-The approach presented in this paper yields novel insights into the evolution of density perturbations in both the pre-decoupling and post-decoupling eras.
-This is achieved by examining the evolution of energy density and particle number density perturbations, as well as their interactions.
-A detailed explanation of this can be found in Sec.
-XII.
-II.2 Perturbation Theory
-The equations governing the evolution of the unperturbed universe include three scalars: the energy density, the particle number density, and the expansion.
-As will be demonstrated in Sec.
-III, the perturbations to these three scalars transform under general linear coordinate transformations in the same way.
-This ultimately provides a solution to the gauge problem.
-As will be demonstrated in Sec.
-IV, there are only three sets of three gauge-invariant quantities that can be constructed from these scalars and their gauge-dependent perturbations.
-Each of these sets contains precisely one gauge-invariant quantity that is equal to zero.
-It will be shown that if the gauge-invariant perturbation of the expansion is zero, then it follows from the nonrelativistic limit in Sec.
-IX that the other two gauge-invariant quantities, denoted by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, are the true physical perturbations of the energy density and the particle number density, respectively.
-The subsequent step is to derive the evolution equations for $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
-In order to do this, it is necessary to employ a suitable coordinate system.
-In the context of general relativity, there is no a priori preferred coordinate system.
-Furthermore, the perturbations of the energy and particle number densities are both gauge-invariant, allowing for the freedom to choose any coordinate system.
-In order to facilitate the derivation of the aforementioned equations, it is typical to select a coordinate system that aligns with the specific characteristics of the problem in question.
-In Section Sec.
-V, a detailed justification is provided for the selection of synchronous coordinates.
-In Sec.
-VI, the Einstein equations, conservation laws, and their linearized counterparts with respect to this coordinate system are presented.
-In Sec.
-VII, the synchronous reference system is employed to decompose the coupled system of linearized Einstein equations and conservation laws into three independent systems.
-These systems describe the evolution of tensor, vector, and scalar perturbations, respectively.
-In the extant literature, it is commonly assumed that the decomposition of the linearized equations can be achieved by decomposing only the perturbation of the metric tensor into its tensor, vector, and scalar parts.
-However, this is insufficient.
-Additionally, the perturbation of the spatial Ricci tensor, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}$, must similarly be decomposed into its tensor, vector, and scalar parts, as demonstrated in Ref.
-[34].
-In the case of tensor and vector perturbations, the perturbation of the spatial Ricci scalar, or the spatial curvature perturbation, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, is identically zero.
-In contrast, for scalar perturbations, the value of ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ is not zero.
-In addition, it will be demonstrated that the Newtonian potential is encapsulated within the expression for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$.
-This implies that only scalar perturbations are associated with density perturbations, specifically the physical density perturbations, represented by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
-In Sec.
-VIII, the system of equations (37) is derived which describes the evolution of scalar perturbations.
-This system incorporates not only the usual conservation laws for energy density and momentum, but also a conservation law for the particle number density.
-Furthermore, it is found that this system comprises an algebraic energy density constraint equation (37a) that incorporates the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, in addition to an evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$.
-The latter equation is the covariant divergence of the three linearized momentum constraint equations (12b).
-In the nonrelativistic limit (i.e., the limit in which the pressure vanishes), the algebraic energy density constraint equation (37a) and the evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, when combined, yield the time-independent Poisson equation of Newtonian gravitation with source term $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x})$, as detailed in Sec.
-IX.
-Consequently, the constraint equations (37a)–(37b) are of great importance in the evolution of density perturbations and in the derivation of the nonrelativistic limit.
-The constraint equations presented in this specific formulation are notably absent from the extant literature on cosmological perturbation theories.
-This is the precise reason why an exact and correct nonrelativistic limit for a cosmological perturbation theory has not yet been discovered.
-The absence of a clearly defined nonrelativistic limit makes it impossible to interpret gauge-invariant quantities in previous perturbation theories.
-The issue is further discussed in Sec.
-IX, which addresses the nonrelativistic limit.
-In Sec.
-X, the primary result is presented, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/n_{{\scriptscriptstyle(0)}}$ in closed, flat, and open flrw universes.
-As the quantities $\delta_{\varepsilon}$ and $\delta_{n}$ are gauge-invariant, the evolution equations will have the same mathematical form in all coordinate systems and yield solutions that are independent of the choice of a coordinate system.
-In Sec.
-XII, the recently developed perturbation theory is applied to a flat flrw universe in its phases preceding and following the decoupling of radiation and matter.
-In order to proceed with the analysis, it is necessary to derive gauge-invariant thermodynamic quantities.
-These quantities will be presented in Sec.
-XI.
-Finally, in Sec.
-XIII, a summary of the results is provided.
-The paper is accompanied by three appendices.
-In Appendix A, the primary result, Eqs.
-(60), is derived using the background equations (10), and the equations for scalar perturbations (37).
-The latter system represents a novel result and is indispensable for the proposed methodology.
-A proof of its correctness is provided in Appendices B and C.
-In Appendix C, the systems of equations (10) and (37) are employed once more.
-In this instance, however, the evolution equations (154) and (156) are recovered for the gauge-dependent quantity defined by $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, in a flat flrw universe.
-From these equations, the well-known physical solutions for large-scale perturbations in a flat flrw universe are derived.
-These include the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe following the decoupling of matter and radiation.
-A comparison of Eqs.
-(85) and (155a) and Eqs.
-(109) and (157a) reveals that the novel perturbation theory encompasses all previously established and well-known physical solutions for large-scale density perturbations.
-This constitutes a significant result, as the aforementioned solutions are widely documented in the literature.
-It follows that a comprehensive and accurate perturbation theory must also yield these solutions.
-The homogeneous part of Eqs.
-(154a) and (156a) constitutes the conventional relativistic equation (150).
-Evidently, the latter equation is incomplete and has the gauge mode $\propto t^{-1}$ as solution.
-As will be demonstrated in Sec.
-IX, in the nonrelativistic limit, relativistic gauge transformations converge to Newtonian gauge transformations.
-Consequently, the quantity $\delta$ is also dependent on the choice of coordinates in Newtonian gravitation.
-Accordingly, it possesses no physical meaning.
-It can be concluded that the conventional relativistic evolution equation (150), is inadequate for examining the evolution of density fluctuations in the universe.
-In the extant literature, the relativistic equation (150) is derived from Newtonian gravitation under two conditions: first, the expansion of the universe must be taken into account, and second, the equation of state must be nonrelativistic.
-The presence of the gauge mode in the solution of the aforementioned equation, when derived from the Newtonian theory of gravity, is the reason why Newtonian gravitation is not applicable to the study of density fluctuations in an evolving universe.
-III Gauge Problem of Cosmology
-Einstein’s equations and conservation laws are invariant under general coordinate transformations $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$.
+2.10 on relativistic hydrodynamics, and Refs.
+[32, 43].
+This results in random, nonadiabatic density perturbations, which are crucial for the formation of structures in the universe.
+The aforementioned considerations are consistent with thermodynamic principles, namely that energy density $\varepsilon$ and pressure $p$ are functions of particle number density $n$ and temperature $T$, i.e.,
+$\varepsilon=\varepsilon(n,T),\quad p=p(n,T).$ (1)
+In order to facilitate the calculations, it is advantageous to eliminate the temperature from the given equations of state. This yields the equation of state for the pressure:
+$p=p(n,\varepsilon),$ (2)
+which will be used in the subsequent analysis.
+IV Gauge Problem and its Solution
+Einstein’s equations and conservation laws are invariant under coordinate transformations $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$.
 This signifies that there are no preferred coordinate systems.
-Consequently, the linearized Einstein equations and conservation laws are invariant under a general linear infinitesimal space-time transformation, which is usually referred to as a gauge transformation.
-This transformation is given by
+Consequently, the linearized Einstein equations and conservation laws are gauge-invariant, meaning they are invariant under linear, infinitesimal space-time transformations, also known as gauge transformations, which are given by
 $x^{\mu}\rightarrow x^{\prime\mu}=x^{\mu}-\xi^{\mu}(t,\bm{x}),$ (3)
 where the gauge functions $\xi^{\mu}(t,\bm{x})$ are four arbitrary infinitesimal functions of time, $x^{0}=ct$ with $c$ the speed of light, and space, $\bm{x}=(x^{1},x^{2},x^{3})$, coordinates.
+The gauge problem of cosmology stems from the linear transformation (3) and the linearity of the linearized Einstein equations and conservation laws.
+This problem will now be examined.
 Consider a closed, flat, or open flrw universe.
@@
 Let $S_{\scriptscriptstyle(0)}(t)$ denote the quantities $\varepsilon_{\scriptscriptstyle(0)}$, $n_{\scriptscriptstyle(0)}$, and $\theta_{\scriptscriptstyle(0)}$.
-These quantities satisfy the background, that is to say, the unperturbed Einstein equations and conservation laws.
+These quantities satisfy the background equations, that is to say, the unperturbed Einstein equations and conservation laws.
 Let $S_{\scriptscriptstyle(1)}(t,\bm{x})$ be their perturbed counterparts $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, which satisfy the linearized equations.
-As a consequence of the linearity of these equations, new solutions that are physically equivalent can be generated.
-These solutions are given by (see Weinberg [31], Sec.
+As a consequence of the linearity of these equations, new solutions that have precisely the same physical content can be generated.
+These new solutions are given by (see Weinberg [42], Sec.
 10.9 for a detailed explanation)
-$S^{\prime}_{\scriptscriptstyle(1)}=S_{\scriptscriptstyle(1)}+\mathcal{L}_{\xi}S_{\scriptscriptstyle(0)}=S_{\scriptscriptstyle(1)}+\xi^{0}\dot{S}_{\scriptscriptstyle(0)}.$ (3)
+$S^{\prime}_{\scriptscriptstyle(1)}=S_{\scriptscriptstyle(1)}+\mathcal{L}_{\xi}S_{\scriptscriptstyle(0)}=S_{\scriptscriptstyle(1)}+\xi^{0}\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,\theta.$ (5)
 In this expression, the operator $\mathcal{L}_{\xi}$ is the Lie derivative with respect to the infinitesimal four-vector $\xi^{\mu}$.
@@
 The solutions $S_{\scriptscriptstyle(1)}$ and $S^{\prime}_{\scriptscriptstyle(1)}$ contain the so-called gauge modes which are given by $\hat{S}_{\scriptscriptstyle(1)}=\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$.
-This indicates that the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are dependent on the choice of a system of reference, rendering them gauge-dependent and, consequently, devoid of physical significance.
-This is the gauge problem of cosmology, whereby the coordinate artifacts and the underlying physics are inextricably linked in the solution of the linearized equations.
-Even when physical initial values are imposed in order to solve the linearized equations, the resulting solution lacks physical meaning.
-It is therefore imperative to resolve the gauge problem in order to facilitate an investigation into the evolution of density perturbations in the universe.
-IV Solving the Gauge Problem of Cosmology
-In contrast to the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, their physical counterparts are independent of the choice of a coordinate system.
-This is to say that they are gauge-invariant.
-As the physics of the density perturbations is concealed within the general solution of the linearized Einstein equations and conservation laws, it follows that the physical counterparts of $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, must be linear combinations of gauge-dependent solutions to these equations.
-In order to identify these quantities, it is necessary to consider the fact that a quantity is gauge-invariant if it is independent of the coordinate transformation parameter $\xi^{\mu}$, as outlined by Stewart [3].
-This implies that the quantity must be invariant under a linear coordinate transformation (1).
-As is evident from (3), the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are independent of the spatial components, $\xi^{i}$, of the transformation, but not of the temporal component, $\xi^{0}$.
-Given that the transformation of these quantities under the general infinitesimal transformation (1) is identical to that described by (3), it is possible to construct gauge-invariant quantities by combining these gauge-dependent quantities in a way that eliminates the gauge modes $\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$.
+This indicates that the solutions $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ to the linearized equations are gauge-dependent.
+Since these solutions depend on the choice of a reference frame, they are devoid of physical significance.
+Therefore, $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ do not represent physical phenomena.
+This is the gauge problem of cosmology, whereby the coordinate artifacts and the underlying physics are inextricably linked in the solution to the linearized equations.
+Now that the gauge problem of cosmology has been understood, a solution can be found.
+The physics of the perturbations is hidden in the solutions to the linearized Einstein equations and conservation laws.
+Consequently, quantities that describe physical phenomena must be linear combinations of gauge-dependent solutions to these equations.
+Since the transformation of the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ under the infinitesimal transformation (3) is identical to that described by (5), we can construct gauge-invariant quantities by combining these three gauge-dependent quantities in a way that eliminates the gauge modes $\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$.
 This results in three distinct sets of linear combinations.
 In each of these sets, precisely one gauge-invariant quantity is identically zero.
-Given that the focus is on energy density and particle number density perturbations, the only possible set is
+Since we are only interested in energy density and particle number density perturbations, the only possible set is
 $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\varepsilon_{\scriptscriptstyle(1)}-\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad n_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=n_{\scriptscriptstyle(1)}-\dfrac{\dot{n}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},$ (6a)
 $\theta_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\theta_{\scriptscriptstyle(1)}-\dfrac{\dot{\theta}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}=0.$ (6b)
-Local density perturbations exert no influence on the expansion of the universe, as evidenced by (4b). The transformation rule (3), illustrates that the quantities (4) are indeed gauge-invariant.
-In Sec.
-IX, it is demonstrated that in the nonrelativistic limit, where the pressure is considered to be zero, the linearized conservation laws lack physical significance and are decoupled from the linearized energy density constraint equation and momentum constraint equations.
-The combination of these constraint equations results in the time-independent Poisson equation of Newtonian gravitation with a source term $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})$.
-Furthermore, the special relativistic relation $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}mc^{2}$, where $m$ is the rest mass, is recovered.
-In light of the aforementioned considerations, it can be concluded that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the actual physical perturbations of the energy density and the particle number density, respectively, in the context of general relativity.
-The sole method for constructing gauge-invariant quantities for the energy density and particle number density perturbations is provided by the expressions (4).
-As these quantities also possess the correct nonrelativistic limit, it can be concluded that the gauge problem of cosmology has been resolved.
-This outcome is an inevitable result of the mathematical consistency of general relativity.
-That is to say, the general theory of relativity allows for only one gauge-invariant quantity that accurately represents a physical density perturbation.
-V Selection of a Reference System
+From these definitions, it can be inferred that the general form of a gauge-invariant quantity describing a physical phenomenon is given by
+$q^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}:=q_{{\scriptscriptstyle(1)}}-\dfrac{\dot{q}_{{\scriptscriptstyle(0)}}}{\dot{\theta}_{{\scriptscriptstyle(0)}}}\theta_{{\scriptscriptstyle(1)}},$ (7)
+where $q$ stands for energy density $\varepsilon$, particle number density $n$, temperature $T$, pressure $p$, or the entropy per particle $s$.
+The quantities (6a) represent density perturbations and (6b) represents the perturbation to the expansion of the universe. This will be demonstrated in Sec. X on the Newtonian limit.
+V Selection of a Reference Frame
 The quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ are defined without the use of a coordinate system.
-However, to derive evolution equations for these quantities, it is first necessary to establish a system of reference.
+However, to derive evolution equations for these quantities, it is first necessary to choose a suitable reference frame.
 Since general relativity is covariant, and the quantities in question are gauge-invariant, it follows that the evolution equations for these quantities can be derived in any chosen coordinate system.
 The selection of an appropriate coordinate system for this problem is dependent on two criteria, which are outlined below.
-Firstly, the interpretation of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ requires taking the nonrelativistic limit.
+Firstly, to show that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent density perturbations requires taking the Newtonian limit.
 In the context of general relativity, coordinate transformations are in general space-time transformations, specifically $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$.
 In the linear case, space-time transformations are given by (3).
-In Newtonian gravitation, the space and time transformations are regarded as distinct and independent entities.
-Consequently, when taking the nonrelativistic limit, the relativistic space-time transformations must be automatically separated into independent space and time transformations.
-Secondly, it would be advantageous to have a coordinate system that would facilitate the derivation of the evolution equations.
+In Newtonian gravity, space and time are different and independent entities.
+Consequently, in the Newtonian limit, the relativistic space-time transformations (3) must be automatically separated into independent space and time transformations.
+Secondly, it would be preferable to have a coordinate system that would facilitate the derivation of the perturbation theory.
 It will now be demonstrated that both of these requirements can be met by employing the same coordinate system.
-In Newtonian gravitation, where space and time are treated as distinct and independent entities, all coordinate systems are inherently synchronous.
-In view of the nonrelativistic limit, it can be seen that a synchronous reference system represents the optimal choice.
+Firstly, in Newtonian gravity, where space and time are treated as distinct and independent entities, all coordinate systems are inherently synchronous.
+In view of the Newtonian limit, it can be seen that a synchronous reference frame represents the optimal choice.
 In these coordinates, the metric tensor $g_{\mu\nu}(t,\bm{x})$ of the flrw universes is given by
@@
 The metric tensor of the three-dimensional maximally symmetric subspaces of constant time is denoted by $\tilde{g}_{ij}$.
-The Killing equations, $\xi_{\mu;\nu}+\xi_{\nu;\mu}=0$, and (5) demonstrate that the gauge functions $\xi^{\mu}(t,\bm{x})$ in the transformation (1) become
+The Killing equations $\xi_{\mu;\nu}+\xi_{\nu;\mu}=0$ and (8) demonstrate that the gauge functions $\xi^{\mu}(t,\bm{x})$ in the transformation (3) become
 $\xi^{0}=\psi(\bm{x}),\quad\xi^{i}=\tilde{g}^{ik}(\bm{x})\dfrac{\partial\psi(\bm{x})}{\partial x^{k}}\int\!\!\frac{{\mathrm{d}}t}{a^{2}(t)}+\chi^{i}(\bm{x}),$ (9)
-if only transformations between synchronous coordinates are allowed.
+if only transformations between synchronous coordinates are allowed (see Ref.
+[42], Sec.
+10.9).
 In expressions (9), $\psi(\bm{x})$ and $\chi^{i}(\bm{x})$ are four arbitrary infinitesimal functions of the spatial coordinates.
-In Sec.
-IX, it will be proven that in the nonrelativistic limit, the function $\psi(\bm{x})$ is an arbitrary infinitesimal constant.
-This means that the transformations of space and time are independent of each other in this limit.
-The second requirement will now be considered.
-Synchronous coordinates possess the property that the space-space components, $R_{ij}$, of the four-dimensional Ricci curvature tensor, $R_{\mu\nu}$, are partitioned into two distinct parts.
-One part exclusively comprises the time derivatives of the spatial metric tensor, $g_{ij}$, while the second part is the Ricci curvature tensor of the three-dimensional subspaces, denoted by ${{}^{3}\!}R_{ij}$.
+Secondly, in synchronous coordinates the space-space components $R_{ij}$ of the four-dimensional Ricci tensor $R_{\mu\nu}$ are partitioned into two distinct parts.
+One part exclusively comprises the time derivatives of the spatial metric tensor $g_{ij}$ while the second part is the Ricci tensor ${{}^{3}\!}R_{ij}$ of the three-dimensional subspaces.
 This is demonstrated in Ref.
 [44], § 97.
-As the perturbations of the spatial metric tensor and the spatial Ricci tensor are also tensors (see Sec.
-VI.2), the covariant decomposition theorems in Refs.
-[34, 36, 3] may be employed to decompose these perturbed tensors into their scalar, vector, and tensor parts.
-This decomposition permits the system of linearized Einstein equations and conservation laws to be decomposed into three independent systems, each of which describes the evolution of scalar, vector, and tensor perturbations.
-As will be demonstrated in Sec.
-VII, only scalar perturbations are associated with density perturbations.
-Accordingly, it is sufficient to consider the system for scalar perturbations (37).
-The aforementioned system is more straightforward to utilize than the original set of equations (12), thus facilitating the derivation of the evolution equations for density fluctuations in synchronous coordinates in comparison to any other coordinate system.
-In view of the fact that synchronous coordinates are likewise compatible with Newton’s theory of gravitation, they will be employed in the derivation of the primary result (60).
+This implies that the spatial metric and Ricci tensor perturbations are both tensors.
+Furthermore, the spatial part of the fluid four-vector is a vector in synchronous coordinates.
+This allows us to apply the decomposition theorem for second-rank, symmetric three-tensors and the Helmholtz theorem to decompose the fluid vector.
+Using these theorems, it is shown that the linearized Einstein equations and conservation laws can be decomposed into three mutually exclusive systems: one for tensor perturbations, one for vector perturbations, and one for scalar perturbations.
+Next, it will be demonstrated that only scalar perturbations are associated with density perturbations.
+Therefore, when studying density perturbations, it is sufficient to consider the evolution equations for scalar perturbations.
+This system is easier to use than the original set of equations, thus facilitating the derivation of a cosmological perturbation theory.
+Because synchronous coordinates are also compatible with Newtonian gravity, they will be used from this point onward.
 VI Einstein Equations and Conservation Laws
-In the preceding section, it was concluded that synchronous coordinates are the optimal choice for addressing the problem at hand.
-In this section, the background Einstein equations and conservation laws, along with their linearized counterparts, will be expressed in this particular coordinate system.
+As was concluded in the previous section, synchronous coordinates are the optimal choice for deriving a perturbation theory.
+This section presents the background Einstein equations and conservation laws, as well as their linearized counterparts, in this coordinate system.
 The background energy density constraint equation and its linearized counterpart are expressed in a manner consistent with the contracted Bianchi identities.
-These identities state that the Riemann curvature tensor satisfies the equation $R^{\mu}{}_{\nu;\mu}=\tfrac{1}{2}R_{,\nu}$.
+These identities state that the Riemann tensor satisfies the equation $R^{\mu}{}_{\nu;\mu}=\tfrac{1}{2}R_{,\nu}$, where a comma denotes ordinary differentiation.
 This can be expressed in an equivalent manner as the four-divergence of the Einstein tensor, i.e., $G^{\mu\nu}{}_{;\nu}=0$}.
-This result demonstrates that the constraint equations contain at most first-order time derivatives of the metric, which is precisely what is required for the development of the novel perturbation theory.
+This result demonstrates that the constraint equations contain at most first-order time derivatives of the metric, which is precisely what is required for the development of a cosmological perturbation theory.
 For further details see Ref.
@@
 [44], § 92 and § 95.
-It is understood from thermodynamic principles that the energy density, $\varepsilon$, and the pressure, $p$, are functions of both the particle number density, $n$, and the temperature, $T$:
-$\varepsilon=\varepsilon(n,T),\quad p=p(n,T).$ (7)
-In order to facilitate the calculations, it is advantageous to eliminate the temperature from the given equations of state. This yields the equation of state for the pressure:
-$p=p(\varepsilon,n),$ (8)
-which will be utilized in the subsequent analysis.
 VI.1 Background Equations
@@
 $\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (11d)
-It can be demonstrated that the $G_{0i}$ constraint equations and the $G_{ij}$ dynamical equations with $i\neq j$ are satisfied identically.
-The $G_{ii}$ dynamical equations are equivalent to the time derivative of the $G_{00}$ constraint equation (Friedmann equation) given by equation (10a).
-Therefore, the $G_{ij}$ dynamical equations can be omitted.
-The cosmological constant, $\Lambda$, the gravitational constant, $G_{\text{N}}$, and the speed of light, $c$, are fundamental constants.
-The overdot is used to denote differentiation with respect to $x^{0}=ct$.
+The $G_{0i}$ constraint equations and the $G_{ij}$ dynamical equations with $i\neq j$ are satisfied identically.
+The $G_{ii}$ dynamical equations are identical to the time derivative of the $G_{00}$ constraint equation (Friedmann equation) (11a).
+Therefore, the $G_{0i}$ and $G_{ij}$ equations can be omitted.
+The cosmological constant $\Lambda$, the gravitational constant $G_{\text{N}}$, and the speed of light $c$ are fundamental constants.
+The overdot denotes differentiation with respect to $x^{0}=ct$.
 The Hubble function $H$ is defined by $H:=\dot{a}/a$.
 For flrw universes, the Hubble function is given by $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, where $\theta_{\scriptscriptstyle(0)}$ is the background value of the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$, where $u^{\mu}$ is the four-velocity $u^{\mu}:=c^{-1}U^{\mu}$, normalized to unity, $u^{\mu}u_{\mu}=1$.
-A semicolon is used to denote covariant differentiation with respect to the background metric tensor $g_{{\scriptscriptstyle(0)}\mu\nu}$ (5).
+A semicolon denotes covariant differentiation with respect to the background metric tensor $g_{{\scriptscriptstyle(0)}\mu\nu}$ (8).
 The equations (11c) and (11d) represent the energy density conservation law and the particle number conservation law, respectively.
-It is crucial to highlight that the variable $w$ in Eq.
-(10c) is an abbreviation for $p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, and it does not represent the equation of state.
-The spatial parts of the background Riemann tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}$, the Ricci curvature tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}$, and its contraction, ${{}^{3}\!}R_{\scriptscriptstyle(0)}$, are given by the following expressions:
+The variable $w$ in Eq.
+(11c) is an abbreviation for $p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, and it does not represent an equation of state.
+The spatial parts of the background Riemann tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}$, the Ricci tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}$, and its contraction ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ are given by
 ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}=\tilde{R}^{i}{}_{jkl}=K(\delta^{i}{}_{k}\tilde{g}_{jl}-\delta^{i}{}_{l}\tilde{g}_{jk}),$ (12a)
@@
 The value of $K$ determines the nature of the flrw universe.
-The universe is open for $K=-1$, flat for $K=0$, and closed for $K=+1$.
+The universe is open for K=1K=-1, flat for $K=0$, and closed for $K=+1$.
 VI.2 Linearized Equations
-The system of linearized Einstein equations and conservation laws for closed, flat, and open flrw universes is succinctly expressed as follows:
+The system of linearized Einstein equations and conservation laws for closed, flat, and open flrw universes is given by
 $H\dot{h}^{k}{}_{k}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\kappa\varepsilon_{\scriptscriptstyle(1)},$ (13a)
@@
 $\ddot{h}^{i}{}_{j}+3H\dot{h}^{i}{}_{j}\,$+
-$\qquad\delta^{i}{}_{j}H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}=-\kappa\delta^{i}{}_{j}(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}),$ (12c)
+[…]
 $\dot{\varepsilon}_{\scriptscriptstyle(1)}+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})+\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)}=0,$ (13d)
-$\frac{1}{c}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}\right)\,-$
+[…]
 $\qquad g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}=0,$ (13e)
 $\dot{n}_{\scriptscriptstyle(1)}+3Hn_{\scriptscriptstyle(1)}+n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}=0.$ (13f)
-In these equations, the following notation is employed for the perturbed metric tensor: $h_{\mu\nu}:=-g_{{\scriptscriptstyle(1)}\mu\nu}$ and $h^{\mu\nu}:=g^{\mu\nu}_{{\scriptscriptstyle(1)}}$.
+In these equations, the perturbed metric tensor is given by $h_{\mu\nu}:=-g_{{\scriptscriptstyle(1)}\mu\nu}$ and $h^{\mu\nu}:=g^{\mu\nu}_{{\scriptscriptstyle(1)}}$.
 In light of the use of synchronous coordinates, it follows that $h_{00}=0$ and $h_{0i}=0$.
-In addition, the raising and lowering of indices is performed by the spatial background metric tensor as defined in expression (5), namely, $h^{i}{}_{j}=g^{ik}_{\scriptscriptstyle(0)}h_{kj}$ where $g^{ik}_{\scriptscriptstyle(0)}=-\tilde{g}^{ik}/a^{2}$.
-A vertical bar denotes covariant differentiation with respect to the spatial background metric tensor $g_{{\scriptscriptstyle(0)}ij}=-a^{2}\tilde{g}_{ij}$ or, equivalently, with respect to $\tilde{g}_{ij}$.
+Raising and lowering of indices is performed by the spatial background metric tensor as defined in expression (8), namely, $h^{i}{}_{j}=g^{ik}_{\scriptscriptstyle(0)}h_{kj}$ where $g^{ik}_{\scriptscriptstyle(0)}=-\tilde{g}^{ik}/a^{2}$.
+A vertical bar denotes covariant differentiation with respect to the spatial background metric tensor $g_{{\scriptscriptstyle(0)}ij}=-a^{2}\tilde{g}_{ij}$.
 In the context of flrw universes, where $\Gamma^{k}_{{\scriptscriptstyle(0)}ij}=\tilde{\Gamma}^{k}{}_{ij}$, the operations of taking the time derivative and the covariant derivative are shown to commute.
-Equation (12a) represents the linearized energy density constraint equation (linearized Friedmann equation), while Eqs.
-(12b) correspond to the linearized momentum constraint equations.
+Equation (13a) represents the linearized energy density constraint equation (linearized Friedmann equation), while Eq.
+(13b) corresponds to the linearized momentum constraint equation.
 The linearized dynamical equations are presented in Eqs.
 (13c).
-Equations (12d) and (12e) represent the linearized energy density conservation law and the linearized energy-momentum conservation laws, respectively.
+Equations (13d) and (13e) represent the linearized energy density conservation law and the linearized momentum conservation law, respectively.
 Finally, the linearized particle number density conservation law is given by Eq.
 (13f).
-As previously stated in Sec.
-II, the pressure perturbation, the spatial curvature perturbation, and the covariant divergence of the spatial fluid velocity play a pivotal role in a cosmological perturbation theory.
-In consequence, the expressions for these quantities are now derived.
+Pressure perturbations, spatial curvature perturbations, and the covariant divergence of the fluid velocity play pivotal roles in the evolution of density perturbations.
+Therefore, these quantities must be included in the perturbation theory.
+The next three subsections will derive expressions for these quantities.
 VI.2.1 Pressure Perturbation
-The gauge-dependent perturbation to the pressure (8) is given by
+The gauge-dependent quantity $p_{{\scriptscriptstyle(1)}}$, which is associated with pressure (2), is given by
 $p_{{\scriptscriptstyle(1)}}=p_{n}n_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}},$ (14a)
 $p_{n}:=\left(\dfrac{\partial p}{\partial n}\right)_{\varepsilon},\quad p_{\varepsilon}:=\left(\dfrac{\partial p}{\partial\varepsilon}\right)_{n}.$ (14b)
-The quantity representing the true physical perturbation, designated as $p^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, will be derived in Sec. XI on thermodynamics.
+The quantity $p^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ representing the pressure perturbation will be derived in Sec. XII on thermodynamics.
 VI.2.2 Spatial Curvature Perturbation
-Lifshitz’ expression for the perturbed connection coefficients presented in Ref. [2], Eq. (I.3) and Ref. [31], Eq. (10.9.1) is a tensor, which is given by
+Lifshitz’s expression for the perturbed connection coefficient presented in Ref. [2], Eq. (I.3) and Ref. [42], Eq. (10.9.1) is a tensor, which is given by
 $\Gamma^{k}_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j}+h_{lj\mid i}-h_{ij\mid l}).$ (15)
@@
 ${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=\Gamma^{k}_{{\scriptscriptstyle(1)}ij\mid k}-\Gamma^{k}_{{\scriptscriptstyle(1)}ik\mid j}.$ (16)
-By combining (14) and (15), the following expression for the perturbed spatial Ricci tensor may be derived:
+By combining (15) and (16), the following expression for the perturbed spatial Ricci tensor can be derived (see Ref. [42], Eq. (10.9.3):
 ${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j\mid k}+h_{lj\mid i\mid k}-h_{ij\mid l\mid k}-h_{lk\mid i\mid j}).$ (17)
-By raising the index $i$, one arrives at the following result, employing (11b):
+By raising the index $i$ and using (12b), one arrives at
 ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}:=$ $\,(g^{ik}\,{{}^{3}\!}R_{kj})_{\scriptscriptstyle(1)}=g^{ik}_{\scriptscriptstyle(0)}\,{{}^{3}\!}R_{{\scriptscriptstyle(1)}kj}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}$
-$=$ $-\tfrac{1}{2}g^{il}_{\scriptscriptstyle(0)}(h^{k}{}_{l\mid j\mid k}+h^{k}{}_{j\mid l\mid k}-h^{k}{}_{k\mid l\mid j})$
+[…]
 $+\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}h^{i}{}_{j\mid k\mid l}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}$. (18)
-By employing $g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid j\mid k}=g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid k\mid j}$, one arrives at the contraction of the perturbed spatial Ricci tensor:
-${{}^{3}\!}R_{\scriptscriptstyle(1)}:={{}^{3}\!}R^{k}_{{\scriptscriptstyle(1)}k}=g_{\scriptscriptstyle(0)}^{ij}(h^{k}{}_{k\mid i\mid j}-h^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{k}{}_{k}.$ (18)
+By using $g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid j\mid k}=g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid k\mid j}$, one arrives at the contraction of the perturbed spatial Ricci tensor:
+${{}^{3}\!}R_{\scriptscriptstyle(1)}:={{}^{3}\!}R^{k}_{{\scriptscriptstyle(1)}k}=g_{\scriptscriptstyle(0)}^{ij}(h^{k}{}_{k\mid i\mid j}-h^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{k}{}_{k}$.
+(19)
 In Sec.
-VII it will be demonstrated that the expression (18) represents the local perturbation to the spatial curvature ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}$, or the spatial curvature perturbation, induced by local density perturbations.
-The spatial curvature perturbation ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ is incorporated into the evolution equations (60), as shown in Eq.
+VII it will be demonstrated that the expression (19) represents the local perturbation to the spatial curvature ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}$ induced by local density perturbations.
+The quantity ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ will be used in Sec.
+IX to eliminate the metric perturbations and their spatial derivatives from the momentum constraint equation (13b).
+It is incorporated into the perturbation theory using Eq.
 (127).
-VI.2.3 Covariant Divergence of the Spatial Fluid Velocity
-In the background flrw universe, the fluid four-velocity is given by $u^{\mu}_{\scriptscriptstyle(0)}=\delta^{\mu}{}_{0}$, so that the perturbation $\theta_{\scriptscriptstyle(1)}$ to the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$ is
+VI.2.3 Covariant Divergence of the Fluid Velocity
+[…]
 $\theta_{\scriptscriptstyle(1)}=\vartheta_{\scriptscriptstyle(1)}-\tfrac{1}{2}\dot{h}^{k}{}_{k},\quad\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k},$ (20)
-where $\vartheta_{\scriptscriptstyle(1)}$ is the covariant divergence of the spatial part $u^{i}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity $u^{\mu}_{\scriptscriptstyle(1)}$.
-In deriving (19), the following identity is employed: $(u^{k}{}_{|k})_{{\scriptscriptstyle(1)}}=u^{k}_{{\scriptscriptstyle(1)}|k}$, which follows from $u^{i}_{{\scriptscriptstyle(0)}}=0$.
-The covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the evolution equations (60), as shown in Eq.
+where $\vartheta_{\scriptscriptstyle(1)}$ is the covariant divergence of the fluid vector.
+The quantity $\theta_{{\scriptscriptstyle(1)}}$ will be used in Sec.
+IX to eliminate the time derivative of the metric perturbation from the energy density constraint equation (13a) in favor of $\vartheta_{{\scriptscriptstyle(1)}}$.
+The latter quantity is incorporated into the perturbation theory using Eq.
 (132).
-VII Decomposition of Spatial Tensors
-The following important theorem is used to decompose the linearized Einstein equations and conservation laws into three independent systems.
-In Refs.
-[34, 36, 3], it is demonstrated that any symmetric spatial tensor ${\cal{T}}_{ij}$ of rank two can be unambiguously decomposed in a covariant manner into three irreducible components:
+VII Decomposition of Spatial Tensors and Vectors
+The system of linearized Einstein equations and conservation laws (13) is not suitable for deriving the perturbation theory because it contains solutions related to and unrelated to density perturbations.
+Therefore, our goal is to transform the system (13) into one that is exclusively linked to density perturbations and excludes all other solutions.
+To that end, we will use two key theorems to break the linearized equations down into three independent systems.
+These theorems are presented in the next two subsections.
+VII.1 Tensor Decomposition
+Refs. [45, 46, 47] demonstrate that any rank-two symmetric spatial tensor ${\cal{T}}_{ij}$ can be decomposed into three irreducible components in a covariant manner:
 ${\cal{T}}^{i}{}_{j}={\cal{T}}^{i}_{\parallel\,j}+{\cal{T}}^{i}_{\perp\,j}+{\cal{T}}^{i}_{\ast\,j}.$ (21)
-The constituents possess the following properties:
+The components have the following properties:
 ${\cal{T}}^{k}_{\perp\,k}=0,\quad{\cal{T}}^{k}_{\ast\,\,k}=0,\quad{\cal{T}}^{k}_{\ast\;i\mid k}=0,\quad{\cal{T}}^{i}_{\parallel\,j}=\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j}$, (22)
-where $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$ are two independent potentials. Consequently, the component $h^{i}_{\parallel\,j}$ can be expressed in terms of the two independent potentials as follows:
-$h^{i}_{\parallel\,j}=\frac{2}{c^{2}}\bigl{(}\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j}\bigr{)}.$ (22)
-The factor $2/c^{2}$ is included in view of the nonrelativistic limit.
-Evidently, from (21), it can be inferred that an expression analogous to (22) exists for the tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\parallel\,j}$.
-In the context of cosmological perturbation theory, only the contraction ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ of the latter tensor is required.
-Upon substituting (22) into (18), the following result is obtained:
+where $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$ are two independent potentials.
+Consequently, the component $h^{i}_{\parallel\,j}$ of the perturbed metric can be expressed in terms of the two independent potentials as follows:
+[…]
+The factor $2/c^{2}$ is included in view of the Newtonian limit. By substituting (23) into (19), the following result is obtained:
 ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}=\dfrac{2}{c^{2}}\Bigl[2\phi^{|i}{}_{|i}$ $+\zeta^{|k}{}_{|k}{}^{|i}{}_{|i}-\zeta^{|k|i}{}_{|k|i}$
-$+\tfrac{1}{3}\,{{}^{3}\!}R_{{\scriptscriptstyle(0)}}(3\phi+\zeta^{|k}{}_{|k})\Bigr{]},$ (23)
+[…]
 and for the perturbed expansion (20) it is found
-$\theta_{{\scriptscriptstyle(1)}}=\vartheta_{{\scriptscriptstyle(1)}}-\dfrac{1}{c^{2}}\bigl{(}3\dot{\phi}+\dot{\zeta}^{|k}{}_{|k}\bigr{)}.$ (24)
-As evidenced by (VII), the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ encompasses two distinct potentials $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$.
-In the nonrelativistic limit, as will be discussed in detail in Sec.
-IX, the potential $\phi$ becomes independent of time and equal to the Newtonian potential, while the potential $\zeta$ becomes inconsequential.
-The spatial component $\bm{u}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity can be decomposed as follows
+[…]
+The expression (VII.1) reveals that two potentials, $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$, are associated with cosmological density perturbations in open and closed flrw universes.
+For a flat flrw universe, however, density perturbations are characterized by one potential, $\phi(t,\bm{x})$.
+This is because, for scalar perturbations, the quantity $\theta_{{\scriptscriptstyle(1)}}$ can be eliminated from expressions (6a) and Eqs.
+(37), as was done in (122) and (123).
+As will be demonstrated in Sec. X, in the Newtonian limit, the potential $\phi(t,\bm{x})$ becomes identical to the Newtonian potential, $\phi(\bm{x})$.
+VII.2 Vector Decomposition
+Since $\bm{u}_{\scriptscriptstyle(1)}$ is a vector in three-space, it can be decomposed using the Helmholtz decomposition as follows:
 $\bm{u}_{{\scriptscriptstyle(1)}}=\bm{u}_{{\scriptscriptstyle(1)}\parallel}+\bm{u}_{{\scriptscriptstyle(1)}\perp},$ (26)
 where the components $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ and $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ have the properties
-~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel},\quad\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{{\scriptscriptstyle(1)}\perp}.$ (26)
-In these expressions, the generalized vector differential operator, denoted by $\bm{\tilde{\nabla}}$, is defined by the relation ~${\tilde{\nabla}_{i}v_{k}:=v_{k\mid i}}$.
-The expressions (25)–(26) are referred to as the Helmholtz decomposition.
-The following section will demonstrate that the decomposition of ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}$ and $h^{i}{}_{j}$ as given by (20) and the properties (21) are consistent with the the Riemann tensor (11a) and the momentum constraint equations (12b).
-It is important to note that the velocity vector $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is not uniquely determined by the system of equations (12).
-Since $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is irrotational, it is possible to supplement it with the gradient of an arbitrary function.
+$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel},\quad\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{{\scriptscriptstyle(1)}\perp}.$ (27)
+[…]
+Since $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is irrotational, it can be supplemented with the gradient of an arbitrary function.
 Given that the system of equations (13) is invariant under the gauge transformation (3), where the gauge function $\xi^{\mu}(t,\bm{x})$ is given by (9), it follows that
 $u^{\prime}_{{\scriptscriptstyle(1)}\parallel i}=u_{{\scriptscriptstyle(1)}\parallel i}+\psi_{|i},$ (28)
-is also a solution of the system (12). As a result, the gauge modes associated with $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ are given by
+is also a solution to the system (13). As a result, the gauge modes associated with $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ are given by
 $\hat{u}_{{\scriptscriptstyle(1)}\parallel i}=\psi_{|i},\quad\hat{u}^{i}_{{\scriptscriptstyle(1)}\parallel}=g_{{\scriptscriptstyle(0)}}^{ik}\psi_{|k}=-\dfrac{1}{a^{2}}\tilde{g}^{ik}\psi_{|k}.$ (29)
-This result is of particular significance in the context of the derivation of the nonrelativistic limit, as presented in Sec. IX.
-VII.1 Decomposition of the Linearized Equations
-The system of equations (12) is split into three independent systems by using the decomposition (20) with the properties (21) for $h_{ij}$ and ${{}^{3}\!}R_{ij}$, as well as (25)–(26).
+This result will be used in Sec. X on the Newtonian limit.
+VIII Decomposition of the Linearized Equations
+The system of equations (13) is split into three independent systems by using the decomposition (21) with the properties (22) for $h_{ij}$ and ${{}^{3}\!}R_{ij}$, as well as the decomposition (26) with the properties (27) for $\bm{u}_{{\scriptscriptstyle(1)}}$.
 The solutions to these systems are conventionally designated as tensor perturbations $\ast$, vector perturbations $\perp$, and scalar perturbations $\parallel$.
 In the following three subsections, it is demonstrated that scalar perturbations are the only ones associated with density perturbations $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
-VII.1.1 Tensor Perturbations
-From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\ast}=0$.
-This, in combination with Eq.
-(12a), yields ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$.
-Eqs.
-(12b) yield the result that ${\bm{u}_{\scriptscriptstyle(1)}=\bm{0}}$.
-This, in turn, implies, with (19) that $\theta_{{\scriptscriptstyle(1)}}=0$.
+VIII.1 Tensor Perturbations
+From the expressions (19) and (22), it follows that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\ast}=0$, and Eq.
+(13b) yields $\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$}, so that, with (20), $\theta_{{\scriptscriptstyle(1)}}=0$.
+Equation (13a) yields $\varepsilon_{{\scriptscriptstyle(1)}}=0$}.
 From (6a), it follows that $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$}.
 Given that $\theta_{{\scriptscriptstyle(1)}}=0$, Eqs.
-(10d) and (12f) are identical.
-This implies that $n_{{\scriptscriptstyle(1)}}=0$.
-This, in turn, implies that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$.
-Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$.
-In consequence, the evolution equations for tensor perturbations are as follows:
+(11d) and (13f) are identical, so $n_{{\scriptscriptstyle(1)}}=0$.
+This, in turn, implies with (6a), that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$.
+Finally, the result obtained from either (13d) or the contraction of (13c) is that $p_{{\scriptscriptstyle(1)}}=0$.
+Consequently, the evolution equations for tensor perturbations are as follows:
 $\ddot{h}^{i}_{\ast\,j}+3H\dot{h}^{i}_{\ast\,j}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\ast j}=0.$ (30)
 Given the form of these equations, tensor perturbations are typically referred to as gravitational waves.
-VII.1.2 Vector Perturbations
-From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=0$. This implies that
-$h^{kl}_{\perp\mid k\mid l}=0.$ (30)
-Moreover, it is found from Eq.
-(12a), that ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$.
-Raising the index $i$ of Eqs.
-(12b) with $g^{ij}_{\scriptscriptstyle(0)}$, and subsequently taking the covariant derivative with respect to the index $j$, one finds, using that $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$,
-$\dot{h}^{kj}_{\perp{\mid k\mid j}}+2Hh^{kj}_{\perp{\mid k\mid j}}=-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{j}_{{\scriptscriptstyle(1)}\mid j}.$ (31)
-The combination of (30) and (31) results in ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=0$.
-This indicates that only the component $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ remains.
-From (19) one finds that $\theta_{{\scriptscriptstyle(1)}}=0$.
-This, combined with $\varepsilon_{{\scriptscriptstyle(1)}}=0$, yields $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$.
-From $\theta_{{\scriptscriptstyle(1)}}=0$ one finds that Eq.
-(12f) is identical to the background equation (10d), implying that $n_{{\scriptscriptstyle(1)}}=0$.
-This, in turn, implies with (4a) that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$.
-Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$.
+VIII.2 Vector Perturbations
+Expressions (19) and (22), imply that, for vector perturbations, we have ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=-h^{kj}_{\perp|k|j}$.
+Raising the index $i$ of the momentum constraint equation (13b) with $g^{ij}_{\scriptscriptstyle(0)}$, and subsequently taking the covariant derivative with respect to the index $j$, one finds, using that $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$,
+$\dot{h}^{kj}_{\perp{\mid k\mid j}}+2Hh^{kj}_{\perp{\mid k\mid j}}=-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{j}_{{\scriptscriptstyle(1)}\perp\mid j}.$ (31)
+The right-hand side of this equation vanishes, as can be seen from (27).
+Consequently, the covariant divergence of the momentum constraint equation (31) can only be satisfied if ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=0$.
+The consequences are as follows.
+Equation (13a) implies that $\varepsilon_{{\scriptscriptstyle(1)}}=0$}, and (20) yields $\theta_{{\scriptscriptstyle(1)}}=0$.
+The latter identity shows that Eq.
+(13f) is identical to the background equation (11d), so $n_{{\scriptscriptstyle(1)}}=0$}.
+From (6a) we find that $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$.
+Finally, either (13d) or the contraction of (13c) implies that $p_{{\scriptscriptstyle(1)}}=0$.
 Therefore, the system of equations for vector perturbations is given by
 $\!\dot{h}^{k}_{\perp i\mid k}+2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}\perp i}=0,$ (32a)
-$\!\ddot{h}^{i}_{\perp j}+3H\dot{h}^{i}_{\perp j}+2R^{i}_{{\scriptscriptstyle(1)}\perp j}=0,$ (32b)
+$\!\ddot{h}^{i}_{\perp j}+3H\dot{h}^{i}_{\perp j}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\perp j}=0,$ (32b)
 $\!\frac{1}{c}\frac{{\text{d}}}{{\text{d}}t}\Bigl(\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}\Bigr)+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}=0.$ (32c)
 According to (27), vector perturbations are also referred to as rotational perturbations.
-VII.1.3 Scalar Perturbations
-In this particular instance one has ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}\neq 0$.
-By taking the covariant derivative of (12b) with respect to the index $j$ and subsequently substituting the expression (22), the following result is obtained:
+VIII.3 Scalar Perturbations
+In this particular instance one has ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}\neq 0$, as follows from expression (VII.1).
+By taking the covariant derivative of the momentum constraint equation (13b) with respect to the index $j$ and subsequently substituting the expression (23), the following result is obtained:
 $2\dot{\phi}_{\mid i\mid j}+\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}=\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}i\mid j}.$ (33)
-Upon interchanging the indices $i$ and $j$ and subtracting the result from Eqs. (33) one finds
+By interchanging the indices $i$ and $j$ and subtracting the result from Eqs. (33) one finds
 $\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}=-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}),$ (34)
 where it is used that $\dot{\phi}_{\mid i\mid j}=\dot{\phi}_{\mid j\mid i}$ and $\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}=\dot{\zeta}^{\mid k}{}_{\mid k\mid j\mid i}$.
-By rearranging the covariant derivatives, Eqs.
-(34) can be transformed in the following form:
-$(\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid j\mid k})-(\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i\mid k})\,+$
+Rearranging the second-order covariant derivatives converts Eq.
+(34) into a form with only commutators on the left-hand side:
+[…]
 $\qquad(\dot{\zeta}^{\mid k}{}_{\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i})_{\mid k}\,=$
 $\qquad-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}).$ (35)
-The expressions for the commutator of second-order covariant derivatives (see Ref. [31], Sec. 6.5) are given by
+The commutators of second-order covariant derivatives (see Ref. [42], Sec. 6.5) are given by
 $A^{i}{}_{j\mid p\mid q}-A^{i}{}_{j\mid q\mid p}=A^{i}{}_{k}\,{{}^{3}\!}R^{k}_{{\scriptscriptstyle(0)}jpq}-A^{k}{}_{j}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq},$ (36a)
 $B^{i}{}_{\mid p\mid q}-B^{i}{}_{\mid q\mid p}=B^{k}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq}.$ (36b)
-Upon substituting the background Riemann tensor (11a), it is found that the left-hand sides of Eqs.
-(35) vanished identically.
-This implies that ~×$\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$, so that only the component $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ remains.
-Since ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}\neq 0$ implies that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$, it can thus be concluded that only scalar perturbations are coupled to density perturbations.
-VIII Evolution Equations for Scalar Perturbations
-As demonstrated in the preceding section, the evolution of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is determined by the background equations (10) and the equations that govern the evolution of scalar perturbations.
-In this section, the findings of the preceding section are utilized to reformulate the system (12) into a novel set of equations that exclusively describe the evolution of scalar perturbations.
-Given the subsequent focus on scalar perturbations, the subscript $\parallel$ is omitted.
+Upon substituting the background Riemann tensor (12a), the left-hand side of Eq.
+(35) vanishes identically.
+This implies that $\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$, so that only the component $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ remains.
+Since $\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}\neq 0$ implies that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$, it can thus be concluded that only scalar perturbations are coupled to density perturbations.
+IX Evolution Equations for Scalar Perturbations
+As demonstrated in the preceding section, the evolution of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is determined by the background equations (11) and the equations that describe the evolution of scalar perturbations.
+In this section, the findings from the previous section are used to reformulate the system (13) into a new system of equations that exclusively describes the evolution of scalar perturbations.
+Given the subsequent focus on density perturbations, the subscript $\parallel$ is omitted.
 The evolution equations for scalar perturbations are as follows:
@@
 $\dot{n}_{\scriptscriptstyle(1)}=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (37e)
-The symbol ~$\tilde{\nabla}^{2}$ denotes the Laplace-Beltrami operator with respect to the metric $\tilde{g}_{ij}(\bm{x})$ of three-dimensional subspaces of constant time:
-~$\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}:=\tilde{g}^{ij}p_{{\scriptscriptstyle(1)}\mid i\mid j},\quad g^{ij}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|i|j}=-\dfrac{\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}}{a^{2}}.$ (38)
+[…]
+The symbol $\tilde{\nabla}^{2}$ denotes the Laplace-Beltrami operator with respect to the metric $\tilde{g}_{ij}(\bm{x})$ of three-dimensional subspaces of constant time:
+$\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}:=\tilde{g}^{ij}p_{{\scriptscriptstyle(1)}\mid i\mid j},\quad g^{ij}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|i|j}=-\dfrac{\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}}{a^{2}}.$ (38)
 The parameter $\beta$ is defined as follows:
 $\beta^{2}:=\dfrac{\dot{p}_{{\scriptscriptstyle(0)}}}{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}.$ (39)
-By employing the expression $\dot{p}_{\scriptscriptstyle(0)}=p_{n}\dot{n}_{\scriptscriptstyle(0)}+p_{\varepsilon}\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and eliminating the time derivatives of $\varepsilon_{\scriptscriptstyle(0)}$ and $n_{\scriptscriptstyle(0)}$ with the aid of the conservation laws (10c) and (10d), the following result is obtained:
+Using the expression $\dot{p}_{\scriptscriptstyle(0)}=p_{n}\dot{n}_{\scriptscriptstyle(0)}+p_{\varepsilon}\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and eliminating the time derivatives of $\varepsilon_{\scriptscriptstyle(0)}$ and $n_{\scriptscriptstyle(0)}$ with the aid of the conservation laws (11c) and (11d), the following result is obtained:
 $\beta^{2}=p_{\varepsilon}+\dfrac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (40)
 where $p_{\varepsilon}$ and $p_{n}$ are given by (14b).
-The derivation of Eqs.
-(37) will now be conducted.
-To initiate this process, the term $\dot{h}^{k}{}_{k}$ is eliminated from Eq.
-(12a) using the expression (19), which results in the algebraic energy density constraint equation (37a).
-Given that only $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is associated with scalar perturbations, it is possible to simplify the momentum constraint equations (12b) and the momentum conservation laws (12e) by employing the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}:=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}$, (19), as an alternative to $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$.
-By multiplying both sides of Eqs.
+We will now derive Eqs. (37). To begin, the term $\dot{h}^{k}{}_{k}$ is eliminated from Eq. (13a) using the expression (20). This results in the algebraic energy density constraint equation (37a).
+Given that only $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is associated with scalar perturbations, it is possible to simplify the momentum constraint equation (13b) and the momentum conservation law (13e) by using the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}:=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}$}, (20), as an alternative to $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$.
+By multiplying both sides of Eq.
 (13b) by $g^{ij}_{\scriptscriptstyle(0)}$ and taking the covariant derivative with respect to the index $j$, one finds
@@
 Indeed, differentiating (19) with respect to time, and using the relation $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$ and (11b), yields the following result:
-${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\dot{h}^{k}{}_{k}.$ (42)
+${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\dot{h}^{k}{}_{k}$.
+(42)
 By combining (41) and (42) and using (20) to eliminate $\dot{h}^{k}{}_{k}$, one obtains Eq.
 (37b).
-In consequence, the three $G^{0}_{{\scriptscriptstyle(1)}i}$ momentum constraint equations (12b) have been reformulated in the form of a first-order ordinary differential equation (37b) for the perturbed spatial Ricci scalar (18).
-The momentum conservation laws (12e) are rewritten by performing the differentiation with respect to time and using (10c), then dividing by $\varepsilon_{{\scriptscriptstyle(0)}}(1+w)$.
-The result is as follows:
+Consequently, the $G^{0}_{{\scriptscriptstyle(1)}i}$ momentum constraint equation (13b) has been reformulated as a first-order ordinary differential equation (37b) for the perturbed spatial Ricci scalar (VII.1) expressed in the potentials.
+The momentum conservation law (13e) is reformulated by first performing the differentiation with respect to time.
+Then, we use Eq.
+(11c) to eliminate $\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$.
+Next, dividing the result by $\varepsilon_{{\scriptscriptstyle(0)}}(1+w)$ yields
 $\dot{u}^{i}_{{\scriptscriptstyle(1)}}+H(2-3w)u^{i}_{{\scriptscriptstyle(1)}}+\dfrac{\dot{w}}{1+w}u^{i}_{{\scriptscriptstyle(1)}}-\dfrac{g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}}{\varepsilon_{{\scriptscriptstyle(0)}}(1+w)}=0.$ (43)
@@
 $\dot{w}=3H(1+w)(w-\beta^{2}).$ (44)
-Upon substituting this expression into Eq. (43), it can be seen that the momentum conservation laws can be reformulated as
+Upon substituting this expression into Eq. (43), the momentum conservation law becomes
 $\dot{u}^{i}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})u^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}=0.$ (45)
-By taking the covariant divergence of (45) with respect to the background metric tensor (5), and employing (19) and (38), one obtains Eq. (37d).
-As demonstrated in Appendix B, the dynamical equations (12c) are superfluous since the system (37) comprises the constraint equations and conservation laws, and thus its solution automatically satisfies the dynamical equations.
-Consequently, the derivation of the equations (37) has been successfully completed.
-In Appendix A, the primary result, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}/n_{{\scriptscriptstyle(0)}}$, will be derived using the systems (10) and (37).
-Prior to proceeding, it is crucial to demonstrate that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the true physical energy density and particle number density perturbations, respectively.
-This necessitates the nonrelativistic limit being employed for the interpretation of these quantities.
-IX Nonrelativistic Limit
-The standard nonrelativistic limit is defined by four fundamental criteria (see, e.g., Ref. [37], Sec. 4.1):
-  1. 1.
-    The subspaces of constant time are flat.
-  2. 2.
-    When the pressure is reduced to zero, there are no local fluid flows.
-  3. 3.
-    Relativistic gauge transformations are reduced to Newtonian gauge transformations.
-  4. 4.
-    The gravitational field of a density perturbation is required to be constant.
-The third requirement is of considerable importance, as it permits the selection of a coordinate system in Newtonian gravitation.
-This fact is not referenced in the existing literature.
-The fourth requirement is a direct consequence of the first and second requirement.
-In the case of flat three-spaces, it is evident that ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$.
-In a flat flrw universe, covariant derivatives are equivalent to ordinary derivatives, thereby simplifying the expression (VII) to the following form:
+Taking the covariant divergence of (45) with respect to the spatial background metric tensor (8), and using (20) and (38) yields Eq. (37d).
+As demonstrated in Appendix B, the dynamical equation (13c) is redundant since the system (37) comprises the constraint equations and conservation laws.
+Thus, the derivation of Eqs.
+(37) has been completed.
+X Newtonian Limit
+As demonstrated in Sec.
+IV, the gauge-invariant quantities, $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, which are intended to describe density perturbations, can only be defined in one way.
+This section proves that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent perturbations to energy and particle number densities, respectively.
+In Newtonian gravity, space is Euclidean.
+Consequently, the Newtonian limit can only be realized in a flat flrw universe.
+In this case, the Newtonian limit is reached when gravity is weak and the velocities of the particles are negligible compared to the speed of light.
+These conditions are met in a linear perturbation theory and when pressure approaches zero.
+In a flat flrw universe where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, covariant derivatives are equal to ordinary derivatives. This simplifies the expression (VII.1) to
 ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}=\dfrac{4}{c^{2}}\phi^{\mid k}{}_{\mid k}=-\dfrac{4}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}.$ (46)
-The symbol $\nabla^{2}$ represents the conventional Laplace operator.
-Upon substituting the expression (46) into the perturbation equations (37), and employing the relation $H:=\dot{a}/a$, the following result is obtained:
+The symbol $\nabla^{2}$ represents the conventional Laplace operator. By substituting the expression (46) into the perturbation equations (37) and using that $H:=\dot{a}/a$, we obtain the result:
 $-H\vartheta_{\scriptscriptstyle(1)}$ $=-\dfrac{1}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)},\quad\kappa=8\pi G_{\text{N}}/c^{4},$ (47a)
@@
 $\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)},\quad\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (47e)
-In order to derive the $G^{0}_{{\scriptscriptstyle(1)}0}$ constraint equation (47a) from Eq.
-(37a), it is first necessary to eliminate the function $\varepsilon_{{\scriptscriptstyle(1)}}$ using the expression (4a).
+To derive the $G^{0}_{{\scriptscriptstyle(1)}0}$ constraint equation (47a) from Eq.
+(37a), the function $\varepsilon_{{\scriptscriptstyle(1)}}$ must first be eliminated using the expression (6a).
 The result is
 $2H(\theta_{{\scriptscriptstyle(1)}}-\vartheta_{{\scriptscriptstyle(1)}})=-\dfrac{2}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\kappa\left(\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\dfrac{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}{\dot{\theta}_{{\scriptscriptstyle(0)}}}\theta_{{\scriptscriptstyle(1)}}\right).$ (48)
-Subsequently, the Friedmann equation (10a) with the Ricci scalar ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ set to zero, the energy conservation law (10c), and the relation between the Hubble parameter and the expansion scalar, namely, $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, are employed to obtain the result $\dot{\varepsilon}_{\scriptscriptstyle(0)}/\dot{\theta}_{\scriptscriptstyle(0)}=2H/\kappa$.
-Upon substituting this expression into (48), the desired result, Eq.
-(47a), is obtained.
-With the condition of ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, the first requirement is satisfied.
-Subsequently, the second requirement will be implemented. Upon substituting $p=0$, i.e., $w=0$ and $\beta=0$, into the momentum conservation laws (45), the result is
-$\dot{u}^{i}_{{\scriptscriptstyle(1)}}=-2Hu^{i}_{{\scriptscriptstyle(1)}}.$ (49)
-By lowering the index $i$ with the background metric $g_{{\scriptscriptstyle(0)}ij}$ given by (5), and using the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$, one finds
-$\dot{u}_{{\scriptscriptstyle(1)}i}=0.$ (50)
-These equations admit only functions of the spatial coordinates as solutions.
-As these equations are part of the system (12), it can be concluded that the gauge modes, $\psi_{|i}$, given by (28), are the only solutions of equations (50).
-Consequently, it can be deduced that
-$p\rightarrow 0\quad\Rightarrow\quad u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0.$ (51)
-In the absence of pressure, local pressure gradients are nonexistent. Consequently, the emergence of local fluid flows is precluded. This implies that
-$u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0\quad\Rightarrow\quad\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})\rightarrow\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x}).$ (52)
-Therefore, in the event of a zero fluid flow, the density perturbations are observed to remain static. It can be concluded that the second requirement is fulfilled by the limits (51)–(52).
-Given that the physical parts of the spatial components of the fluid four-velocity are identically zero, only the gauge modes $\psi_{|i}$ remain.
-It can thus be concluded with confidence that the gauge modes may be taken to be equal to zero without any loss of physical information.
-Taking $\hat{u}_{{\scriptscriptstyle(1)}i}=\psi_{|i}=0$ implies that $\psi$ is an arbitrary infinitesimal constant.
-Upon substituting $\psi=C$ into the expressions (6), the result is that the relativistic gauge transformation between synchronous coordinates reduces in the limits given in (51) to the gauge transformation of Newtonian gravitation:
+Setting ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$ in the Friedmann equation (11a) and differentiating the result with respect to time, and using the relation $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, we obtain $\dot{\varepsilon}_{\scriptscriptstyle(0)}/\dot{\theta}_{\scriptscriptstyle(0)}=2H/\kappa$.
+Substituting this expression into Eq.
+(48) yields the desired result, Eq.
+(47a).
+The Newtonian limit can now be derived using the system (47).
+In a universe filled with a nonrelativistic fluid, we have $w\ll 1$ and $\beta^{2}\ll 1$.
+Neglecting the quantities $w$ and $\beta^{2}$ with respect to the constants of order one in the momentum conservation law (45), yields
+$\dot{u}^{i}_{\scriptscriptstyle(1)}+2Hu^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}-\dfrac{p_{{\scriptscriptstyle(1)}|j}}{\varepsilon_{\scriptscriptstyle(0)}}=0.$ (49)
+The second equation is derived by lowering the index $i$ in the first equation using the background metric $g_{{\scriptscriptstyle(0)}ij}$ and the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$.
+Accordingly, the evolution of density perturbations is highly dependent upon the presence of pressure gradients, represented by $p_{{\scriptscriptstyle(1)}|j}$, where $p_{{\scriptscriptstyle(1)}}$ is given by (14).
+In a pressureless fluid, i.e., when $w=0$ and $\beta=0$, however, the result is
+$\dot{u}^{i}_{{\scriptscriptstyle(1)}}+2Hu^{i}_{{\scriptscriptstyle(1)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}=0.$ (50)
+The second equation has solutions that are only functions of the spatial coordinates.
+Since these equations are part of the system (13) that is invariant under the gauge transformation (3) with $\xi^{\mu}(t,\bm{x})$ given by (9), it can be inferred that the gauge modes given by (29), are the only solutions to equations (50).
+Therefore, it can be concluded that
+$p\rightarrow 0\quad\Rightarrow\quad u_{{\scriptscriptstyle(1)}i}^{{\textnormal{phys}}}(t,\bm{x})\rightarrow 0.$ (51)
+In other words, in the absence of pressure, local pressure gradients $p_{{\scriptscriptstyle(1)}|i}$ are nonexistent, so local fluid flows are impossible.
+Thus, in the Newtonian limit, density perturbations are independent of time, that is, they are static.
+The remaining part of this section examines the consequences of the limit (51).
+Since the components (51) of the fluid velocity are zero, only the gauge modes $\hat{u}_{{\scriptscriptstyle(1)}i}=\psi_{|i}$ remain.
+We can set these gauge modes to zero without losing any physical information.
+Substituting $\psi_{|i}=0$ into the expressions (9) yields $\xi^{0}=C$ and $\xi^{i}=\chi^{i}$, implying that the relativistic gauge transformation (3) reduces to the gauge transformation of Newtonian gravity:
 $t\rightarrow t^{\prime}=t-C,\quad x^{i}\rightarrow x^{\prime i}=x^{i}-\chi^{i}(\bm{x}).$ (52)
-This implies that within the Newtonian theory of gravitation, there is the possibility of shifting the time coordinate and selecting the spatial coordinates at will.
-In accordance with the expressions (53), the third requisite is thereby fulfilled.
-The fourth and final requirement will now be discussed.
-Given that $u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}=0$ and $\hat{u}^{i}_{{\scriptscriptstyle(1)}}=0$, it follows that $\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k}=0$.
-By combining this result with (52), the linearized Einstein equations and conservation laws (47) are reduced to
-$\nabla^{2}\phi(t,\bm{x})$ $=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}),$ (54a)
+Thus, we arrive at the well-known fact that in Newtonian gravity, we can freely select the spatial coordinates and shift the time coordinate at will, and that space and time are independent of each other.
+Since both $\hat{u}_{{\scriptscriptstyle(1)}i}=0$ and $u_{{\scriptscriptstyle(1)}i}^{{\textnormal{phys}}}=0$ we find from (20) that $\vartheta_{\scriptscriptstyle(1)}=0$.
+By substituting the latter into the system (47) we obtain
+$\nabla^{2}\phi(t,\bm{x})$ $=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x}),$ (53a)
 $\nabla^{2}\dot{\phi}(t,\bm{x})$ $=0,$ (53b)
@@
 $\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (53d)
-The constraint equations (54a) and (54b) are combined to yield the following result:
-$\nabla^{2}\phi(\bm{x})=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t_{0})\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}).$ (55)
-The potential $\varphi(\bm{x})$ is defined as $\varphi(\bm{x}):=\phi(\bm{x})/a^{2}(t_{0})$, which results in the Poisson equation of Newtonian gravitation:
-$\nabla^{2}\varphi(\bm{x})=4\pi G_{\text{N}}\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{c^{2}}.$ (56)
-With (56), the fourth and final requirement has been met.
-In the nonrelativistic limit, the background equations (10) are given by
+The constraint equations (53a) and (53b) imply
+$a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=a^{2}(t_{0})\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t_{0},\bm{x}).$ (54)
+This equation applies at any given time $t_{0}$. From the Friedmann equation (11a), it follows that, for ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, one may take $a(t_{0})=1$, so that
+$a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}).$ (55)
+This expression is consistent with the limit (51) that results in static density perturbations. Combining (53a), (53b), and (55) yields
+$\nabla^{2}\phi(\bm{x})=4\pi G_{\text{N}}\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{c^{2}},$ (56)
+which is the Poisson equation of Newtonian gravity.
+Finally, the meaning of $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is established. For a universe filled with a pressureless fluid, the background equations (11) are given by
 $3H^{2}=\kappa\varepsilon_{\scriptscriptstyle(0)}+\Lambda,\quad\dot{\varepsilon}_{\scriptscriptstyle(0)}=-3H\varepsilon_{\scriptscriptstyle(0)},\quad\dot{n}_{\scriptscriptstyle(0)}=-3Hn_{\scriptscriptstyle(0)}.$ (57)
 Consequently, the conservation laws imply that $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$.
-As may be observed, Eqs.
-(54c)–(54d) yield solely gauge modes as solutions: specifically, $\hat{\varepsilon}_{\scriptscriptstyle(1)}=C\dot{\varepsilon}_{\scriptscriptstyle(0)}$, $\hat{n}_{{\scriptscriptstyle(1)}}=C\dot{n}_{{\scriptscriptstyle(0)}}$, and $\hat{\theta}_{{\scriptscriptstyle(1)}}=C\dot{\theta}_{{\scriptscriptstyle(0)}}$, where $C$ is an infinitesimal constant of the gauge transformation (53) of Newtonian gravitation.
+As can be verified by substitution, Eqs.
+(53c)–(53d) yield solely gauge modes as solutions: specifically, $\hat{\varepsilon}_{\scriptscriptstyle(1)}=C\dot{\varepsilon}_{\scriptscriptstyle(0)}$, $\hat{n}_{{\scriptscriptstyle(1)}}=C\dot{n}_{{\scriptscriptstyle(0)}}$, and $\hat{\theta}_{{\scriptscriptstyle(1)}}=C\dot{\theta}_{{\scriptscriptstyle(0)}}$, where $C$ is an infinitesimal constant of the gauge transformation (52) of Newtonian gravity.
 This, combined with $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$ yields $\varepsilon_{{\scriptscriptstyle(1)}}=n_{{\scriptscriptstyle(1)}}mc^{2}$.
-Combining the latter two relations with the definitions (4a) one finds the well-known special relativistic relation between the rest mass density and energy density:
+Combining the latter two relations with the definitions (6a) one finds
 $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})mc^{2}.$ (58)
-The equations (54c) and (54d) represent the residual form of the conservation laws (47c) and (47e) in the nonrelativistic limit.
-It is evident that equations Eqs.
-(54c) and (54d) have no physical significance.
-These equations are decoupled from the physical constraint equations (54a) and (54b) and thus do not form part of Newtonian gravitation.
-Therefore, the potential $\zeta$ which appears in $\theta_{{\scriptscriptstyle(1)}}$, (24), is no longer relevant.
-In view of the aforementioned considerations, it becomes evident that solely the Poisson equation (56) and the special relativistic relation (58) have survived.
-Following an examination of the four points regarding the nonrelativistic limit, the following conclusion is reached.
-It has been demonstrated that the system of linearized Einstein equations and conservation laws for scalar perturbations (37), when combined with the gauge-invariant quantities (4a), reduces in the nonrelativistic limit to the Poisson equation (56) of Newtonian gravitation and the special relativistic relation (58).
-This indicates that the gauge-invariant quantities $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, defined by expressions (4a), represent the local perturbations of the energy density and the particle number density, respectively, in Newtonian gravitation.
-Therefore, it can be stated that $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ represent the actual local density perturbations of the energy density and the particle number density in general relativity.
-In conclusion, since the gauge-invariant perturbation of the expansion, $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ (4b), is identically zero, it can be stated that density perturbations have no influence on the expansion of the universe.
-The subsequent section of this study will proceed to derive the evolutionary equations for $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$.
-These equations permit the investigation of the evolution of density fluctuations in closed, flat, and open flrw universes independently of the choice of coordinate system.
-X Evolution Equations for Density Fluctuations
-In light of the preceding analysis, it can be concluded that the quantities represented by $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, are the real physical perturbations to the energy density and particle number density, respectively.
-Furthermore, it has been demonstrated that the evolution equations for scalar perturbations (37), when combined with the background equations (10), govern the evolution of the energy density perturbation and the particle number density perturbation.
-In Appendix A, the derivation of evolution equations for their corresponding density fluctuations will be presented.
-These density fluctuations are defined as follows:
-$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{\scriptscriptstyle(0)}(t)},\quad\delta_{n}(t,\bm{x}):=\dfrac{n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{n_{\scriptscriptstyle(0)}(t)}.$ (59)
-The final result is a system of evolution equations for the density fluctuations $\delta_{\varepsilon}$ and $\delta_{n}$, in closed, flat, and open flrw universes. These equations are given by
+This is the well-known special relativistic relation between rest mass and energy.
+The equations (53c) and (53d) represent the residual form of the conservation laws (47c) and (47e) in the Newtonian limit.
+Since Eqs.
+(53c) and (53d) only have gauge modes as solutions, they clearly have no physical significance.
+These equations are decoupled from the physical constraint equations (53a) and (53b) and thus do not form part of Newtonian gravity.
+Therefore, the potential $\zeta$ which appears in $\theta_{{\scriptscriptstyle(1)}}$ (25) is no longer relevant.
+This is consistent with the comment just below (25).
+Thus, in the Newtonian limit, only the Poisson equation (56) and the special relativistic relation (58) remain.
+We now come to a final conclusion.
+The analysis above demonstrates that, in the Newtonian limit, only the Poisson equation (56) and the relation (58) remain.
+Therefore, $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ (6a) represent the perturbations in energy and particle number density, respectively, in Newtonian gravity.
+Consequently, they also represent perturbations in energy and particle number density in general relativity.
+Finally, we note that, since $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$ (6b), density perturbations do not affect the expansion of the universe.
+In the next subsections, relative density perturbations in the Newtonian limit are discussed, and the perturbation theory based on Newtonian gravity is examined in the context of relativity theory.
+X.1 Relative Density Perturbation in the Newtonian Limit
+The expression (55) implies that
+$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{a^{2}(t)}.$ (59)
+Since Eq. (11c) with $w=0$ implies $\varepsilon_{{\scriptscriptstyle(0)}}\propto a^{-3}$, it follows that the relative density perturbation $\delta_{\varepsilon}$ evolves according to
+$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{{\scriptscriptstyle(0)}}(t)}\propto\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})a(t).$ (60)
+Thus, due to the expansion of the universe, the relative density perturbation $\delta_{\varepsilon}$ for a static density perturbation $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})$ increases proportionally to the scale factor $a(t)$ of the universe.
+In other words, a static density perturbation becomes relatively denser when viewed against the decreasing energy density $\varepsilon_{{\scriptscriptstyle(0)}}$ in the background universe.
+X.2 Newtonian Gravity
+As demonstrated in Appendix C, the relative perturbation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ is gauge-dependent.
+The gauge mode associated with $\delta$ is given by (152a).
+For $w\ll 1$, it is a solution to the homogeneous part of Eq.
+(157a).
+This gauge mode does not disappear from the solution because $H\neq 0$ in a nonempty universe.
+Consequently, the quantity $\delta$ depends on the choice of coordinates and therefore has no physical meaning.
+As can be seen from Eqs.
+(57), $H\neq 0$ even in the Newtonian limit.
+Therefore, $\delta$ must also depend on the choice of coordinates in that limit and consequently has no physical significance.
+This is consistent with the fact that, in the Newtonian limit, the relativistic gauge transformation (3) converges to the Newtonian gauge transformation (52).
+Given that the system (151) reduces to the system (157) for a nonrelativistic fluid, it can be concluded that the system (157) is inadequate for examining the evolution of density perturbations.
+The existing literature derives Eq.
+(156), from Newtonian gravity under two conditions: the expansion of the universe must be considered and the equation of state must be nonrelativistic, i.e., $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}\ll 1$ and $\beta^{2}\ll 1$.
+See, for example, Ref.
+[7], Sec.
+6.2, Ref.
+[42], Sec.
+15.9, and Ref.
+[48], Eq.
+(3.17).
+However, the gauge transformation (52) shows that $\delta$ depends on the choice of coordinates in Newtonian gravity.
+Therefore, it also lacks physical significance in Newtonian gravity.
+Furthermore, a comparison of the conventional equation (156) and the exact equation (157a) reveals that equation (156) is incomplete.
+Consequently, the conventional evolution equation (156) — whether derived from general relativity or Newtonian gravity modified to account for the expansion of the universe — also has gauge-dependent solutions.
+These findings indicate that Newtonian gravity is inadequate in accurately describing the evolution of cosmological density perturbations.
+Therefore, there exists no distinction between sub- and super-horizon density perturbations.
+XI Relativistic Cosmological Perturbation Theory
+As demonstrated in Sec.
+VIII, the evolution equations for scalar perturbations (37) when combined with the background equations (11) describe the evolution of energy density and particle number density perturbations.
+In Appendix A, the derivation of evolution equations for their corresponding relative density perturbations defined by
+$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{\scriptscriptstyle(0)}(t)},\quad\delta_{n}(t,\bm{x}):=\dfrac{n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{n_{\scriptscriptstyle(0)}(t)},$ (61)
+will be presented. The result is a system of evolution equations for the relative density perturbations $\delta_{\varepsilon}$ and $\delta_{n}$, in closed, flat, and open flrw universes, given by
 $\ddot{\delta}_{\varepsilon}+b_{1}\dot{\delta}_{\varepsilon}+b_{2}\delta_{\varepsilon}$ $=b_{3}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right),$ (62a)
 $\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right)$ $=\frac{3Hn_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right).$ (62b)
-The coefficients $b_{1}$, $b_{2}$ and $b_{3}$ are as follows:
+The background equations (11) and an equation of state (2) for the pressure determine the coefficients $b_{1}$, $b_{2}$, and $b_{3}$. These coefficients are given by
 $b_{1}=$ $\,\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{H}-2\dfrac{\dot{\beta}}{\beta}-H(2+6w+3\beta^{2})$
@@
 $-{{}^{3}\!}R_{\scriptscriptstyle(0)}\left(\tfrac{1}{2}w+\dfrac{H^{2}(1+6w)(1+3\beta^{2})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)$
-$-\beta^{2}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right),$ (61b)
+[…]
 $b_{3}=$ $\,\Biggl[\dfrac{-18H^{2}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\Biggl(\varepsilon_{\scriptscriptstyle(0)}p_{\varepsilon n}(1+w)+\dfrac{2p_{n}}{3H}\dfrac{\dot{\beta}}{\beta}$
 $+p_{n}(p_{\varepsilon}-\beta^{2})+n_{\scriptscriptstyle(0)}p_{nn}\Biggr)+p_{n}\Biggr]$
-×$\times\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right).$ (61c)
+[…]
 In these expressions the partial derivatives of the pressure, $p_{\varepsilon}$ and $p_{n}$, are defined by (14b).
 The second-order partial derivatives are defined by $p_{nn}:={\partial^{2}p}/{\partial n^{2}}$ and $p_{\varepsilon n}:={\partial^{2}p}/{\partial\varepsilon\,\partial n}$.
-It is essential to recognize that, with regard to the nonrelativistic limit, the aforementioned equations apply solely in the case of $p\not\equiv 0$.
-XI Thermodynamic Quantities
-The combined first and second laws of thermodynamics are expressed in terms of density fluctuations, denoted by $\delta_{\varepsilon}$ and $\delta_{n}$ and gauge-invariant expressions are provided for the local pressure and temperature perturbations.
-XI.1 Thermodynamic Laws
+The local curvature perturbation, ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, caused by a local density perturbation as well as the divergence, ${\vartheta}_{{\scriptscriptstyle(1)}}$, of the local fluid velocity caused by local pressure gradients, are incorporated into system (62) via expressions (127) and (132), respectively.
+The system of equations (62) allows us to study the evolution of the relative density perturbations (61) in closed $K=+1$, flat $K=0$, and open K=1K=-1 flrw universes, where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=-6K/a^{2}$.
+Sec.
+XIII will demonstrate that $p_{n}<0$ before and after the decoupling of radiation and matter.
+For $p_{n}>0$, however, the coefficient in Eq.
+(62b) is positive, causing $\delta_{n}-\delta_{\varepsilon}/(1+w)$ to grow exponentially, which is not physically plausible.
+In the system (62), Eq.
+(62a) is the evolution equation for energy density perturbations.
+Since $p_{n}<0$ for an equation of state $p=p(n,\varepsilon)$, the right-hand side of Eq.
+(62a) is nonzero.
+This indicates that density perturbations are nonadiabatic.
+Density perturbations are adiabatic when the source term of Eq.
+(62a) vanishes, i.e., when
+$\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}=0.$ (64)
+Therefore, Eq. (62b) can be regarded as the entropy equation. An explanation is provided by Sec. XII.1.
+Since $p_{n}<0$, the entropy equation (62b) shows that
+$\left|\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right|\rightarrow 0.$ (65)
+This means that relative perturbations in energy density are coupled with those in particle number density.
+Since general relativity is indifferent to the composition of matter, this is true for cdm as well as for ordinary matter, where it is assumed that cdm interacts only via gravity.
+Therefore, the contraction of cdm before decoupling was impossible.
+This precludes the formation of potential wells into which ordinary matter could fall after decoupling to form structures.
+Consequently, the initiation of structure formation by cdm after decoupling is also impossible.
+XII Thermodynamic Quantities
+To link an equation of state to Eqs.
+(62), one needs gauge-invariant expressions for the entropy, temperature, and pressure perturbations.
+These quantities must be expressed in terms of the relative density perturbations $\delta_{\varepsilon}$ and $\delta_{n}$.
+This section presents the derivation of these expressions.
+XII.1 Entropy Perturbations
 The combined first and second laws of thermodynamics for a simple, single-species system is given by (see, for example, Ref. [49], Sec. 2.1)
@@
 $\mu=\dfrac{\varepsilon+p}{n}-Ts.$ (68)
-Eliminating $\mu$ in (63) with the aid of (64) reveals that the combined first and second laws of thermodynamics  (62) can be expressed in a form that does not include $\mu$ or $N$ (see Ref. [33]):
+Eliminating $\mu$ in (67) with the aid of (68) reveals that the combined first and second laws of thermodynamics (66) can be expressed in a form that does not include $\mu$ and $N$ (see Ref. [32]):
 $T{\mathrm{d}}s={\mathrm{d}}\Bigl(\dfrac{\varepsilon}{n}\Bigr)+p\,{\mathrm{d}}\Bigl(\dfrac{1}{n}\Bigr).$ (69)
 From the background equations (11) and the thermodynamic law (69) it can be demonstrated that $\dot{s}_{\scriptscriptstyle(0)}=0$}, indicating that the expansion of the universe takes place without generating entropy.
-From the transformation (3) it can be concluded that $s_{\scriptscriptstyle(1)}=s_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ is automatically gauge-invariant.
-By employing (138) and $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$, the thermodynamic relation (65) can be rewritten as
-$T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n^{2}_{\scriptscriptstyle(0)}}\left(n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\right).$ (66)
-By employing the definitions (59), the following is obtained:
+From the definition of gauge-invariant quantities (7), it can be concluded that $s_{\scriptscriptstyle(1)}=s_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ is automatically gauge-invariant.
+By using (136) and $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$, the thermodynamic relation (69) can be rewritten as
+[…]
+By using the definitions (61), the following is obtained:
 $T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n_{\scriptscriptstyle(0)}}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (71)
-Therefore, $s_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}$ represents the local perturbation to the entropy per particle.
-XI.2 Temperature and Pressure Perturbations
-Given that both the background temperature, denoted by $T_{{\scriptscriptstyle(0)}}$, and the pressure, represented by $p_{{\scriptscriptstyle(0)}}$, are scalars, their gauge-invariant temperature and pressure perturbations can be defined in a manner analogous to that of the expressions (4):
+Therefore, Eq. (62b) can be regarded as the evolution equation for the entropy per particle.
+XII.2 Temperature and Pressure Perturbations
+Since the background temperature $T_{{\scriptscriptstyle(0)}}$ and pressure $p_{{\scriptscriptstyle(0)}}$ are scalars, their gauge-invariant perturbations are defined using the definition (7):
 $T^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=T_{\scriptscriptstyle(1)}-\dfrac{\dot{T}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad p^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=p_{\scriptscriptstyle(1)}-\dfrac{\dot{p}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}.$ (72)
+The expression for $T^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ will be used in Sec.
+XIII.2.
 In order to arrive at the gauge-invariant counterpart of (14a), it is first necessary to eliminate $\varepsilon_{{\scriptscriptstyle(1)}}$ and $n_{{\scriptscriptstyle(1)}}$ from (14a) using (6a).
-The utilization of $\dot{p}_{{\scriptscriptstyle(0)}}=p_{n}\dot{n}_{{\scriptscriptstyle(0)}}+p_{\varepsilon}\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and (68), results in
+Using $\dot{p}_{{\scriptscriptstyle(0)}}=p_{n}\dot{n}_{{\scriptscriptstyle(0)}}+p_{\varepsilon}\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and (72), yields
 $p^{\textnormal{phys}}_{\scriptscriptstyle(1)}=p_{n}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (73)
-Eliminating $p_{\varepsilon}$ with the aid of (40) and utilizing the definitions provided in (59) for density fluctuations results in the following expression:
+Eliminating $p_{\varepsilon}$ with the aid of (40) and using the definitions provided in (61) for relative density perturbations results in the following expression:
 $p_{\scriptscriptstyle(1)}^{\textnormal{phys}}=\beta^{2}\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}+n_{\scriptscriptstyle(0)}p_{n}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (74)
 The first term represents the adiabatic component of the pressure perturbation, while the second term denotes the nonadiabatic component.
-XII Application: Structure Formation in a Flat FLRW Universe
+XIII Application: Flat FLRW Universe
 The evolution equations presented in (62), are applicable to open, flat, and closed flrw universes.
-Henceforth, the discussion will be focused on the flat, ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, universe.
-Given that the initial density fluctuations occurred in the early universe when $\Lambda\ll\kappa\varepsilon_{\scriptscriptstyle(0)}$, the cosmological constant, $\Lambda$, will be neglected.
-Consequently, the background equations, as given in Eq.
-(10), are reduced to
+We will now focus on studying density perturbations in a flat flrw universe, for which ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$.
+Given that the initial density perturbations occurred in the early universe when $\Lambda\ll\kappa\varepsilon_{\scriptscriptstyle(0)}$, the cosmological constant $\Lambda$ will be neglected.
+Consequently, the background equations (11), are reduced to
 $3H^{2}$ $=\kappa\varepsilon_{\scriptscriptstyle(0)},\quad$ $\kappa$ $=8\pi G_{\text{N}}/c^{4},$ (75a)
@@
 $\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (75c)
-By employing the Friedmann equation (71a), the coefficients (61) of Eq. (60a) result in
+By using the Friedmann equation (75a), the coefficients (63) of Eq. (62a) result in
 $b_{1}$ $=H(1-3w-3\beta^{2})-2\dfrac{\dot{\beta}}{\beta},$ (76a)
@@
 where $\nabla^{2}$ represents the conventional Laplace operator.
-XII.1 Era before Decoupling of Matter and Radiation
+XIII.1 Era before Decoupling of Matter and Radiation
 In this era the primordial fluid is a mixture of radiation and matter, wherein the contribution of matter to the pressure is negligible.
-Consequently, the equations of state (see, for example, Ref.
-[8], Eq.
-(5.49), Ref.
-[10], Eq.
-(1.78), and Ref.
-[17], § V-1) are given by:
+Consequently, the equations of state are given by [see Ref.
+[13], § V-1]:
 $\varepsilon=a_{\text{B}}T_{\gamma}^{4}+nmc^{2},\quad p=\tfrac{1}{3}a_{\text{B}}T_{\gamma}^{4}.$ (77)
@@
 $p_{n}=-\tfrac{1}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{1}{3}.$ (79)
-For the parameter $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and the nonadiabatic speed of sound $\beta$, (40), the following expressions are derived:
+Since $p_{n}<0$, energy density perturbations are coupled to particle number density perturbations according to (65).
+For the parameter $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and the nonadiabatic speed of sound $\beta$, (40), it is found
 $w=\frac{\tfrac{1}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}},\quad\beta^{2}=\frac{1}{3}-\frac{\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{\tfrac{4}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}}.$ (80)
-The ensuing subsections provide a concise overview of the radiation-dominated and matter-dominated eras preceding decoupling.
-XII.1.1 Radiation-dominated Era
 The universe was radiation-dominated when
 $a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}\gg n_{{\scriptscriptstyle(0)}}mc^{2}.$ (81)
-In this case one has $\beta^{2}\approx\tfrac{1}{3}$, and $w\approx\tfrac{1}{3}$, implying that $\dot{\beta}\approx 0$ and $p_{\varepsilon}\approx w$.
-Upon substituting the aforementioned values and expressions (75) into the coefficients (72), the following result is obtained:
-$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}$
-$\qquad\qquad\qquad\;\,\approx\dfrac{-\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\dfrac{\nabla^{2}}{a^{2}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)},$ (78a)
-$\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}\approx-\tfrac{3}{4}H\dfrac{n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}.$ (78b)
-From (77), it can be deduced that the pressure term on the left-hand side of equation (78a) is, in absolute value, significantly larger than the source term.
-Consequently, the source term may be neglected.
-In the radiation-dominated phase, the system of equations (60) can be expressed as follows:
-$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}\approx 0,$ (79a)
-$\left|\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\right|\rightarrow 0.$ (79b)
-The limit in Eq. (79b) is a consequence of the fact that the coefficient in Eq. (78b) is negative, since $p_{n}<0$.
-In order to solve Eq. (79a), it is necessary to employ the solutions of the background equations (71). These solutions are provided by:
+In this case one has $\beta^{2}\approx\tfrac{1}{3}$, and $w\approx\tfrac{1}{3}$, implying that $\dot{\beta}\approx 0$.
+Upon substituting these values and expressions (79) into the coefficients (76), and using (81) the result is
+$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}=0.$ (82)
+In order to solve Eq. (82), it is necessary to use the solutions to the background equations (75). These solutions are given by
 $H=\tfrac{1}{2}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{3}{4}(ct)^{-2},\quad a\propto t^{1/2}.$ (83)
-The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{0}\geq 1$. This definition implies that
+The dimensionless time, denoted by the symbol $\tau$, is defined by $\tau:=t/t_{0}\geq 1$. This definition implies that
 $\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{0}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl(2H(t_{0})\bigr)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (84)
-By employing the Helmholtz equation $\nabla^{2}\delta_{\varepsilon}=-|\bm{q}|^{2}\delta_{\varepsilon}$ and the result derived in (81), it can be shown that the equation (79a) can be rewritten as
-$\delta_{\varepsilon}^{\prime\prime}-\dfrac{1}{2\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{\mu_{\text{r}}^{2}}{4\tau}+\dfrac{1}{2\tau^{2}}\right)\delta_{\varepsilon}=0.$ (82)
-In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{r}}$ is defined as follows:
+By using the Helmholtz equation $\nabla^{2}\delta_{\varepsilon}=-|\bm{q}|^{2}\delta_{\varepsilon}$ and the result derived in (84), it can be shown that the equation (82) can be rewritten as
+$\delta_{\varepsilon}^{\prime\prime}-\dfrac{1}{2\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{\mu_{\text{r}}^{2}}{4\tau}+\dfrac{1}{2\tau^{2}}\right)\delta_{\varepsilon}=0,$ (85)
+where the prime denotes differentiation with respect to the $\tau$. With $|\bm{q}|=2\pi/\lambda_{0}$, the parameter $\mu_{\text{r}}$ is defined by
 $\mu_{\text{r}}:=\dfrac{2\pi}{\lambda_{0}}\dfrac{1}{H(t_{0})}\dfrac{1}{\sqrt{3}}.$ (86)
-In this expression, $\lambda_{0}:=\lambda a(t_{0})$ represents the physical scale of a fluctuation at $t=t_{0}$, with the magnitude of the fluctuation given by $\lvert\bm{q}\rvert=2\pi/\lambda_{0}$.
+In this expression, $\lambda_{0}:=\lambda a(t_{0})$ represents the physical scale of a density perturbation at $t=t_{0}$.
 In order to solve Eq.
@@
 Upon transforming back to $\tau$, the general solution of Eq.
-(82) with constants of integration $A_{1}(\bm{q})$ and $A_{2}(\bm{q})$ is obtained, namely
-$\delta_{\varepsilon}(\tau,\bm{q})=\bigl{[}A_{1}(\bm{q})\sin\left(\mu_{\text{r}}\sqrt{\tau}\right)+A_{2}(\bm{q})\cos\left(\mu_{\text{r}}\sqrt{\tau}\right)\bigr{]}\sqrt{\tau},$ (84)
-It can be inferred from this solution that, during the radiation-dominated era of the universe, density fluctuations oscillated with an amplitude that grew in a manner proportional to the square root of time.
-In the case of large-scale fluctuations, that is, ${\lambda_{0}\rightarrow\infty}$, which leads to $\mu_{\text{r}}\rightarrow 0$, (83), the general solution of Eq. (82) is given by
+(85) with constants of integration $A_{1}(\bm{q})$ and $A_{2}(\bm{q})$ is obtained:
+$\delta_{\varepsilon}(\tau,\bm{q})=\bigl[A_{1}(\bm{q})\sin\left(\mu_{\text{r}}\sqrt{\tau}\right)+A_{2}(\bm{q})\cos\left(\mu_{\text{r}}\sqrt{\tau}\right)\bigr]\sqrt{\tau}.$ (87)
+Since $\vartheta_{{\scriptscriptstyle(1)}}$ is included in Eq.
+(82) it yields oscillating relative density perturbations with an increasing amplitude.
+In contrast, the standard equation (153), for which $\vartheta_{{\scriptscriptstyle(1)}}=0$ (i.e., no fluid flow) yields oscillating solutions with a constant amplitude.
+In the case of large-scale perturbations, that is, $\lambda_{0}\rightarrow\infty$}, which leads to $\mu_{\text{r}}\rightarrow 0$, (86), the general solution to Eq. (85) is given by
 $\delta_{\varepsilon}(\tau)=-\bigl[\delta_{\varepsilon}(1)-2\delta^{\prime}_{\varepsilon}(1)\bigr]\tau+2\bigl[\delta_{\varepsilon}(1)-\delta^{\prime}_{\varepsilon}(1)\bigr]\tau^{1/2}.$ (88)
-This solution, with the exception of the precise factors of proportionality, has been derived by a significant number of authors, as evidenced in Refs. [2, 39, 40, 41, 42].
-A comparison of the solutions (85) and (155a) reveals that the solution ${\propto\tau}$ is a consequence of the homogeneous part of equation (154a).
-The solution $\propto\tau^{1/2}$ is the particular solution of the inhomogeneous equation.
-This solution can only be obtained if the covariant divergence, $\vartheta_{{\scriptscriptstyle(1)}}$, is taken into account, as demonstrated in Appendix C.1.
-The minus sign preceding the solution $\propto\tau$ is a direct consequence of the fact that ${\vartheta_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}>0}$, as follows from (155b).
-The presence of the gauge function $\psi$ precludes the derivation of the solution (85) from (154).
-As demonstrated in Appendix A, $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the left-hand side of (78a), as evidenced by the derivation of (60a).
-XII.1.2 Matter-dominated Era
-The radiation-dominated phase came to an end and the matter-dominated phase commenced at $z(t_{\text{eq}})=3387$, as indicated in Table 1, well before decoupling [43], when
-$a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}=n_{{\scriptscriptstyle(0)}}mc^{2}.$ (86)
-Given that $p_{n}<0$, it can be shown from Eq. (60b) that
-$\left|\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right|\rightarrow 0.$ (87)
-It can thus be seen that fluctuations in the energy density are coupled to those in the particle number density.
-In light of the fact that the evolution equations (60) are general relativistic, it follows that the weak equivalence principle is valid.
-This implies that the evolution of fluctuations in the particle number density is independent of the composition of matter.
-Therefore, if cdm consists of particles that interact via gravitation, the fluctuations in the cdm are also linked to the fluctuations in the energy density.
-This implies that it is impossible for cdm to have contracted before decoupling.
-Consequently, cdm cannot have initiated the formation of structure after decoupling.
-XII.2 Era after Decoupling of Matter and Radiation
-Once protons and electrons have combined to form hydrogen, the radiation pressure will be negligible, and the equations of state will be those of a nonrelativistic monatomic perfect gas with three degrees of freedom.
-These are given by the equations of state [see Ref.
+A comparison of the solutions (88) and (155a) reveals that the solution proportional to $\tau$ in (88) is also a solution to the homogeneous part of equation (154a).
+The solution proportional to $\tau^{1/2}$ is the particular solution to the inhomogeneous equation.
+This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.1.
+The minus sign preceding the solution proportional to $\tau$ is a consequence of the fact that $\vartheta_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}>0$}, as follows from (155b).
+The presence of the gauge function $\psi(\bm{x})$ precludes the derivation of the solution (88) from (154).
+XIII.2 Era after Decoupling of Matter and Radiation
+Once protons and electrons have combined to form hydrogen, radiation pressure will be negligible and the equations of state will be those of a nonrelativistic, monatomic, perfect gas with three degrees of freedom.
+Therefore, pressure can be neglected in the background universe.
+However, the momentum conservation law (49) indicates that, in order to study the evolution of density perturbations, one must consider pressure perturbations, since density perturbations are static (51) in a pressureless fluid.
+According to the literature [see Ref.
+[32], Eq.
+(13), and Ref.
 [42], Eqs.
-(15.8.20)–(15.8.21), and Ref.
-[33], Eq.
-(13)]:
+(15.8.20)–(15.8.21)], we use the following equations of state for a nonrelativistic gas:
 $\varepsilon(n,T)=nmc^{2}+\tfrac{3}{2}nk_{\text{B}}T,\quad p(n,T)=nk_{\text{B}}T.$ (89)
@@
 The rest mass energy density is $nmc^{2}$ and the kinetic energy density is $\tfrac{3}{2}nk_{\text{B}}T$.
-The elimination of the temperature, $T$, from the equations of state (88), yields the equation of state for the pressure [see Ref.
-[31], Eq.
-(2.10.27)]:
+The elimination of the temperature $T$ from the equations of state (89), yields the equation of state for the pressure [see Ref. [42], Eq. (2.10.27)]:
 $p(n,\varepsilon)=\tfrac{2}{3}(\varepsilon-nmc^{2}).$ (90)
-The partial derivatives are determined by the utilization of the expressions (13b) and are given by
+The partial derivatives are determined using the expressions (14b) and are given by
 $p_{n}=-\tfrac{2}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{2}{3}.$ (91)
@@
 Upon substituting (89), (91) and (92) into (40), one arrives at the well-known result [see Ref. [42], Eq. (15.8.22)]:
-$\quad\beta\approx\dfrac{v_{\text{s}}}{c}=\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}}.$ (92)
-In this expression, $v_{\text{s}}$ represents the adiabatic speed of sound.
-From $\beta^{2}\approx\tfrac{5}{3}w$, it is found that Eq.
-(44) results in $\dot{w}\approx-2Hw$.
+$\quad\beta\approx\dfrac{v_{\text{s}}}{c}=\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}},$ (93)
+where $v_{\text{s}}$ represents the adiabatic speed of sound.
+From $\beta^{2}\approx\tfrac{5}{3}w$, we find from Eq.
+(44) that $\dot{w}\approx-2Hw$.
 Therefore, with $H:=\dot{a}/a$, it can be concluded that $w\propto a^{-2}$.
@@
 $T_{\scriptscriptstyle(0)}\propto a^{-2}.$ (94)
-In light of the relatively minor values of both $w$ and $\beta^{2}$, it is now possible to justify the equations of state (88).
-From the proportionality (93) and the fact that $n_{{\scriptscriptstyle(0)}}\propto a^{-3}$, as can be seen from Eq.
-(71c), it follows from (88) that ${p_{{\scriptscriptstyle(0)}}\propto a^{-5}}$.
-Given that pressure decayed rapidly and was already negligible in comparison to the energy density at the time of decoupling, it can be concluded that the unperturbed pressure had a minimal impact on structure formation.
-It is therefore reasonable to neglect the quantities $w$ and $\beta^{2}$ with respect to the constants of order one that appear in the momentum conservation laws (45).
-In doing so, one finds
-$\dot{u}^{i}_{\scriptscriptstyle(1)}+2Hu^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}-\dfrac{p_{{\scriptscriptstyle(1)}|j}}{\varepsilon_{\scriptscriptstyle(0)}}=0.$ (94)
-The second equation is derived by lowering the index $i$ in the first equation using the background metric $g_{{\scriptscriptstyle(0)}ij}$ and the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$.
-Accordingly, the evolution of density fluctuations is highly dependent upon the presence of pressure gradients, represented by $p_{{\scriptscriptstyle(1)}|j}$.
-In the event that the pressure gradients are identically zero, Eqs.
-(94) reduce to the form of Eqs.
-(49) and (50) in the nonrelativistic limit.
-This implies that the physical component of the fluid velocity is zero, and consequently, density perturbations cease to evolve.
-Given that structure formation in the universe was predominantly determined by pressure gradients, it is necessary to consider the kinetic energy density and pressure in addition to the energy density.
-This is the reason why the particle number density must be taken into account in the equations of state.
-In conclusion, it can be inferred from the equations (94) that the gauge modes (28) are zero.
-This result demonstrates that the relativistic gauge transformations (1), where $\xi^{\mu}$ is given by (6), are equivalent to the Newtonian gauge transformations (53).
-This conclusion is consistent with expectations, given that the equations of state (88) are nonrelativistic.
-The subsequent step is to derive the evolution equations for density perturbations.
+The subsequent step is to derive the evolution equations for relative density perturbations.
 The proportionality (94) implies with (93) that $\dot{\beta}/\beta=-H$.
@@
 $\dfrac{1}{c}\dfrac{\mathrm{d}}{{\mathrm{d}}t}\left(\delta_{n}-\delta_{\varepsilon}\right)=-2H\left(\delta_{n}-\delta_{\varepsilon}\right),$ (95b)
-In this calculation, the approximation is made that $w$ and $\beta^{2}$ are negligible with respect to the constants of order one.
+where it is used that $w$ and $\beta^{2}$ are negligible with respect to constants of order one.
 Given that the term $(\delta_{n}-\delta_{\varepsilon})$ occurs within the source term of Eq.
@@
 By making the substitution $H:=\dot{a}/a$, the following result is obtained:
-$\delta_{n}-\delta_{\varepsilon}\propto a^{-2}.$ (96)
-In order to relate this proportionality to thermodynamic quantities, it is necessary to express the perturbed equation of state for the energy density in terms of density fluctuations.
+$(\delta_{n}-\delta_{\varepsilon})\propto a^{-2}.$ (96)
+In order to relate this proportionality to thermodynamic quantities, it is necessary to express the perturbed equation of state for the energy density in terms of relative density perturbations.
 It can be deduced from $\varepsilon=\varepsilon(n,T)$ that
@@
 By multiplying (97a) by $\theta_{\scriptscriptstyle(1)}/\dot{\theta}_{\scriptscriptstyle(0)}$ and subtracting the result from (97b), one obtains the following result:
-$\varepsilon_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}.$ (98)
-This result is based on the use of expressions (4a) and (68). By employing the expression for $\varepsilon$ in (88) to eliminate the partial derivatives, one obtains the perturbed equation of state:
+$\varepsilon_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}^{{\textnormal{phys}}},$ (98)
+where expressions (6a) and (72) have been used. By using the expression for $\varepsilon$ in (89) to eliminate the partial derivatives, one obtains the perturbed equation of state:
 $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}mc^{2}+\tfrac{3}{2}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}k_{\text{B}}T_{\scriptscriptstyle(0)}+\tfrac{3}{2}n_{\scriptscriptstyle(0)}k_{\text{B}}T^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (99)
-By dividing expression (99) by $\varepsilon_{\scriptscriptstyle(0)}$, and using the exact value of the ratio $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, the perturbed equation of state for the energy density expressed in fluctuations (59) is obtained:
+Upon dividing (99) by $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}+\tfrac{3}{2}n_{{\scriptscriptstyle(0)}}k_{\text{B}}T_{{\scriptscriptstyle(0)}}$ and using the exact value of the ratio $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ (92), the perturbed equation of state for the energy density expressed in relative density perturbations (61) is obtained:
 $\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})=-\tfrac{3}{2}w(t)\delta_{T}(t,\bm{x}).$ (100)
-In this context, the quantity $\delta_{T}$ is defined by $\delta_{T}:=T^{\textnormal{phys}}_{\scriptscriptstyle(1)}/T_{\scriptscriptstyle(0)}$.
+The quantity $\delta_{T}$ is defined by $\delta_{T}:=T^{\textnormal{phys}}_{\scriptscriptstyle(1)}/T_{\scriptscriptstyle(0)}$.
 The solution to Eq.
-(95b) is obtained by combining (96) and (100) and using the approximate value of $w$, (91), which implies that $w\propto a^{-2}$.
-The result is as follows:
+(95b) is obtained by combining (96) and (100).
+Using the approximate value of $w$ implies that $w\propto a^{-2}$ [see (94)], making $\delta_{T}$ independent of time.
+Therefore, the solution to Eq.
+(95b) is given by
 $\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})\approx-\tfrac{3}{2}w(t)\delta_{T}(t_{\text{dec}},\bm{x}).$ (101)
-The quantity $\delta_{T}(t_{\text{dec}},\bm{x})$ represents the temperature fluctuation of matter at the time $t_{\text{dec}}$, which marks the decoupling of matter from radiation.
-Since $w\ll 1$, (101) is in alignment with (87).
-According to (67), (91) and (101) the entropy per particle is as follows:
+The quantity $\delta_{T}(t_{\text{dec}},\bm{x})$ represents the relative temperature perturbation of matter at the time $t_{\text{dec}}$, which marks the decoupling of matter from radiation.
+According to (71), (92) and (101) the entropy per particle is:
 $s^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})\approx\tfrac{3}{2}k_{\text{B}}\delta_{T}(t_{\text{dec}},\bm{x}).$ (102)
-Therefore, it can be concluded that the quantity $\delta_{T}$ is random and that density fluctuations are nonadiabatic if $\delta_{T}$ is nonzero.
-Upon substituting (90), (92), (101), and the approximation ${w\ll 1}$, the local pressure fluctuation, defined by $\delta_{p}:=p^{\textnormal{phys}}_{\scriptscriptstyle(1)}/p_{\scriptscriptstyle(0)}$, can be calculated from expression (70).
+Thus, the quantity $\delta_{T}$ is random. As a result, it is generally not equal to zero, which implies that density perturbations are generally nonadiabatic.
+Using (91), (93), and (101), the local relative pressure perturbation $\delta_{p}:=p^{\textnormal{phys}}_{\scriptscriptstyle(1)}/p_{\scriptscriptstyle(0)}$ can be calculated from expression (74).
 One obtains
 $\delta_{p}(t,\bm{x})\approx\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})+\delta_{T}(t_{\text{dec}},\bm{x}).$ (103)
-In this context, the terms $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ and $\delta_{T}(t_{\text{dec}},\bm{x})$ represent the adiabatic and nonadiabatic pressure fluctuations, respectively.
-The solution to Eq.
-(95b) has been determined.
-This allows for the reformulation of the second-order equation (95a) to facilitate analysis of the evolution of density fluctuations.
+In this context, the terms $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ and $\delta_{T}(t_{\text{dec}},\bm{x})$ represent the adiabatic and the random, nonadiabatic relative pressure perturbation, respectively.
+The solution (101) to Eq.
+(95b) allows for the reformulation of the second-order equation (95a).
 To that end, the solutions to the background equations (75) are required.
@@
 $H=\tfrac{2}{3}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{4}{3}(ct)^{-2},\quad a\propto t^{2/3}.$ (104)
-The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{\text{dec}}\geq 1$. This definition implies that
+The dimensionless time, denoted by the symbol $\tau$, is defined by $\tau:=t/t_{\text{dec}}\geq 1$. This definition implies that
 $\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{\text{dec}}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl(\tfrac{3}{2}H(t_{\text{dec}})\bigr)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (105)
-By employing Eqs.
-(71a) and (101), as well as the expressions (92), (104) (105), the Helmholtz equation $\nabla^{2}\delta=-|\bm{q}|^{2}\delta$, and $w\propto a^{-2}$, it can be demonstrated that Eq.
+By using Eqs.
+(75a) and (101), as well as the expressions (93), (104), (105), the Helmholtz equation $\nabla^{2}\delta=-|\bm{q}|^{2}\delta$, and $w\propto a^{-2}$, it can be demonstrated that Eq.
 (95a) can be expressed in the following manner:
-$\delta_{\varepsilon}^{\prime\prime}+\dfrac{2}{\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{4}{9}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}-\dfrac{10}{9\tau^{2}}\right)\delta_{\varepsilon}=-\dfrac{4}{15}\dfrac{\mu^{2}_{\text{m}}}{\tau^{8/3}}\delta_{T}(t_{\text{dec}},\bm{q}).$ (106)
-In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{m}}$ is defined as follows:
+$\delta_{\varepsilon}^{\prime\prime}+\dfrac{2}{\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{4}{9}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}-\dfrac{10}{9\tau^{2}}\right)\delta_{\varepsilon}=-\dfrac{4}{15}\dfrac{\mu^{2}_{\text{m}}}{\tau^{8/3}}\delta_{T}(t_{\text{dec}},\bm{q}),$ (106)
+where the prime denotes differentiation with respect to $\tau$. With $\lvert\bm{q}\rvert=2\pi/\lambda_{\text{dec}}$, the parameter $\mu_{\text{m}}$ is defined by
 $\mu_{\text{m}}:=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{H(t_{\text{dec}})}\dfrac{v_{\text{s}}(t_{\text{dec}})}{c}.$ (107)
-In this expression, $\lambda_{\text{dec}}:=\lambda a(t_{\text{dec}})$ represents the physical scale of a fluctuation at $t=t_{\text{dec}}$.
-The magnitude of the fluctuation is given by $\lvert\bm{q}\rvert=2\pi/\lambda_{\text{dec}}$.
-Furthermore, one has $v_{\text{s}}(t_{\text{dec}})/c\approx\beta(t_{\text{dec}})$, as stated in expression (92).
-Finally, the expression (103) enables the reformulation of Eq. (106) in a form suitable for the study of the evolution of density fluctuations:
+In this expression, $\lambda_{\text{dec}}:=\lambda a(t_{\text{dec}})$ represents the physical scale of a perturbation at $t=t_{\text{dec}}$, and $v_{\text{s}}(t_{\text{dec}})/c\approx\beta(t_{\text{dec}})$ is given by (93).
+Finally, the expression (103) enables the reformulation of Eq. (106) in a form suitable for the study of the evolution of relative density perturbations:
 $\delta^{\prime\prime}_{\varepsilon}+\frac{2}{\tau}\delta^{\prime}_{\varepsilon}+\dfrac{4}{15}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}\delta_{p}-\dfrac{10}{9\tau^{2}}\delta_{\varepsilon}=0,\quad\tau:=\dfrac{t}{t_{\text{dec}}}\geq 1.$ (108)
 The second term in this equation represents the expansion, the third term is the pressure term, where $\delta_{p}$ is given by (103), and the fourth term represents gravitation.
-In the limit of large scales, that is, $\lambda_{\text{dec}}\rightarrow\infty$, which leads to $\mu_{\text{m}}\rightarrow 0$, the solution of Eq. (108) is
+For large-scale perturbations, i.e., when $\lambda_{\text{dec}}\rightarrow\infty$, implying that $\mu_{\text{m}}\rightarrow 0$, relative pressure perturbations $\delta_{p}$ played a negligible role in their evolution.
+Therefore, the solution to Eq.
+(108) is
 $\delta_{\varepsilon}(\tau)=\bigl[\tfrac{5}{7}\delta_{\varepsilon}(1)+\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr]\tau^{2/3}+\bigl[\tfrac{2}{7}\delta_{\varepsilon}(1)-\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr]\tau^{-5/3}.$ (109)
-Therefore, pressure fluctuations $\delta_{p}$ played a negligible role in the evolution of the large-scale fluctuations.
-In other words, these fluctuations were almost adiabatic and evolved solely under the influence of gravitation and expansion.
-A comparison of the solutions (109) and (157a) reveals that the solution ${\propto\tau^{2/3}}$ is a consequence of the homogeneous part of Eq.
+The solution proportional to $\tau^{-5/3}$ dies out quickly.
+Thus, large-scale relative density perturbations evolve nearly proportionally to $\tau^{2/3}$, which is close to the solution (60) for a pressureless fluid.
+However, since $\vartheta_{\scriptscriptstyle(1)}^{\textnormal{phys}}(\tau)=-\tfrac{7}{9}\tau^{-4/3}$, (158b), there is a pressure-gradient-driven fluid flow.
+Therefore, the solution (109) is invalid for a pressureless fluid.
+As shown in Sec.
+X, the equations for scalar perturbations (47) are, for a pressureless fluid, reduced to a single equation, namely the time-independent Poisson equation (56).
+A comparison of the solutions (109) and (158a) reveals that the solution proportional to $\tau^{2/3}$ in (109) is also a solution to the homogeneous part of Eq.
 (157a).
-The solution $\propto\tau^{-5/3}$ is the particular solution to the inhomogeneous equation.
+The solution proportional to $\tau^{-5/3}$ is the particular solution to the inhomogeneous equation.
 This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.2.
-In contrast to the standard equation (150), the quantity $\vartheta_{{\scriptscriptstyle(1)}}$ is included in Eq.
-(95a).
+Since in the standard equation (156) $\vartheta_{{\scriptscriptstyle(1)}}=0$ (i.e., no fluid flow), it does not produce growing density perturbations.
 The presence of the gauge constant $C$, precludes the derivation of solution (109) from the system (157).
-XII.2.1 Cosmological Quantities
-To investigate the growth of density fluctuations, it is first necessary to express the parameter $\mu_{\text{m}}$ (107) in observable quantities.
+XIII.2.1 Cosmological Quantities
+Table 1: Planck satellite resultsy
+$z(t_{\text{dec}})$ = 1090
+$z(t_{\text{eq}})$ = 3387
+$cH(t_{\text{p}})$ = $67.66\,\text{km}\,\text{s}^{-1}\text{Mpc}^{-1}$
+$T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})$ = $2.725\,\text{K}$
+$t_{\text{p}}$ = $13.79\,\text{Gyr}$
+$\lvert\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})\rvert \lesssim$ $10^{-5}$
+To investigate the evolution of relative density perturbations using Eq.
+(108), the parameter $\mu_{\text{m}}$ (107) must first be expressed in terms of observable quantities, namely the redshift $z(t_{\text{dec}})$ at the time of decoupling, the present value of the Hubble parameter $cH(t_{\text{p}})$, and the present value of the temperature $T_{{\scriptscriptstyle(0)}}(t_{\text{p}})$ of the cosmological fluid.
 The redshift $z(t)$ is defined by the expression
@@
 where $a(t_{\text{p}})$ is the present value of the scale factor.
-In the case of a flat flrw universe, it is permissible to take $a(t_{\text{p}})=1$.
-Using the background solutions (104), it is possible to express the relevant parameters of the universe in the redshift.
-The result is
+The Friedmann equation (11a) shows that, for a flat flrw universe, one may set $a(t_{\text{p}})=1$.
+Using the background solutions (104), it is possible to express the relevant parameters of the universe in the redshift:
 $H(t)$ $=H(t_{\text{p}})\bigl(z(t)+1\bigr)^{3/2},$ (111a)
@@
 $n_{{\scriptscriptstyle(0)}}(t)$ $=n_{{\scriptscriptstyle(0)}}(t_{\text{p}})\bigl(z(t)+1\bigr)^{3}.$ (111d)
-In deriving relation (111c), it is utilised that, subsequent to decoupling, $T_{{\scriptscriptstyle(0)}\gamma}\propto a^{-1}$. The dimensionless time, defined as $\tau:=t/t_{\text{dec}}$, is given by
-$\tau=\left(\dfrac{z(t_{\text{dec}})+1}{z(t)+1}\right)^{3/2},$ (112)
-where relation (111b) with $z(t_{\text{p}})=0$ is used. Upon substituting the result (92) into expression (107) and utilizing the relations (111), the following is obtained:
+In deriving relation (111c), it is used that, subsequent to decoupling, $T_{{\scriptscriptstyle(0)}\gamma}\propto a^{-1}$.
+Upon substituting (93) into expression (107) and using the relations (111), we obtain
 $\mu_{\text{m}}=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{cH(t_{\text{p}})}\dfrac{1}{z(t_{\text{dec}})+1}\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})}{m}},$ (112)
@@
 In this manner, the parameter $\mu_{\text{m}}$ has been expressed in observable quantities.
-Upon substituting the numerical values from Tables 1 and 2, one obtains the result:
+Upon substituting the numerical values from Tables 1 and 2, one obtains
 $\mu_{\text{m}}=\dfrac{16.48}{\lambda_{\text{dec}}},\quad\lambda_{\text{dec}}\text{ in pc}.$ (113)
-It can thus be concluded that the parameter $\mu_{\text{m}}$ is dependent solely on the initial scale $\lambda_{\text{dec}}$ of a density fluctuation.
-Table 1: Planck satellite results
-$z(t_{\text{dec}})$ = 1090
-$z(t_{\text{eq}})$ = 3387
-$cH(t_{\text{p}})$ = $67.66\,\text{km}\,\text{s}^{-1}\text{Mpc}^{-1}$
-[…]
-[…]
-$\lvert\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})\rvert$ $\lesssim$ $10^{-5}$
-XII.2.2 Initial Values
-In order to solve Eq.
-(108), it is necessary to determine the initial values of the quantities $\delta_{\varepsilon}$ and $\delta^{\prime}_{\varepsilon}$.
-Furthermore, it is imperative to ascertain the values of the local nonadiabatic random pressure fluctuation $\delta_{T}$ in expression (103).
-Planck Satellite
-The Planck observations [43] of the fluctuations $\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})$ in the background radiation temperature imply that $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$.
-In the absence of knowledge regarding the initial growth rate, it is assumed that $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$.
-Therefore, the initial values for Eq.
-(108) are as follows:
-$\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\lesssim 10^{-5},\quad\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0.$ (115)
-As demonstrated in the calculations presented in Sec.
-XII.2.3, the outcome of Eq.
-(108) is largely independent of the initial value for $\delta_{\varepsilon}$, provided that it satisfies the condition of having a maximum value of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\leq 10^{-4}$.
-Nonadiabatic Pressure Fluctuations
-The values of the quantities $\delta_{\varepsilon}(t,\bm{q})$ and $\delta_{n}(t,\bm{q})$ are related to the local nonadiabatic random pressure fluctuation $\delta_{T}(t_{\text{dec}},\bm{q})$ by Eq.
-(101).
-Immediately following the decoupling, this equation reads:
-$\delta_{n}(t_{\text{dec}},\bm{q})-\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx-4.1\times 10^{-10}\delta_{T}(t_{\text{dec}},\bm{q}),$ (116)
-where it has been assumed that at decoupling, the matter temperature was equal to the radiation temperature, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$.
-Expressions (91) and (111c) were employed in conjunction with the values presented in Tables 1 and 2.
-The transition of the universe from the era preceding decoupling to the era following decoupling was rapid and chaotic.
-A notable decline was observed in the mean particle velocity, and the pressure.
-Furthermore, the values of both $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ exhibited minor irregularities on the surface of last scattering.
-Therefore, it is probable that after the transition, the initial values of the difference between fluctuations, represented by ${(\delta_{n}-\delta_{\varepsilon})(t_{\text{dec}},\bm{q})}$, were randomly distributed among all density fluctuations.
-From expression (116), it can be seen that very small differences between the initial energy density and particle number density fluctuations result in relatively large positive or negative fluctuations in the nonadiabatic component of the pressure fluctuations in expression (103).
-Therefore, the nonadiabatic pressure fluctuations, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, are randomly (102) distributed among the multitude of density fluctuations.
+It can thus be concluded that the parameter $\mu_{\text{m}}$ is dependent solely on the initial scale $\lambda_{\text{dec}}$ of a density perturbation.
+Finally, the dimensionless time, defined as $\tau:=t/t_{\text{dec}}$, is expressed in terms of the redshift, given by:
+$\tau=\left(\dfrac{z(t_{\text{dec}})+1}{z(t)+1}\right)^{3/2},$ (114)
+where relation (111b) with $z(t_{\text{p}})=0$ is used.
 Table 2: Physical constants
@@
 $a_{\text{B}}$ = $7.5657\times 10^{-16}\,\text{J}\,\text{m}^{-3}\,\text{K}^{-4}$
-XII.2.3 Structure Formation
-The subsequent analysis will employ Eq.
-(108) to examine the evolution of density fluctuations.
-This equation demonstrates that the evolution of a fluctuation $\delta_{\varepsilon}(t,\bm{q})$ with an initial scale $\lambda_{\text{dec}}$, was influenced by three key factors: the expansion of the universe, fluctuations in pressure, and gravitation.
-Fluctuations in pressure, as indicated by (103), were found to consist of two discrete components: an adiabatic contribution, represented by $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{q})$, and a nonadiabatic contribution, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, where the adiabatic component was initially negligible (115) and the nonadiabatic component was observed to be random.
-[…]
+XIII.2.2 Initial Values
+In order to solve Eq.
+(108), it is necessary to determine the initial values of the quantities $\delta_{\varepsilon}$ and $\delta^{\prime}_{\varepsilon}$.
+In addition, values of the local random, nonadiabatic relative pressure perturbation $\delta_{T}$ in expression (103), are required.
+Planck Satellite
+The Planck observations [50] of the relative perturbations $\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})$ in the background radiation temperature imply that $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$.
+In the absence of knowledge regarding the initial growth rate, it is assumed that $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$.
+Therefore, the initial values for Eq.
+(108) are as follows:
+$\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\lesssim 10^{-5},\quad\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0.$ (115)
+As demonstrated in the calculations presented in Sec.
+XIII.2.3, the outcome of Eq.
+(108) is almost entirely independent of the initial value for $\delta_{\varepsilon}$, provided that it satisfies the condition $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\leq 10^{-4}$.
+Random Nonadiabatic Pressure Perturbations
+The initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ are related to the local random, nonadiabatic relative pressure perturbation $\delta_{T}(t_{\text{dec}},\bm{q})$ by expression (101).
+Since $w\ll 1$}, very small values of the difference $\delta_{n}(t_{\text{dec}},\bm{q})-\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ result in relatively large positive or negative values of $\delta_{T}(t_{\text{dec}},\bm{q})$.
+The fact that $\delta_{T}(t_{\text{dec}},\bm{q})$ takes on random values follows from expression (102).
+However, using Eq.
+(101) the randomness of $\delta_{T}(t_{\text{dec}},\bm{q})$ can also be explained as follows.
+The transition of the universe from the era preceding decoupling to the era following decoupling was rapid and chaotic.
+A notable decline was observed in the mean particle velocity, and the pressure.
+As a consequence, the values of both $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ exhibited minor irregularities on the surface of last scattering.
+Therefore, it is probable that after the transition, the initial values of the difference between the relative density perturbations, represented by $(\delta_{n}-\delta_{\varepsilon})(t_{\text{dec}},\bm{q})$}, were randomly distributed among all density perturbations.
+XIII.2.3 Structure Formation
+The subsequent analysis will use Eq.
+(108) to examine the evolution of relative density perturbations.
+This equation shows that the evolution of a relative perturbation, $\delta_{\varepsilon}(t,\bm{q})$, with an initial scale $\lambda_{\text{dec}}$, was influenced by three factors: the expansion of the universe, pressure perturbations, and gravitation.
+Relative pressure perturbations $\delta_{p}$, as indicated by (103), consisted of two discrete components: an adiabatic contribution, represented by $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{q})$, and a random (102), nonadiabatic contribution, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, where the adiabatic component was initially negligible (115).
+Figure 1: The graphs show the redshift and time when a relative perturbation in the energy density with initial scale $\lambda_{\text{dec}}$, and initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$ starting to grow at an initial redshift of $z(t_{\text{dec}})=1090$ has become nonlinear, i.e., when $\delta_{\varepsilon}(t,\bm{q})=1$.
+The graphs are labeled with the initial values of the random, nonadiabatic pressure perturbations $\delta_{T}(t_{\text{dec}},\bm{q})$.
+For each graph, the Jeans scale is $6.4\,\text{pc}$.
+Since $w\ll 1$, (92), the nonadiabatic component could have taken on relatively large positive or negative values for small differences in the relative perturbations of energy and particle number density, as shown in (101).
+Consequently, there were regions where the total relative pressure perturbation was initially negative after decoupling.
+This allowed density perturbations to grow rapidly, despite the fact that the expansion of the universe opposed growth.
+As the adiabatic component of the relative pressure perturbation increased, the total relative pressure perturbation became positive.
+This significantly reduced the growth rate of density perturbations.
+The initial growth phase was brief yet long enough for a density perturbation to reach the nonlinear regime several hundred million years after the Big Bang.
+Conversely, if the nonadiabatic relative pressure perturbation was positive initially, then the total relative pressure perturbation was also positive.
+This resulted in the formation of voids, which were regions with negative values of $\delta_{\varepsilon}$.
+In these regions, matter was driven to the edges.
 Fig.
-1 provides a clear summary of the evolution of density fluctuations.
+1 provides a summary of the evolution of relative density perturbations.
 It was created using the following methodology.
-For each value of ${\delta_{T}=-0.01}$, $-0.05$, $-0.10$, $-0.15$, and $-0.20$, Eq.
-(108) is numerically solved for a large number of values for the initial fluctuation scale $\lambda_{\text{dec}}$ using the initial values (115).
-[…]
-[…]
-[…]
-$d_{\text{H}}(t_{\text{dec}})=a(t_{\text{dec}})\int_{0}^{ct_{\text{dec}}}\!\!\dfrac{\mathrm{d}\tau}{a(\tau)}=3ct_{\text{dec}}\approx 3.5\times 10^{5}\,\text{pc}.$ (117)
-This value has been calculated using Eq. (111b) and Tables 1 and 2.
+For each value of $\delta_{T}=-0.01$}, 0.05-0.05, 0.10-0.10, 0.15-0.15, and 0.20-0.20, Eq.
+(108) was solved numerically for a large number of values for the initial perturbation scale $\lambda_{\text{dec}}$ using the initial values (115).
+The integration started at $z=1090$ and ended when either $z=0$ or when $\delta_{\varepsilon}=1$.
+Each integration run returned a single point on the graph for a particular choice of the initial scale $\lambda_{\text{dec}}$ when $z>0$ and $\delta_{\varepsilon}=1$.
+In this case, it is clear that the perturbation has become nonlinear within $13.79\,\text{Gyr}$.
+On the other hand, if the integration stopped at $z=0$ and $\delta_{\varepsilon}<1$, then the perturbation had not yet reached its nonlinear phase today.
+Each graph shows the time and scale at which $\delta_{\varepsilon}=1$ for a given value of $\delta_{T}$.
+As illustrated in Fig.
+1, the optimal scale for growth was approximately $6.4\,\text{pc}$.
+Perturbations at this scale entered the nonlinear phase as early as $40$ million years ($z\approx 49$) after the Big Bang.
+Perturbations with scales smaller than $6.4\,\text{pc}$ reached their nonlinear phase at a much later time because their internal gravitation was weaker than for large-scale perturbations.
+Furthermore, pressure perturbations and expansion of the universe led to oscillatory behavior, as shown in Fig.
+2.
+In contrast, perturbations with a scale greater than $6.4\,\text{pc}$ were less affected by pressure perturbations.
+However, because of their larger size, the expansion worked against their growth, so that, despite their stronger gravitation, they also reached the nonlinear phase at a later time.
+Perturbations larger than $70\,\text{pc}$ did not reach their nonlinear regime within $13.79\,\text{Gyr}$.
+Perturbations that became nonlinear within $13.79\,\text{Gyr}$ are within the particle horizon at decoupling given by:
+$d_{\text{H}}(t_{\text{dec}})=a(t_{\text{dec}})\int_{0}^{ct_{\text{dec}}}\!\!\dfrac{\mathrm{d}\tau}{a(\tau)}=3ct_{\text{dec}}\approx 3.5\times 10^{5}\,\text{pc},$ (116)
+where Eq. (111b) and Tables 1 and 2 have been used.
 Figure 2: The graphs show the growth rates $\delta^{\prime}_{\varepsilon}$, with initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$, as function of the redshift $z$, or time in million of years.
-The initial scales $\lambda_{\text{dec}}$ of the fluctuations are measured in parsec.
-The evolution of density fluctuations started at $z=1090$.
-[…]
-[…]
+The initial scales $\lambda_{\text{dec}}$ of the perturbations are measured in parsec.
+The evolution of relative density perturbations started at $z=1090$.
+As Fig.
+2 shows, the largest initial growth rates occurred for perturbations that were smaller than $6.4\,\text{pc}$.
+The smallest perturbations which had become nonlinear within $13.79\,\text{Gyr}$, had a scale of approximately $2\,\text{pc}$, as shown in Fig.
+1.
+Perturbations with scales $2\,{\text{pc}}\lesssim\lambda_{\text{dec}}\lesssim 4.1\,\text{pc}$} oscillated toward the nonlinear phase within $13.79\,\text{Gyr}$.
+After approximately $14$ million years, the total pressure perturbation (103) had become positive and the growth rate had decreased, resulting in gravitation and, to a lesser extent, expansion becoming the primary drivers of the evolution of density perturbations.
+Consequently, the most turbulent phase of density perturbations had concluded, and a phase of gradual and consistent gravitational growth toward the nonlinear regime had commenced.
+As can be seen in Fig.
+1, at a redshift of $z\approx 7$, or $600$ million years after the Big Bang, all perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$ had become nonlinear.
+Assuming a spherical density perturbation with a diameter equal to $\lambda_{\text{dec}}$, the mass of the perturbation at decoupling is given by
 $M(t_{\text{dec}})=\dfrac{4\pi}{3}\bigl(\tfrac{1}{2}\lambda_{\text{dec}}\bigr)^{3}n_{\scriptscriptstyle(0)}(t_{\text{dec}})m.$ (117)
-It is first necessary to determine the particle number density at the time of decoupling denoted by $n_{\scriptscriptstyle(0)}(t_{\text{dec}})$, in order to proceed.
-This quantity is derived from the Friedmann equation (71a).
+The particle number density at the time of decoupling, $n_{\scriptscriptstyle(0)}(t_{\text{dec}})$, can be derived from the Friedmann equation (75a).
 For $w\ll 1$ and at time $t_{\text{p}}$ this equation is given by:
 $3H^{2}(t_{\text{p}})=\kappa n_{{\scriptscriptstyle(0)}}(t_{\text{p}})mc^{2},\quad\kappa=8\pi G_{\text{N}}/c^{4}.$ (118)
-By employing the relation (111d), one obtains
+By using the relation (111d), one obtains
 $n_{\scriptscriptstyle(0)}(t_{\text{dec}})=\frac{3\bigl[cH(t_{\text{p}})\bigr]^{2}\bigl[z(t_{\text{dec}})+1\bigr]^{3}}{8\pi mG_{\text{N}}}.$ (119)
-[…]
-[…]
+The relations (117) and (119) and the constants in Tables 1 and 2 reveal that perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$, had masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$.
+Because of the steepness of the graphs in Fig. 1 for scales below $6.4\,\text{pc}$, this scale is designated as the relativistic Jeans scale. Its value is
 $\lambda_{\text{Jeans}}(t_{\text{dec}}):=\lambda_{\text{Jeans}}a(t_{\text{dec}})\approx 6.4\,\text{pc}.$ (120)
-Accordingly, the Jeans mass at decoupling is thus given by
+Accordingly, the Jeans mass at decoupling is given by
 $M_{\text{Jeans}}(t_{\text{dec}})\approx 2.2\times 10^{4}\,\text{M}_{\odot}.$ (121)
-As evidenced by the preceding analysis, fluctuations in density, with initial values of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$ and $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$, result in the formation of structures in a flat flrw universe, provided that pressure fluctuations are taken into account.
-This marks the conclusion of the analysis concerning the evolution of density fluctuations in a flat flrw universe.
-XIII Summary and Conclusion
-This paper addresses four issues in the field of cosmological perturbation theory for Friedmann-Lemaître-Robertson-Walker (flrw) universes.
-The first and most significant issue concerns the gauge problem.
-It is well-established that the linearized Einstein equations and conservation laws yield solutions that are both physical and nonphysical.
-Consequently, the imposition of physical initial conditions does not result in solutions with physical significance.
-To resolve this issue, it is crucial to consider gauge-invariant quantities, that is, quantities that are independent of the choice of a coordinate system.
-It has been demonstrated that perturbations in the energy density and the particle number density can be constructed in a unique manner.
-In the nonrelativistic limit, it is demonstrated that these quantities represent the actual physical perturbations of the energy density and particle number density in the general theory of relativity.
-Similarly, as has been done for density perturbations, gauge-invariant perturbations in temperature, pressure, and entropy can be constructed.
-This addresses the gauge issue in the field of cosmology.
-The second issue is that of the growth of minor density fluctuations subsequent to the decoupling of matter and radiation.
-In a fluid with zero pressure, density fluctuations do not evolve.
-To address this issue, it is necessary to incorporate the particle number density in addition to the energy density in the linearized Einstein equations and conservation laws.
-This allows for the kinetic energy density and pressure to be accounted for, resulting in the emergence of local pressure gradients.
-These gradients give rise to local fluid flows that are instrumental in the formation of structures after decoupling.
-A third result is the novel perturbation theory (60), which provides a clear and unambiguous description of the evolution of fluctuations in the particle number density the energy density, as well as their interaction in a perfect fluid where the pressure depends on both the particle number density and the energy density.
-The aforementioned equations are not only applicable to a flat flrw universe, but also to open and closed universes.
-Furthermore, the cosmological constant need not be zero.
-It has been demonstrated that Eqs.
-(60) yield all previously established physical solutions for large-scale density perturbations in a flat flrw universe in its radiation-dominated phase and in the era following the decoupling of matter and radiation.
-In the development of the new cosmological perturbation theory, it has been demonstrated that in the nonrelativistic limit, relativistic gauge transformations reduce to Newtonian gauge transformations.
-Consequently, the conventional relativistic evolution equation—whether derived from general relativity or Newtonian gravitation modified to account for the expansion of the universe—also exhibits gauge-dependent solutions.
-This finding indicates that Newtonian gravitation is inadequate in accurately describing the evolution of density perturbations.
-[…]
-The investigation into cosmological density perturbations in a flat flrw universe reveals that if a nonrelativistic monatomic perfect gas with three degrees of freedom accurately represents the cosmic fluid in our universe following the decoupling of matter and radiation, then a perturbation theory based on general relativity provides an explanation for the origin of the first structures in the universe that does not rely on cdm.
-In contrast, the theory considers the local nonadiabatic random pressure fluctuations that emerged during the turbulent transition to the era that followed the decoupling of matter and radiation.
-Subsequent observations with the James Webb Space Telescope are expected to yield novel insights into the conditions that ensued immediately after the decoupling of radiation and matter, as well as the formation of the earliest structures in the universe.
-Appendix A Derivation of the Evolution Equations
+The above analysis shows that, if nonadiabatic pressure perturbations are taken into account, relative density perturbations with initial values $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$ and $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$} can give rise to structure in the early universe after the decoupling of matter and radiation.
+XIV Summary and Conclusion
+This article introduces a relativistic cosmological perturbation theory (62) that describes the evolution of relative energy density and particle number density perturbations in closed ($K=+1$), flat ($K=0$), and open (K=1{K=-1}) Friedmann-Lemaître-Robertson-Walker (flrw) universes, where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=-6K/a^{2}$ is the spatial curvature.
+It consists of two differential equations: a second-order inhomogeneous evolution equation for the relative energy density perturbation and a first-order entropy equation that determines the evolution of the source term of the first equation.
+These equations together describe the evolution of nonadiabatic relative perturbations $\delta_{\varepsilon}(t,\bm{x}):=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})/\varepsilon_{{\scriptscriptstyle(0)}}(t)$ and $\delta_{n}(t,\bm{x}):=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})/n_{{\scriptscriptstyle(0)}}(t)$ in energy and particle number densities, respectively.
+The entropy equation implies that before matter and radiation decoupled, both ordinary matter perturbations and Cold Dark Matter (cdm) perturbations are coupled to energy density perturbations by gravitation.
+Consequently, cdm did not initiate structure formation after decoupling.
+After decoupling, pressure gradients became essential for fluid flow, which is crucial for forming structures, since in a pressureless fluid, density perturbations $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ are static.
+Newtonian gravity is not a suitable framework for studying the evolution of cosmological density perturbations.
+This is because the conventional evolution equation, whether derived from general relativity or Newtonian gravity, yields gauge-dependent solutions.
+Consequently, no distinction exists between sub- and super-horizon density perturbations.
+XIV.1 Results for a Flat FLRW Universe
+In a radiation-dominated universe, relative density perturbations $\delta_{\varepsilon}$ oscillated with an amplitude that increased proportionally to the square root of time.
+After matter and radiation decoupled at $t=t_{\text{dec}}$, the cosmological fluid became a nonrelativistic perfect fluid.
+According to observations made by the Planck satellite [50], relative density perturbations were of the order of $10^{-5}$ at the time of decoupling.
+These perturbations are far too small to explain structures in the universe by gravitational growth, if only adiabatic perturbations are considered.
+This is because the growth of an adiabatic density perturbation slows down due to the increasing pressure perturbation caused by the density perturbation itself.
+For this reason, this article examines nonadiabatic density perturbations.
+This was achieved using an equation of state for a nonrelativistic perfect gas that considers not only the particles’ rest energy, but also their kinetic energy.
+Therefore, in addition to the energy density, the particle number density must also be considered.
+Then, the total pressure is given by $\delta_{p}(t,\bm{x})\approx\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})+\delta_{T}(t_{\text{dec}},\bm{x})$, where $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ is the adiabatic component caused by the density perturbation itself, and a random nonadiabatic component $\delta_{T}(t_{\text{dec}})$.
+In addition, the solution to the entropy equation is identical to the perturbed equation of state of the energy density, expressed in terms of relative density perturbations and the nonadiabatic component of pressure, that is, $\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})\approx-\tfrac{3}{2}w(t)\delta_{T}(t_{\text{dec}},\bm{x})$, where $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}\ll 1$.
+This expression shows that small random fluctuations in both $\delta_{n}(t_{\text{dec}},\bm{x})$ and $\delta_{\varepsilon}(t_{\text{dec}},\bm{x})$ — arising from the rapid and chaotic transition to the post-decoupling era — could lead to significant local random fluctuations in the relative nonadiabatic pressure perturbation, $\delta_{T}(t_{\text{dec}},\bm{x})$, which could take on both positive and negative values.
+If the total relative pressure perturbation $\delta_{p}(t_{\text{dec}},\bm{x})$ was positive at the moment of decoupling, a void was formed and matter moved toward its edges.
+Conversely, if $\delta_{p}(t_{\text{dec}},\bm{x})$ was negative, the density perturbation would undergo a brief period of rapid growth.
+This rapid growth continued until the total relative pressure perturbation $\delta_{p}(t,\bm{x})$ became positive.
+After that, the density perturbation slowly but steadily grew toward the nonlinear phase, which was reached early on.
+Perturbations smaller than $70\,\text{pc}$ would reach their nonlinear phase within $13.79$ billion years or less.
+The relativistic Jeans scale was $6.4\,\text{pc}$, and the corresponding Jeans mass was $2.2\times 10^{4}\,\text{M}_{\odot}$.
+Perturbations of Jeans scale size reached their nonlinear regime at a redshift of $z\approx 49$ or approximately $40$ million years after the Big Bang.
+All density perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$ reached their nonlinear phase within $z\approx 7$ or $600$ million years.
+These scales correspond to masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$.
+This is smaller than what was observed by the James Webb Space Telescope (jwst) [51], which observed mature galaxies.
+Nevertheless, it is reasonable to assume that the masses continued to increase in the nonlinear phase due to their greater gravitational pull.
+XIV.2 Conclusion
+Investigating cosmological density perturbations in a flat flrw universe reveals that if a nonrelativistic, monatomic, perfect gas with three degrees of freedom accurately represents the cosmological fluid in our universe after matter and radiation decoupled, then a perturbation theory based on general relativity can explain the origin of the first structures in the universe without relying on cdm.
+Subsequent observations with the jwst, launched on Christmas Day $2021$, are expected to yield novel insights into the conditions that ensued immediately after the decoupling of radiation and matter, as well as the formation of the earliest structures in the universe.
+These observations will help us determine if general relativity accurately describes the formation of the first structures in the universe.
+Appendix A Derivation of the
+Perturbation Theory
 Table 3: The coefficients $\alpha_{ij}$ figuring in the equations (124)
@@
 $\dfrac{\kappa n_{\scriptscriptstyle(0)}}{2H}$ $3H$ $n_{\scriptscriptstyle(0)}$ $\dfrac{n_{\scriptscriptstyle(0)}}{4H}$
-$\dfrac{p_{\varepsilon}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $\dfrac{p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $H(2-3\beta^{2})$ 00
-$\dfrac{\kappa\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}$ 00 $-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $2H+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{6H}$
-$\dfrac{-\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ 00 $\dfrac{6\varepsilon_{\scriptscriptstyle(0)}H(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ $\dfrac{\tfrac{3}{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$
-In this appendix, the perturbation equations (60) are derived. Prior to undertaking the subsequent analysis, it is necessary to carry out two preliminary steps.
+$\dfrac{p_{\varepsilon}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $\dfrac{p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $H(2-3\beta^{2})$ 0
+$\dfrac{\kappa\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}$ 0 $2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $2H+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{6H}$
+$\dfrac{-\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ 0 $\dfrac{6\varepsilon_{\scriptscriptstyle(0)}H(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ $\dfrac{\tfrac{3}{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$
+In this appendix, the perturbation theory (62) is derived. Before undertaking this task, it is necessary to carry out two preliminary steps.
 Firstly, as $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$, (6b), it is unnecessary to consider the gauge-dependent quantity $\theta_{{\scriptscriptstyle(1)}}$.
-As a result, the system (37) can be rewritten in a more suitable form.
-By eliminating the variable $\theta_{\scriptscriptstyle(1)}$ from the differential equations  (37b)–(37e) using the algebraic constraint equation (37a), a system of four first-order ordinary differential equations is obtained:
+Eliminating the quantity $\theta_{{\scriptscriptstyle(1)}}$ from the differential equations (37b)–(37e) using the algebraic constraint equation (37a) yields a system of four first-order ordinary differential equations suitable for computer algebra software:
 $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})$
@@
 Secondly, the energy density and particle number density perturbations (6a) are expressed in the four quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$.
-Employing the background equations (10) to eliminate all time derivatives and the linearized constraint equation (37a) to eliminate $\theta_{\scriptscriptstyle(1)}$, the following results are obtained:
+Using the background equations (11) to eliminate all time derivatives, and the linearized constraint equation (37a) to eliminate $\theta_{\scriptscriptstyle(1)}$, the following results are obtained:
 $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\dfrac{\varepsilon_{\scriptscriptstyle(1)}{{}^{3}\!}R_{\scriptscriptstyle(0)}-3\varepsilon_{\scriptscriptstyle(0)}(1+w)(2H\vartheta_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (123a)
@@
 Subsequently, Eqs.
-(60) for the fluctuations (59) will be derived in Sec.
+(62) for the relative perturbations (61) will be derived in Sec.
 A.2.
-The expression (44) will be utilized to eliminate the time derivative of $w$.
+The expression (44) will be used to eliminate the time derivative of $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$.
 A.1 Evolution Equation for Energy Density Perturbations
@@
 Step 2.
-In a manner analogous to the approach undertaken in Step 1, the explicit appearance of $\vartheta_{{\scriptscriptstyle(1)}}$ is removed from the system of equations (129).
+In a manner analogous to the approach undertaken in Step 1, the explicit occurrence of $\vartheta_{{\scriptscriptstyle(1)}}$ is removed from the system of equations (128).
 Differentiating Eq.
@@
 $\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{1}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{2}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=a_{3}\left(n_{\scriptscriptstyle(1)}+\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}\varepsilon_{\scriptscriptstyle(1)}\right),$ (134)
-where the coefficients $a_{1}$, $a_{2}$ and $a_{3}$ are given by
-$a_{1}=-\dfrac{s_{3}}{\gamma_{53}}+\dfrac{r}{q_{5}},\quad a_{2}=-\dfrac{s_{3}}{\gamma_{53}q_{5}},\quad a_{3}=\dfrac{\gamma_{52}s_{3}}{\gamma_{53}q_{5}}-\dfrac{s_{2}}{q_{5}}.$ (136)
 As a consequence of the gauge-invariance of the quantity $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, the left-hand side of Eq.
@@
 $\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}=-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}.$ (135)
-It can be seen that the conservation laws (10c) and (10d) and the expressions (4a) lead to the following conclusion:
+The proof of this equality is straightforward but requires significant computational effort.
+The algebraic task was performed using the computer algebra system Maxima [52].
+The conservation laws (11c) and (11d) and the expressions (6a) yield:
 $n_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (136)
-It follows from expressions (137) and (138) that the right-hand side of Eq.
-(135) does not explicitly contain the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$.
-Consequently, it is gauge-invariant.
-Therefore, equation (135) governs the evolution of $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
-The proof of the equality (137) is straightforward but requires significant computational effort. The algebraic task was performed using the computer algebra system Maxima [44].
+Substituting (135) into (134) and using (136), we arrive at the evolution equation for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$:
+[…]
+where the coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by
+$a_{1}=-\dfrac{s_{3}}{\gamma_{53}}+\dfrac{r}{q_{5}},\quad a_{2}=-\dfrac{s_{3}}{\gamma_{53}q_{5}},\quad a_{3}=\dfrac{\gamma_{52}s_{3}}{\gamma_{53}q_{5}}-\dfrac{s_{2}}{q_{5}}.$ (138)
+The derivation of Eq.
+(137) shows that both the perturbation ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ (19) of the spatial curvature and the divergence $\vartheta_{{\scriptscriptstyle(1)}}$ (20) of the fluid three-vector are included in this equation.
 The final step in this process will be deriving an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
@@
 From this system, a second-order equation (137) for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ has been derived.
-It thus follows that the remaining system from which an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived is at most of second order.
-As the gauge-invariant quantities, namely, $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, have been employed, one degree of freedom, specifically the gauge function $\psi(\bm{x})$ in (147), has been eliminated.
+It thus follows that the remaining system (133a)–(133b), and (133e) from which an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived is at most of second order.
+As the gauge-invariant quantities, namely $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, have been used, one degree of freedom, specifically the gauge function $\psi(\bm{x})$ in (147), has been eliminated.
 Consequently, only a first-order evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived.
-In place of deriving an equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, an evolution equation will be derived for the expression (138), which contains $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
-By differentiating the left-hand side of equation (138) with respect to time and employing the background equations (10c) and (10d), the first-order equations (123a) and (123b) and the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$, the following result is obtained:
+In place of deriving an equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, an evolution equation will be derived for the right-hand side of expression (136).
+By differentiating the left-hand side of this expression with respect to time and using the background equations (11c) and (11d), the first-order equations (122a) and (122b) and the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$, the following result is obtained:
 $\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left(n_{\scriptscriptstyle(1)}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}\right)$
@@
 (A.1) by $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
-A.2 Evolution Equations for Density Fluctuations
+A.2 Perturbation Theory
 First Eq.
 (62a) will be derived.
-Upon substituting the expression $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}$ in Eq.
-(135) and dividing by $\varepsilon_{\scriptscriptstyle(0)}$, the result is
+Substituting the expression $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}$ into Eq.
+(137) and dividing by $\varepsilon_{\scriptscriptstyle(0)}$ yields the following result:
 $b_{1}=2\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1},\quad b_{2}=\dfrac{\ddot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1}\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{2},\quad b_{3}=a_{3}\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}.$ (140)
-The coefficients $a_{1}$, $a_{2}$ and $a_{3}$ are given by (136). With the aid of the computational software Maxima [44], the coefficients (61) of the main text have been calculated.
+The coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by (138). With the aid of the computational software Maxima [52], the coefficients (63) of the main text have been calculated.
 Finally, Eq. (62b) will be derived. From the definitions (61), it can be deduced that
 $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}=n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (141)
-By employing the equality (138) and substituting the expression (141) into Eq. (A.1), the following is obtained:
+By using the equality (136) and substituting the expression (141) into Eq. (A.1), the following is obtained:
 $\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right]$
 $=-3H\left(1-\frac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right].$ (142)
-Derivation of Eq. (60b) of the main text was accomplished through the utilization of Eq. (10d).
+Using Eq. (11d) yields Eq. (62b) of the main text.
 Appendix B Verifying the Evolution Equations for Scalar Perturbations
-The present paper does not provide a derivation of the linearized equations (12).
-A demonstration of the validity of the system of equations (37), which is derived from the aforementioned linearized equations and is fundamental to the new perturbation theory, will therefore be given via three different methods.
-The initial step in this process is to utilize the well-established property of the Einstein equations and conservation laws.
-This implies that if the constraint equations are satisfied at an initial time, then the solutions of the dynamical equations and conservation laws will satisfy the constraint equations for all subsequent times.
-A more detailed discussion can be found in Weinberg’s book [31], in Sec.
-7.5 on the Cauchy problem.
-This is also true in reverse: if solutions of the constraint equations and conservation laws can be found for all times, they must necessarily satisfy the dynamical equations.
-The system (37) comprises the conservation laws and constraint equations, and consists of four first-order ordinary differential equations and one algebraic equation.
-It thus follows that the system (37), can be solved without the use of the dynamical equations, since the latter are inherently satisfied by the solution of the aforementioned system.
-Accordingly, the latter of the two observations will be employed to substantiate the validity of equations (37).
-The off-diagonal ($i\neq j$) equations of (12c) are not associated with scalar perturbations.
-For the diagonal ($i=j$) equations, it is sufficient to consider the contraction of the dynamical equations (12c):
+This article does not present a derivation of the linearized equations (13), from which the system of equations (37) for scalar perturbations is derived.
+Since the latter system is key to the cosmological perturbation theory developed in this article, we will demonstrate its validity.
+This will be accomplished using three different methods.
+The first step in this process is to apply a well-known property of Einstein’s equations and conservation laws.
+This property stems from the fact that the constraint equations contain only first-order time derivatives of the metric tensor, while the dynamical equations contain second-order time derivatives.
+It implies that if the constraint equations are satisfied at an initial time, then the solutions to the dynamical equations and conservation laws will also satisfy the constraint equations at all subsequent times.
+A detailed discussion can be found in Sec.
+7.5 of Weinberg’s textbook [42] on the Cauchy problem.
+The reverse is also true: if solutions to the constraint equations and conservation laws can be found for all times, then they necessarily satisfy the dynamical equations.
+We will now use this latter property to validate the system of equations (37) for scalar perturbations.
+The system (37) encompasses the conservation laws and constraint equations.
+It consists of four first-order ordinary differential equations and one algebraic equation.
+Therefore, it can be solved, and its solutions automatically satisfy the system of dynamical equations.
+This fact will be used to prove the correctness of Eqs.
+(37).
+For $i\neq j$, the dynamical equations (13c) are not coupled to the scalar perturbations. For the case of $i=j$, it is sufficient to consider the contraction of (13c):
 $\ddot{h}^{k}{}_{k}+6H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-3\kappa(\varepsilon_{{\scriptscriptstyle(1)}}-p_{{\scriptscriptstyle(1)}}).$ (143)
-Eliminating the quantity $\dot{h}^{k}{}_{k}$ with the help of (19), yields the following result:
+Using (20), yields
 $(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})+6H(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})-{{}^{3}\!}R_{\scriptscriptstyle(1)}=\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}).$ (144)
-Using the energy density constraint equation (37a) to eliminate the second term of (144), results in the following equation:
+Eliminating the second term of equation (144) using the energy density constraint equation (37a) results in
 $\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)}).$ (145)
-This equation is identical to the time derivative of the energy density constraint equation (37a), as will now be demonstrated. Differentiation of (37a) with respect to time yields
+As will now be shown, this equation is identical to the time derivative of the energy density constraint equation (37a). Differentiating equation (37a) with respect to time yields
 $2\dot{H}(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})+2H(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})=\tfrac{1}{2}\,{{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+\kappa\dot{\varepsilon}_{\scriptscriptstyle(1)}.$ (146)
-Eliminating the time derivatives $\dot{H}$, ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$, and $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ with the help of (10a)–(10c), the momentum constraint equation (37b) and the energy density conservation law (37c), respectively, yields the dynamical equation (145).
+Eliminating the time derivatives $\dot{H}$, ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$, and $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ with the help of (11a)–(11c), the momentum constraint equation (37b) and the energy density conservation law (37c) yields the dynamical equation (145).
 Consequently, Eqs.
 (37a)–(37c) are correct.
-As indicated in the aforementioned verification, the correctness of the equations (37d) and (37e) has not yet been verified.
-This will now be addressed by employing the gauge modes.
-It is observed that Eqs.
-(37) remain invariant under the gauge transformations (1) with the gauge functions $\xi^{\mu}(t,\bm{x})$ given by (6).
-Furthermore, the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\theta_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ are known to be gauge-dependent.
+We will now continue with the second method of verifying Eqs.
+(37).
+Equations (37d) and (37e) have not yet been verified by the above procedure.
+This will now be addressed by using the gauge modes.
+Equations (37) are invariant under the gauge transformations (3) with the gauge functions $\xi^{\mu}(t,\bm{x})$ given by (9).
+Furthermore, the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\theta_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ are gauge-dependent.
 It is therefore evident that the gauge modes
-$\hat{S}_{\scriptscriptstyle(1)}$ $=\psi\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,\theta,p,$ (147a)
-$\hat{\vartheta}_{\scriptscriptstyle(1)}$ =~$=-\frac{\tilde{\nabla}^{2}\psi}{a^{2}},$ (147b)
+$\hat{S}_{\scriptscriptstyle(1)}$ $=\psi\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,T,p,\theta,$ (147a)
+$\hat{\vartheta}_{\scriptscriptstyle(1)}$ $=-\frac{\tilde{\nabla}^{2}\psi}{a^{2}},$ (147b)
 ${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H\left(\frac{\tilde{\nabla}^{2}\psi}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\psi\right),$ (147c)
@@
 (37).
-The gauge modes (147a) are derived directly from (3), with the understanding that $\xi^{0}=\psi$.
+The gauge modes (147a) follow from (5), with the understanding that $\xi^{0}=\psi(\bm{x})$.
 The gauge mode (147b) is the covariant divergence of (29).
-The gauge mode (147c) is derived by substituting the gauge modes (147a)–(147b) into the perturbed Friedmann equation (37a):
+As shown above, the linearized Friedmann equation (37a) is correct.
+Therefore, it can be used to derive the gauge mode (147c).
+Substituting the gauge modes (147a)–(147b) into the linearized Friedmann equation results in
 ${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H(\hat{\theta}_{{\scriptscriptstyle(1)}}-\hat{\vartheta}_{{\scriptscriptstyle(1)}})-2\kappa\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$
 $=4H\left(\psi\dot{\theta}_{{\scriptscriptstyle(0)}}+\frac{\tilde{\nabla}^{2}\psi}{a^{2}}\right)-2\kappa\psi\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (148)
-Using that $\theta_{{\scriptscriptstyle(0)}}=3H$, the Friedmann equation (10a), and Eq.
-(10b), one arrives at the gauge mode (147c).
-It will now be demonstrated that the gauge modes given by (147) are indeed solutions of Eqs.
-(37b)–(37e).
-In order to execute this task, the first step entails deriving the time derivatives of the gauge modes:
+Using that $\theta_{{\scriptscriptstyle(0)}}=3H$, the Friedmann equation (11a), Eq.
+(11b), and the energy density conservation law (11c) yields the gauge mode (147c).
+Expressions (147) imply that the quantities $S_{{\scriptscriptstyle(1)}}$, $\vartheta_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ are invariant under spatial gauge transformations, i.e., transformations for which the condition $\psi(\bm{x})=0$ holds.
+However, they are not invariant under space-time gauge transformations.
+It will now be demonstrated that the gauge modes given by (147) are solutions to Eqs. (37b)–(37e). The first step to doing so is deriving the time derivatives of the gauge modes:
 $\dot{\hat{\vartheta}}_{{\scriptscriptstyle(1)}}$ $=-2H\hat{\vartheta}_{{\scriptscriptstyle(1)}},$ (149a)
@@
 $\dot{\hat{n}}_{{\scriptscriptstyle(1)}}$ $=\psi\ddot{n}_{{\scriptscriptstyle(0)}}=-3n_{{\scriptscriptstyle(0)}}\bigl[\dot{H}-3H^{2}\bigr]\psi,$ (149d)
-where Eqs.
-(10b)–(10d) and Eq.
-(44) have been used.
-Subsequent to the aforementioned preparation, the verification of the accuracy of the equations (37b)–(37e) can be undertaken.
-First, by substituting the gauge modes, namely $\hat{\theta}_{{\scriptscriptstyle(1)}}$, $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}\hat{R}_{{\scriptscriptstyle(1)}}$ into Eq.
-(37b), the result is an equation from which the variable $\psi$ disappears.
-Next, dividing the resulting equation by $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$ and multiplying by $\tfrac{3}{2}H$, one arrives at the time derivative of Eq.
+where Eqs. (11b)–(11d) and Eq. (44) have been used. The validation of the Eqs.  (37b)–(37e) can now be conducted.
+First, substitute the gauge modes $\hat{\theta}_{{\scriptscriptstyle(1)}}$, $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}\hat{R}_{{\scriptscriptstyle(1)}}$ into Eq.
+(37b), to obtain an equation from which the gauge function $\psi$ disappears.
+Next, dividing this equation by $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$ and multiplying the result by $\tfrac{3}{2}H$ yields the time derivative of Eq.
 (11a).
-Second, the substitution of the gauge modes $\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$, $\hat{\theta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq.
-(37c) results in the elimination of the term with $\dot{H}$.
-Next, dividing, the resulting equation by $\psi$ and employing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ followed by utilizing Eq.
+Second, substitute the gauge modes $\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$, $\hat{\theta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq.
+(37c) eliminates the term with $\dot{H}$.
+Next, dividing the resulting equation by $\psi$ and using the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ followed by the use of Eq.
 (11c) yields an identity.
-Third, by substituting the gauge modes $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq.
-(37d), and utilizing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and subsequently using Eq.
+Third, substitute the gauge modes $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq.
+(37d), and using the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and subsequently using Eq.
 (11c) one arrives at an identity.
-Finally, upon substituting the gauge modes $\hat{n}_{{\scriptscriptstyle(1)}}$ and $\hat{\theta}_{{\scriptscriptstyle(1)}}$ into Eq.
+Finally, by substituting the gauge modes $\hat{n}_{{\scriptscriptstyle(1)}}$ and $\hat{\theta}_{{\scriptscriptstyle(1)}}$ into Eq.
 (37e) the term with $\dot{H}$ vanishes.
@@
 (37b)–(37e).
-Consequently, it can be concluded that these equations are accurate.
+Consequently, it can be concluded that these equations are correct.
 Given the established validity of Eq.
-(37a), it can be deduced that the system (37) is also accurate.
-As a final check, it is observed in Appendix C that for a flat flrw universe, the evolution equations (152) for the gauge-dependent density fluctuation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ are obtained from Eqs.
+(37a), it can be deduced that the system (37) is accurate as well.
+As a final check, it is observed in Appendix C that for a flat flrw universe, the evolution equations (151) for the gauge-dependent relative perturbation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ are obtained from Eqs.
 (37).
-Of particular note is the fact that the well-known evolution equations for $\delta$ are recovered in the cases of a radiation-dominated universe (154) and a universe after decoupling of matter and radiation (156).
+In fact, the well-known evolution equations for $\delta$ are obtained in the cases of a radiation-dominated universe (154) and a universe after matter and radiation decouple (157).
 Consequently, it can be concluded that the system of equations (37) has been verified in three distinct ways.
-Appendix C Conventional Evolution Equation for Cosmological Density Fluctuations
-The usual evolution equation for density perturbations in a flat, ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}=0$, flrw universe is given by
-$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\frac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)(1+3w)\right)\delta=0,$ (150)
-where $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$.
-The derivation of this equation has four primary objectives.
-Firstly, it demonstrates the correctness of the system (37), at least in the context of a flat flrw universe.
-Secondly, it is demonstrated that the standard equation (150) is incomplete.
-That is, its solution is a linear combination of only one physical mode and one gauge mode, independent of the scale of a fluctuation.
-Consequently, Eq.
-(150) does not give an adequate description of cosmological density perturbations.
-Thirdly, it is shown that the novel perturbation theory (60) provides all well-known physical solutions for large-scale perturbations in a flat flrw universe.
-Specifically, it produces the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe after decoupling of matter and radiation.
-Finally, it will be demonstrated that Eq.
-(150), when derived from the Newtonian theory of gravitation, yields solutions that remain dependent on the selection of the coordinates.
-In order to derive equation (150), it is necessary to utilize the background equations (71) and the equations (123) for scalar perturbations.
-These two systems of equations are complete, since they contain all conservation laws and constraint equations.
-In this appendix, it is assumed that $c=1$.
-Equation (150) has been derived from an equation of state $p=p(\varepsilon)$ in the existing literature.
-Therefore, the particle number conservation law (71c) and its linearized counterpart (123b) are not required.
-Substituting $p_{n}=0$ into (13) and (40) yields $p_{{\scriptscriptstyle(1)}}=p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}}=\beta^{2}\varepsilon_{{\scriptscriptstyle(1)}}$.
-According to the definition $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$, Eqs.
+Appendix C Conventional Evolution Equation for Relative Density Perturbations
+In this appendix the usual perturbation equations for relative perturbations in a flat, ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, flrw universe is compared with the perturbation theory developed in this article.
+To that end, the evolution equation for the relative perturbation $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$ is derived using the systems of background equations (75) and evolution equations (122) for scalar perturbations.
+In this appendix, we assume $c=1$ and $\Lambda=0$.
+In the existing literature, the particle number density is usually not taken into account.
+Therefore, we take the equation of state for the pressure $p=p(\varepsilon)$, so that $p_{n}=0$}.
+From (14) and (40) we find that $p_{{\scriptscriptstyle(1)}}=p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}}=\beta^{2}\varepsilon_{{\scriptscriptstyle(1)}}$.
+Using the definition $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$, Eqs.
 (122) can be rewritten as follows:
@@
 ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}=0,$ (150c)
-where Eq. (71b) has been used.
-In order to arrive at a generalization of (150), the following steps must be taken.
-First, it is necessary to differentiate equation (151a) with respect to time.
-Subsequently, the time derivatives of $w$, $H$, $\varepsilon_{\scriptscriptstyle(0)}$, $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ must be eliminated.
-This can be achieved by utilizing the equations (44), (71a) and (71b) and the perturbation equations (151b) and (151c).
-The final step in the process is the elimination of ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ through the application of (151a).
-This results in the following system of equations:
+where Eq.
+(75b) has been used.
+First, differentiate equation (150a) with respect to time.
+Then, eliminate the time derivatives of $w$, $H$, $\varepsilon_{\scriptscriptstyle(0)}$, $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ using equations (44), (75a) and (75b) and the perturbation equations (150b) and (150c).
+Finally, eliminate ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ using Eq.
+(150a).
+This results in the system of equations:
 $\ddot{\delta}+2H\dot{\delta}\bigl(1+3\beta^{2}-3w\bigr)$
-$\qquad-\biggl{[}\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\Bigl{(}(1+w)(1+3w)\Bigr{.}\Bigr{.}$
+[…]
 + $\qquad+\,4w-6w^{2}+12\beta^{2}w-4\beta^{2}-6\beta^{4}\Bigr)-6\beta\dot{\beta}H\biggr]\delta$
@@
 $\dot{\vartheta}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\dfrac{\beta^{2}}{1+w}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (151b)
-In this procedure the computer algebra program Maxima [44] has been used.
-The gauge modes (147) are for the system (152) given by
+In this procedure the computer algebra program Maxima [52] has been used. The gauge modes (147) are for the system (151) given by
 $\hat{\delta}(t,\bm{x}):=\dfrac{\psi(\bm{x})\dot{\varepsilon}_{\scriptscriptstyle(0)}(t)}{\varepsilon_{\scriptscriptstyle(0)}(t)}=-3H(t)\psi(\bm{x})\bigl(1+w(t)\bigr),$ (152a)
 $\hat{\vartheta}_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}(t)}.$ (152b)
-The fact that the gauge modes (153) satisfy the system (152) has been shown in Appendix B.
-It can be inferred from the set of equations (152) and the gauge modes (153) that:
-  • The solution of the homogeneous part of Eq.
-(152a) is a linear combination of a physical mode and the gauge mode (153a).
-The particular solution is obtained by solving Eq.
-(152b), and contains the gauge mode (153b).
-  • It is impossible to determine the arbitrary infinitesimal gauge function $\psi(\bm{x})$ through physical means.
-Therefore, the imposition of physical initial values on the system (152) does not yield a physical solution.
-  • In the small-scale limit, the gauge modes (153) do not vanish.
-This means that in the small-scale limit gauge-dependent quantities do not become gauge-invariant.
-Consequently, $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ does not become equal to its Newtonian counterpart $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, given by (56) and (58).
-This is consistent with the gauge transformation (53) of the Newtonian theory of gravitation.
-  • A notable distinction between the conventional equation (150) and Eq.
-(152a) is that the former is homogeneous, whereas Eq.
-(152a) is inhomogeneous.
-This is because the system (152) is derived from the complete set of constraint equations and conservation laws.
-In the case where the right-hand side of Eq.
-(152a) is absent, as is the case in Eq.
-(150), then (153a) is a gauge mode if $\nabla^{2}\psi=0$.
-Consequently, $\psi=C$ is an arbitrary infinitesimal constant, so that the gauge mode (153a) does not vanish.
-  • In the universe following decoupling, the values of $w\ll 1$ and $\beta^{2}\ll 1$ are negligible, allowing for the subsequent neglect of $w$ in relation to the constants of order one in Eq.
-(150).
-This is equivalent to deriving the usual equation (150) from the Newtonian theory of gravitation adapted to the expansion of the universe.
-See, for example, Ref.
-[10], Sec.
-6.2, Ref.
-[31], Sec.
-15.9, and Ref.
-[45], Eq.
-(3.17).
-However, the gauge transformation (53) in the nonrelativistic limit ensures that the gauge mode (153a) does not vanish.
-From these points it is clear that the conventional equation (150) is unsuitable to study the evolution of density perturbations in the universe.
-Given that Eq.
-(150) can be derived from Newtonian gravitation for $w\ll 1$, it follows that the Newtonian theory is unable to provide an adequate description of density perturbations in the universe.
-In the ensuing two sections, a comparison is presented between Eqs.
-(79) and (95), with the system of equations (152).
-This comparison will be conducted in the case of large-scale fluctuations.
-It will become evident that the gauge modes preclude the attainment of physical solutions.
+From the Friedmann equation, $3H^{2}=\kappa\varepsilon_{\scriptscriptstyle(0)}$, it follows that $H\neq 0$ for $\varepsilon_{{\scriptscriptstyle(0)}}\neq 0$.
+This implies that the gauge mode $\hat{\delta}(t,\bm{x})$ is nonzero, regardless of the scale of the density perturbation.
+Consequently, the general solution of the system (151) has no physical meaning.
+Therefore, the system (151) cannot be used to study the evolution of density perturbations.
 C.1 Radiation-Dominated Era
-In this era, the pressure is given by a linear barotropic equation of state $p=\tfrac{1}{3}\varepsilon$, so that $\beta^{2}=\tfrac{1}{3}$ and ${w=\tfrac{1}{3}}$.
-In this case Eqs.
-(152) result in the system
+In this era, the pressure is given by a linear barotropic relativistic equation of state $p=\tfrac{1}{3}\varepsilon$, so that $\beta^{2}=\tfrac{1}{3}$ and $w=\tfrac{1}{3}$}.
+The standard equation is given by
+$\ddot{\delta}+2H\dot{\delta}-\left(\dfrac{1}{3}\frac{\nabla^{2}}{a^{2}}+\tfrac{4}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=0.$ (153)
+In contrast, Eqs. (151) result in the system
 $\ddot{\delta}+2H\dot{\delta}-\left(\dfrac{1}{3}\frac{\nabla^{2}}{a^{2}}+\tfrac{4}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=-\tfrac{4}{3}H\vartheta_{\scriptscriptstyle(1)},$ (154a)
 $\dot{\vartheta}_{\scriptscriptstyle(1)}+H\vartheta_{\scriptscriptstyle(1)}+\dfrac{1}{4}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (154b)
-The gauge modes (153) are solutions of the system (154) for $w=\tfrac{1}{3}$, as can be verified by substitution.
-For large-scale fluctuations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution of the system (154) is
+The gauge modes (152) are solutions to the system (154) for $w=\tfrac{1}{3}$, as can be verified by substitution.
+For large-scale perturbations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution to the system (154) is
 $\delta(t,\bm{x})=(c_{1}t-2\psi(\bm{x})t^{-1})+\tfrac{9}{8}t^{1/2},$ (155a)
 $\vartheta_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}}+\tfrac{9}{8}t^{-1/2},$ (155b)
-where it is used that $H=1/2t$ and $\kappa\varepsilon_{\scriptscriptstyle(0)}=3/4t^{2}$.
-The expression between parentheses in (155a) is the solution of the homogeneous part of Eq.
+where it is used that $H=\tfrac{1}{2}t^{-1}$ and $\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{3}{4}t^{-2}$.
+The expression between parentheses in (155a) is the solution to the homogeneous part of Eq.
 (154a).
 The particular solution $\tfrac{9}{8}t^{1/2}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\tfrac{9}{8}t^{-1/2}$.
-The solution (85) derived from the novel perturbation equation (82) is not derivable from the standard equations (154) due to the presence of the gauge function $\psi$ in the solution (155a).
-Since $\nabla^{2}\delta^{\textnormal{phys}}$ could have been large for small-scale fluctuations, it may have had a large influence on $\vartheta^{{\textnormal{phys}}}_{\scriptscriptstyle(1)}$ and this could have, in turn, a major impact on the evolution of $\delta^{\textnormal{phys}}$.
-That is why (79a) yields oscillating density fluctuations (84) with an increasing amplitude, instead of a constant amplitude which follows from (150).
+The right-hand side of Eq.
+(154a) is not equal to zero, which means that Eq.
+(153) is incomplete.
+As a result, it does not provide all physical solutions.
+Since $\nabla^{2}\delta^{\textnormal{phys}}$ could have been large for small-scale perturbations, it may have had a large influence on $\vartheta^{{\textnormal{phys}}}_{\scriptscriptstyle(1)}$ and this could have, in turn, a major impact on the evolution of $\delta^{\textnormal{phys}}$.
+Because the quantity $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into Eq.
+(82), oscillating relative density perturbations (87) with an increasing amplitude are obtained instead of the constant amplitude that follows from (153).
 C.2 Era after Decoupling of Matter and Radiation
+The standard equation is given by
+$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=0,$ (156)
+where $\beta$ is given by (93).
 In this era, the equation of state for the pressure is according to thermodynamics given by (89).
@@
 $\dot{\vartheta}_{\scriptscriptstyle(1)}+2H\vartheta_{\scriptscriptstyle(1)}+\beta^{2}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (157b)
-The gauge modes (153) are solutions of the system (156) for $w\ll 1$ and $\nabla^{2}\psi=0$, as can be verified by substitution.
+The gauge modes (152) are solutions to the system (157) for $w\ll 1$ and $\nabla^{2}\psi=0$, as can be verified by substitution.
 Consequently, for the system (157) $\psi$ is an arbitrary infinitesimal constant $C$.
-This implies that $\vartheta_{\scriptscriptstyle(1)}=\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is a physical quantity, since its gauge mode $\hat{\vartheta}_{\scriptscriptstyle(1)}$, (153), vanishes identically.
-Consequently, for large-scale fluctuations the solution of the system (156) is
+This implies that $\vartheta_{\scriptscriptstyle(1)}=\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is a physical quantity, since its gauge mode $\hat{\vartheta}_{\scriptscriptstyle(1)}$, (152b), vanishes identically.
+Consequently, the relativistic gauge transformation (3) with $\xi^{\mu}(t,\bm{x})$ given by (9) reduces to the Newtonian gauge transformation (52).
+This is to be expected, since a cosmological fluid for which $w\ll 1$} and $\beta^{2}\ll 1$ can be described by nonrelativistic equations of state (89).
+For large-scale perturbations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution to the system (157) is
 $\delta(t)=(c_{1}t^{2/3}-2Ct^{-1})+\tfrac{7}{9}t^{-5/3},$ (158a)
 $\vartheta_{\scriptscriptstyle(1)}^{\textnormal{phys}}(t)=-\tfrac{7}{9}t^{-4/3},$ (158b)
-where $H=2/3t$, $\kappa\varepsilon_{\scriptscriptstyle(0)}=4/3t^{2}$, and $\beta^{2}\propto t^{-4/3}$ has been used.
-The latter proportionality follows from (91) and (92), $T_{\scriptscriptstyle(0)}\propto a^{-2}$ and (104).
-The expression between parentheses in (157a) is the solution of the homogeneous part of Eq.
+where $H=\tfrac{2}{3}t^{-1}$, $\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{4}{3}t^{-2}$, and $\beta^{2}\propto t^{-4/3}$ has been used.
+The latter proportionality follows from (92) and (93), $T_{\scriptscriptstyle(0)}\propto a^{-2}$, (94), and $a\propto t^{2/3}$, (104).
+The expression between parentheses in (158a) is the solution to the homogeneous part of Eq.
 (157a).
 The particular solution $\tfrac{7}{9}t^{-5/3}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\tfrac{7}{9}t^{-4/3}$.
-The solution (109) derived from the novel perturbation equation (108) is not derivable from the standard equations (156) due to the presence of the gauge constant $C$ in the solution (157a).
-Since $\psi=C$, the gauge transformation (1) with $\xi^{\mu}(t,\bm{x})$ given by (6) reduces to the Newtonian gauge transformation (53).
-This is to be expected, since a cosmological fluid for which ${w\ll 1}$ and $\beta^{2}\ll 1$ can be described by nonrelativistic equations of state (88).
+The right-hand side of Eq. (157a) is nonzero for all scales. Therefore, Eq. (156) is incomplete and does not yield all physical solutions.